A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 April 2024 Version 1.18

Title word cross-reference

J [532]. M [722, 513].

-divergence [532]. -quantile [722, 513].

10th [306]. 12th [44, 300]. 1950s [591]. 1960s [591]. 19th [300, 680]. 1st [686].

2001 [213]. 2012 [300, 306]. 21st [474]. 29th [306]. 2nd [208, 165, 285, 281, 142, 83, 556, 736, 204, 145].

3rd [205, 206, 290, 82, 54].

5th [291, 167].

A. [466]. ability [362, 390, 516]. abolition [543]. absence [614]. absent [674]. absenteeism [773]. abstinence [703]. abuse [318, 176]. Acceptance [231]. Access [780, 508]. accessible [149]. accidents [571]. accounting [593]. accuracy [378, 577, 224, 335]. achievement [721, 679]. achievements [211]. across [334]. Act [636]. activity [95]. Actuarial [873, 468]. acute [333]. Adaptive [608, 792, 575]. added [679, 581]. Addison [65]. Address [569]. addressing [679]. adjust [364, 42]. Adjusting [480, 154]. adjustment [71, 617]. adjustments [322]. administrative [94, 784, 338, 573, 689, 722, 678]. adolescence [748, 418]. adolescent [619]. adult [522]. adulthood [748]. Advanced [444, 201, 399, 872]. Advances [412]. advantage [719]. Adversarial [603]. adverse [382]. advice [865, 492]. affect [541, 671]. affinity [187]. Africa [5, 407]. after [212, 257, 525, 10]. aftermath [219]. again [92]. Age [781, 375, 337, 788, 435, 758, 428, 530, 701, 268, 255, 753, 483, 546, 713]. Age-Period-Cohort [781]. age-specific [435, 753]. age-stratified [255]. aged [573]. Ageing [745]. agencies [745]. Aggregate [339, 129]. aggregated [724]. Agreement [10, 21]. Ahamada [309]. air [794, 127, 678, 450, 404]. Aitchison [621]. Akaike [648, 45]. alcoholism [703]. Alfred [330, 800]. Algebra [782]. algorithm [215]. all-time [453]. Allen [225]. allocation [178, 176, 177]. allocations [767]. alternating [268]. Altman [760]. American [647]. among [449]. analyse [571]. Analyses [55, 393, 69, 248, 405]. Analysing [502, 257, 327, 337, 434]. Analysis [806, 781, 448, 208, 171, 856, 286, 165, 287, 607, 535, 562, 285, 471, 148, 281, 858, 859, 861, 850, 862, 491, 538, 412, 844, 227, 169, 736, 867, 557, 25, 204, 870, 600, 834, 394, 38, 72, 338, 185, 126, 774, 71, 670, 454, 378, 420, 857, 827, 475, 529, 353, 748, 40, 726, 835, 152, 432, 788, 355, 702, 825, 268, 667, 374, 869, 513, 414, 237, 147, 24, 201, 144, 559, 603, 442, 472, 805, 51, 558, 469, 229, 79]. analytic [382]. Analytics [847]. Analyzing [863, 493]. Anatolyev [280]. anchored [774]. Anchoring [184, 179, 181, 180, 376]. Andersen [109]. Anderson [584, 163]. Andrew [134, 137]. Animal [844]. animals [129]. Ann [627]. Anthony [276, 802, 277]. anthropometry [701]. anticipation [258]. Antithetic [374]. AO [551]. application [457, 836, 580, 506, 322, 241, 257, 514, 185, 774, 695, 190, 723, 154, 544, 596, 832, 786, 428, 701, 484, 753, 578, 259, 377]. Applications [165, 56, 447, 762, 577, 617, 85, 164, 850, 634]. Applied [424, 739, 573, 205, 207, 807, 206, 32]. Applying [171]. Approach [845, 11, 115, 447, 600, 128, 394, 318, 511, 650, 212, 599, 750, 499, 619, 375, 327, 785, 420, 611, 124, 729, 748, 449, 833, 669, 235, 786, 221, 332, 757, 430, 699, 572, 725, 335, 711]. approaches [577]. Approaching [285]. approximate [189]. area [646, 829, 650, 514, 500, 157, 364, 648, 455, 122, 647, 486, 640, 73, 484, 718, 215, 649, 639, 717, 594, 642, 725, 505, 528]. areas [94, 643]. Arkansas [674]. Arthur [134, 138, 838, 104, 103]. artwork [190]. Ashby [319]. Ashford [194, 193]. Ashworth [160, 159]. Asia [1]. Ask [751]. Aspects [608]. assess [153]. Assessing [326, 269, 69, 189, 572, 335, 447, 382, 480, 445, 217]. Assessment [296, 823, 242, 408, 700, 698, 575, 721]. assessments [798]. assisted [458]. associated [381, 699]. association [508, 542, 748, 460, 453]. Assortative [363]. asymmetric [542]. asymmetry [597]. Asymptotic [22]. atherosclerosi [508]. attainments [156]. Attention [7]. attractivity [96]. attrition [593, 341]. augmented [454]. Australia [415]. Austrian [363]. Author [630, 14, 102]. authorities [514]. authors [117, 232, 314, 368, 425, 496, 564, 656, 742, 852]. auto [499, 787, 454, 512, 794, 659, 549]. auto-regressive [499, 787, 454, 512, 794, 659, 549]. Autocovariance [779]. Autologistic [700]. automated [697]. auxiliary [755, 668]. Availability [850]. available [95]. Average [410, 670, 465]. award [125]. Aynsley [802].

B

[140, 286, 110, 60, 313, 80, 55, 284, 310, 642]. B-splines [642]. Back [34, 62, 90, 118]. background [790]. Bai [281]. balanced [526]. ballot [437]. Bank [269]. bankruptcy [348]. Barber [54]. Barel [175]. Bartholomew [51, 705]. Baseball [493, 863]. Based [8, 256, 644, 755, 577, 769, 596, 163, 288, 446, 725]. Basic [855, 424, 556, 537]. Basics [844]. basis [129]. Bather [330]. batting [158]. Baye [472]. Bayes [189, 869]. Bayesian [84, 560, 472, 448, 777, 646, 38, 318, 97, 650, 571, 703, 476, 750, 96, 763, 857, 793, 611, 463, 487, 833, 692, 235, 864, 786, 678, 356, 751, 221, 439, 391, 757, 752, 484, 372, 482, 503, 869, 598, 871, 380, 483, 262, 172, 494, 442, 812, 844, 227, 111, 21, 53, 685]. Be [4, 735]. becoming [354]. beer [377]. Beginner [89, 290]. Behavioral [286]. behaviour [754, 258, 221, 752, 674]. behavioural [502, 222]. behaviours [185]. Behind [523]. beholder [416]. Belfast [10]. benchmark [700]. benchmarking [617]. benefit [246, 475, 244, 450]. Berlin [573]. Besag [161, 160]. best [220, 344, 319]. beta [611]. between [508, 542, 382, 480, 477, 748, 520, 271, 69, 515, 498, 414]. between-study [69]. Beyond [638, 516, 523, 297, 781]. Bhargava [197]. Bias [171, 767, 44, 71, 480, 122, 69, 702, 570, 464, 526]. Biases [126, 296, 728, 679]. bidirectional [270]. Big [687, 521, 805, 810, 849, 847, 473, 354, 712]. Binary [79, 514, 793, 596]. Bioconductor [57]. Bioequivalence [60]. Biology [52, 845, 848]. biomarker [577]. Biomedical [810]. Biostatistical [400]. Biostatistics [53, 29]. Biplots [312]. birth [550, 417, 525, 660, 100, 101, 701, 189, 483]. Bissell [800]. Bit [713, 713]. bivariate [599, 817, 154, 434, 832]. Blackwell [134, 135]. block [580, 177]. Blossey [52]. Blunders [686]. boarding [333]. Bob [652]. bodily [415]. Böhning [79]. Book [205, 27, 140, 143, 203, 207, 311, 711, 806, 781, 28, 172, 87, 20, 50, 78, 108, 208, 139, 162, 196, 226, 279, 308, 398, 422, 489, 441, 467, 534, 555, 602, 631, 683, 710, 732, 843, 778, 804, 23, 30, 89, 146, 171, 210, 312, 846, 286, 490, 559, 782, 31, 84, 165, 291, 603, 197, 287, 607, 734, 423, 535, 562, 606, 687, 292, 811, 175, 199, 285, 494, 471, 779, 110, 147, 148, 281, 198, 81, 470, 807, 113, 114, 442, 560, 33, 59, 206]. Book [685, 809, 850, 60, 313, 604, 444, 472, 605, 713, 733, 780, 805, 810, 849, 29, 52, 57, 80, 283, 468, 491, 538, 24, 141, 142, 282, 290, 412, 424, 536, 632, 712, 738, 173, 82, 163, 200, 229, 289, 608, 812, 845, 201, 309, 735, 54, 86, 561, 401, 51, 684, 83, 844, 85, 227, 116, 492, 169, 228, 493, 26, 539, 556, 443, 558, 634, 739, 55, 111, 202, 686, 736, 808, 848, 58, 21, 61, 230, 231, 79, 112, 164, 166, 167, 280, 288, 469, 537, 557, 847, 737, 88, 25, 56, 144, 32, 22]. Book [109, 168, 170, 204, 284, 399, 400, 633, 115, 174, 209, 310, 145, 53, 714, 200]. bookings [460]. Books [534]. boost [592]. Born [530]. Bound [134, 138]. Boyd [17]. Boyle [198]. Boys [840]. brand [377]. branding [377]. Branscum [227]. Brazilian [640]. breaks [511]. Breast [38, 236]. Bretz [228]. Brexit [771]. Britain [128, 340, 320, 475, 415]. British [616, 591]. Broader [431]. Broemeling [21, 53]. Brother [849]. BSc [245]. Bulk [780]. Buonaccorsi [164]. Bureau [569]. burglary [452]. Burkina [188]. Burt [54]. Business [207, 412, 641, 219, 151, 486, 728, 640, 335].

C [210, 292, 198, 313, 289, 228, 111, 61, 32]. call [327, 529]. campaign [373]. Can [632, 10, 720, 666, 70]. cancer [38, 479, 236]. Candela [28]. capacity [155, 814]. Capital [1]. carbon [352]. career [525]. Carli [445]. Carlo [203, 289, 544]. Carpenter [652]. Carr [310]. Carroll [253]. Case [57, 318, 592, 445, 385, 296, 189, 405, 574]. case-study [318, 385, 189, 405]. Casella [289]. cases [261]. Caseworkers [4]. cash [593]. Castle [282]. casualties [364]. Catching [130]. categorical [692]. Causal [684, 614, 834, 670, 316, 520, 825]. Causality [293, 145]. causes [698]. censored [834, 337, 252, 124, 41]. censoring [154]. Census [213, 569, 345, 120]. central [613]. century [239, 474, 680]. Certain [536]. cervical [479]. cessation [385]. chain [462]. Chakraborti [167]. challenges [211, 689]. Chan [165]. Change [2, 506, 186, 753]. Changepoint [471, 870]. changes [754, 613]. changing [320, 797, 449, 120]. characteristics [222, 326]. characterizing [569]. Charitable [239]. charities [354]. Charles [301, 304, 421, 620]. charts [269]. Chernov [199]. child [318, 131, 364, 414]. childhood [746, 522, 416, 333]. Children [73, 457, 417, 517, 592, 131, 362, 360, 513]. China [833]. chlamydia [748, 217]. choice [543, 192, 668, 254, 671, 181]. choices [671]. Christensen [140, 227]. Christopher [134, 137]. Circles [199]. Circular [866, 199, 561]. citation [498]. cited [212]. City [796]. civic [522]. Clarke [80, 651]. Class [7, 836, 485, 343, 616, 185, 500, 190, 216]. classes [830]. Classical [454]. Classification [59, 738, 824, 392, 610]. classifications [121]. classified [595, 594]. Classifying [795]. Clifford [276, 841, 275]. Climate [492, 865]. Climatology [734]. Clinical [60, 608, 58, 318, 257, 252, 70, 792]. closure [699]. Cluster [859, 558, 393, 23]. Clustering [649, 452, 209]. Clusters [114]. co [745]. co-operation [745]. coarser [678]. coding [675]. coefficient [502]. cognitive [417, 517, 480, 362, 360, 578]. coherent [475]. Cohort [781, 517, 513, 154, 428]. cohorts [408]. cointegrated [659]. collected [43]. Collection [780, 668]. combination [222]. Combining [451, 229, 216, 95, 465, 122, 647, 574]. Committees [58]. common [528]. communities [333]. Community [647, 10, 188, 268, 570]. commuting [671]. company [503]. CompanyX [374]. comparative [121, 248, 515]. compare [334]. comparison [784, 71, 378, 150, 723, 544, 524, 250, 123, 344]. comparisons [383]. competing [38, 579, 252, 723]. competition [542]. competitions [478]. complete [581]. completion [516]. complex [318, 825, 259]. complier [670]. component [477]. components [189]. composite [404]. composition [449]. comprehensive [785]. Computation [87, 527]. Computational [873, 52, 468]. Computations [175]. Computer [85, 492, 865]. Computer-simulation [492, 865]. concentration [795]. concentrations [404]. conceptions [149]. concerns [671]. conditional [256]. conditioning [220]. conditions [157, 335, 506]. Confidence [798, 88]. confidentiality [690]. confounding [364]. Congress [530]. Connected [311]. conscious [222]. Consequences [520, 823, 463]. consistency [447]. Consistent [461]. consolidation [599]. constrained [691, 642]. Constructing [127, 645]. Consumer [687, 323, 823, 518, 37, 835, 697, 61]. consumption [518]. contact [241, 220]. container [791]. content [221]. Contents [369, 426, 497, 565, 657, 743, 853]. Contextual [156]. Continuous [724, 260, 832]. contraceptive [752]. Contracts [5]. contributions [745]. control’ [690, 231, 811]. controlled [723]. convenience [699]. conventional [378]. conversational [666]. conversion [822]. copula [817, 446]. copula-based [446]. Corlett [301, 307]. Corporate [1, 822]. correction [180]. Correlated [739, 455]. Correlates [728, 42]. Correspondence [630, 14, 420, 100, 274, 102, 224, 12, 629]. Corrigendum [854, 655]. Cost [13, 393, 670, 131, 529, 432, 12, 405]. cost-effectiveness [670, 405]. costs [544, 503]. cotinine [131]. Council [366, 419, 99, 223, 299, 86]. Count [490, 860, 523, 673, 462, 832]. count-continuous [832]. counterfactual [243]. Counting [358]. countries [97, 261, 720]. country [241, 431, 248]. counts [500]. county [755]. Course [84, 858, 560, 557, 864, 67, 166]. Covariance [779]. covariate [387, 668]. covariates [829, 397]. Coverage [213, 431, 569]. Cowpertwait [283]. Coxon [276, 277]. Crawley [200]. Creating [835]. Credit [61, 822, 823, 818, 820, 819, 815]. Creighton [82]. cricket [464, 158]. crime [531, 507, 452, 790]. Crimean [655, 402]. Crisan [284]. crises [409]. criterion [648]. Critical [13, 12]. critique [177]. crop [755]. Cross [10, 354, 241, 595, 150, 339, 520, 726, 701, 248, 792, 594]. cross-classified [595, 594]. Cross-Community [10]. cross-country [241, 248]. cross-fertilization [792]. cross-lagged [520]. cross-national [726]. cross-sectional [354, 150, 339, 701]. Crossroads [86]. Crowds [311]. Cryer [165]. CStat [245]. Cultural [65]. Culture [846]. cumulative [662]. Cunliffe [301, 300]. cure [38, 618]. customer [334]. Cyber [535]. Cyber-security [535]. cycle [415]. cytology [378].

D [311, 28, 31, 84, 165, 291, 292, 29, 163, 229, 54, 51, 231, 79, 288, 144, 170, 284, 310, 53]. D. [530]. daily [257, 659, 697]. darts [224, 132]. Darwin [64]. DasGupta [22, 166]. Data [31, 287, 607, 562, 687, 147, 148, 858, 442, 860, 861, 9, 805, 849, 80, 862, 538, 412, 712, 83, 844, 227, 668, 169, 493, 736, 867, 79, 557, 847, 144, 633, 310, 600, 828, 749, 403, 814, 502, 485, 94, 784, 595, 257, 703, 545, 338, 393, 474, 696, 327, 568, 500, 337, 429, 577, 363, 857, 458, 451, 397, 770, 573, 521, 689, 529, 747, 723, 459, 70, 124, 544, 524, 381, 748, 462, 432, 832, 692, 863, 756, 345, 43, 728, 431, 640, 100, 101, 436, 321, 701, 680, 693, 374, 671, 645, 482]. data [473, 187, 95, 694, 383, 503, 720, 578, 259, 724, 820, 344, 217, 377, 237, 406, 574, 775, 790, 822, 765, 690, 171, 313, 51, 25, 806, 490, 559, 535, 780, 810, 739, 808, 58, 469]. Dataset [28]. David [134, 282, 585, 799, 801, 135, 421, 705]. Davies [47]. Davis [109]. dead [358]. death [381, 427, 373]. debt [661]. decay [432]. decision [588, 383]. decisions [212]. decomposition [241, 527, 414]. decreased [693]. default [37, 816, 822, 821]. Deficit [7]. Defining [13, 12]. delays [436]. demand [372]. demographic [835, 427, 141]. density [665, 750, 573, 435]. department [336, 176]. Dependence [600, 817, 364]. dependences [392]. dependent [457, 831, 618]. depression [619, 337, 748]. deprivation [768, 364]. Derek [731]. derived [269]. Deriving [486]. Derrick [276, 278]. Describing [9]. Design [213, 82, 394, 218, 420, 793, 437, 520, 323, 531, 792, 725, 146, 288]. design-based [725]. designs [316, 272, 825, 575, 594]. destination [794, 439]. Detecting [595, 100, 101, 501]. Detection [114, 870, 436, 234, 471]. determinants [128, 576]. Determining [519]. deterministic [766]. developing [833, 720]. Development [345, 239, 517, 592, 362, 360]. diagnoses [318]. diagnosis [541]. diagnosis-related [541]. Diagnostic [53]. Diagnostics [438]. Dickinson [167]. did [828, 340]. Dietz [81]. Difference [686]. differences [241, 417, 156, 255, 792, 515, 180]. different [529, 772]. Differential [168, 693]. Differentially [691]. differentials [332]. difficulties [513]. difficulty [218]. Diffusion [433]. Digital [713, 636]. Dimensional [281, 862, 8, 469, 827]. Direct [677, 235]. disability [182, 500, 475, 647]. disabled [600]. disadvantaged [362]. disclosure [690, 693]. discontinuities [505]. discontinuity [375, 793, 437, 543, 531, 272]. Discovering [681]. Discrete [604, 327, 154, 40, 699]. Discrete-valued [604]. discrimination [501]. Discussion [251, 178, 765, 319, 317, 267, 249, 233, 149]. Disease [172, 209, 261, 568, 487, 436, 234]. disease-transmitting [459]. diseases [333]. Disney [803]. Disorder [7]. disparities [189]. displacement [704]. Display [858, 557]. dissimilarity [515]. dissolution [68]. distance [432]. distribution [396, 770, 459, 158, 243, 821]. distributional [332]. Distributions [782, 634, 499, 756, 482, 791]. district [407]. divergence [532]. dividends [72]. divorce [363]. Dixon [48]. Dmitrienko [228]. Do [541, 475, 321, 671, 212, 188, 388, 373]. doctor [751]. Does [361, 592, 526]. domain [641, 181]. domestic [269, 747]. dominant [354]. doubly [834]. Douglas [760]. drinking [703, 185, 268]. driven [438, 458, 240, 570]. drivers [417]. dropout [252]. Drug [846, 664, 723, 381, 758, 751, 405]. DSc [273]. Dudoit [56]. due [511, 487, 184]. Dugard [285]. Dunn [839]. Dunne [625]. Dunstan [319]. duration [128, 834]. Durbin [301, 302]. during [522, 297, 668, 701, 675]. Dutch [650, 481, 766]. dyadic [447]. Dyke [301, 306]. dynamic [293, 37, 768, 827, 433, 434, 258, 501, 662, 313]. Dynamics [182, 619, 353, 339, 757, 676].

Early [758, 185, 123, 389]. earn [828]. earnings [784, 150, 676]. Easley [311]. East [1]. eat [384]. eBay [190]. Ecological [9]. Ecology [845, 116]. econometric [72, 540]. Econometrics [282, 309, 280, 174]. Economic [836, 157, 266, 348, 520, 391, 532, 415, 95, 140]. Economics [197, 207]. Economy [636]. ecotoxicological [242]. Editorial [211, 358, 246, 315, 92, 36, 636, 521, 716, 331, 93, 473, 413, 212]. Editors [370, 212]. edn [205, 208, 165, 291, 285, 281, 206, 142, 290, 82, 54, 83, 556, 167, 204, 145, 736, 686].

Education [5, 183, 746, 121, 338, 477, 40, 669, 716, 198]. Education-related [183]. Educational [413, 721, 156, 73]. Edward [837]. Effect [5, 11, 704, 614, 379, 823, 834, 746, 343, 550, 218, 364, 465, 519, 437, 681, 531, 756, 530, 679, 667, 243, 418, 819]. effectiveness [393, 670, 581, 405]. Effects [11, 116, 461, 595, 365, 96, 619, 670, 548, 465, 677, 258, 615, 455, 507, 824, 390, 272, 668, 404, 73, 578, 513, 430, 594, 666]. efficiency [294]. efficient [224]. effort [322, 673]. ego [447]. Ehrenberg [134, 137]. Eigenvalue [175]. Elbers [644]. elderly [463]. election [92, 519, 698, 674, 766]. Elections [9, 767]. Elegant [147, 148]. Elementary [54]. Elements [83]. elicitation [394, 420]. Embedded [3, 826]. Embedding [769]. emergency [336, 590]. Empirical [9]. employment [523, 271, 237]. end [346, 611]. Endeavours [1]. endemic [261]. endogeneity [695]. engagement [522]. engaging [395]. Engineering [856, 850, 867, 394, 420, 562, 491]. Engineers [205, 206]. England [94, 213, 269, 381, 678, 613, 547, 95, 513]. English [784, 44, 696, 721, 177, 434, 296, 355, 589, 333, 210]. enhanced [577]. Enhancing [437]. Enlightenment [64]. enrolment [356]. ensemble [41]. Enterprise [113]. enterprises [819]. entry [510]. Environmental [292, 408, 450, 671]. environments [762]. epidemic [409]. Epidemics [333]. Epidemiologic [171]. epidemiological [510]. Epidemiology [172, 292]. epilepsy [723]. equal [823]. Equation [290, 578]. Equations [168, 695]. equity [377]. eradication [410]. Ergodic [714]. Erich [74]. Ernest [709]. Error [164, 828, 646, 326, 644, 573, 459, 481, 520, 244, 643]. errors [698, 569, 686]. establishment [826]. estate [214]. estimate [465, 437, 678, 574]. estimates [644, 755, 548, 122, 647, 486, 224, 243].

Estimating [830, 129, 190, 125, 573, 797, 381, 173, 701, 450, 662, 39, 670, 487, 728, 645, 427, 483]. Estimation [98, 71, 280, 646, 294, 461, 94, 650, 334, 750, 514, 831, 644, 454, 641, 573, 463, 154, 124, 342, 455, 588, 268, 484, 718, 215, 649, 639, 725, 505, 528, 684]. estimator [506]. Ethical [805]. Ethics [810]. ethnic [508, 573, 449, 156, 189, 547]. euro [157]. EUROMIND [157]. Europe [98, 182, 745]. European [506, 97, 719, 661, 2, 352, 546]. Evaluating [749, 396, 260, 272, 516]. Evaluation [9, 776, 121, 257, 266, 391, 667]. event [186, 327]. Events [13, 749, 382, 816, 67, 12]. Everitt [286, 55]. Evett [251]. evidence [614, 518, 542, 478, 727, 854, 541, 787, 71, 517, 150, 131, 363, 591, 719, 769, 360, 244, 531, 385, 431, 254, 250, 464, 415, 187, 822, 395, 163]. Evolution [1, 600]. evolving [247]. examination [325]. Example [537, 520, 120]. Examples [858, 560, 557, 42, 864, 82, 168]. exceedance [549]. excellence [716]. exchange [665, 597, 498]. exchangeability [242]. exercise [508]. Exercises [400]. existing [321]. exit [92]. expectancy [39]. Expenditure [591, 518]. expenditures [518, 572]. experience [216]. experiment [576, 769, 826, 323, 720, 674]. Experimental [776, 316, 394, 420, 465]. Experiments [3, 341, 82, 288]. expert [394, 97, 256, 420]. Explained [632]. Explaining [241, 828]. exploration [70]. Exploratory [805]. explore [328]. Exploring [581, 8, 120, 529, 336]. Exposure [539, 822, 487, 428, 678]. Extensions [116]. Extensive [615]. extent [188]. extracts [569]. extreme [187]. eye [416].

F [89, 57, 141, 282, 51, 116, 228, 111]. facilities [508, 822]. facility [543]. facing [383]. factor [502, 382, 454, 528]. factor-augmented [454]. Factors [699, 188, 189, 822, 237]. factors-evidence [822]. Fairness [320]. Falcon [57]. Familial [313]. Family [325, 550, 522, 73, 255, 821]. fan [269]. fans [334]. Farr [579]. Faso [188]. Faustian [36]. Fay [646]. FBA [466, 302]. features [745]. Federer [76]. fee [543]. Fellow [273]. female [236]. fertility [447, 834, 477, 435, 753]. fertilization [792]. Festschrift [282]. fetal [577]. few [383]. FFA [245]. field [322, 720]. fieldwork [323]. Fieller [707]. Fienberg [622]. File [569]. filter [747]. Filtering [284]. Finance [203, 423]. Financial [562, 867, 109, 170, 218, 409, 549, 819, 406]. Finding [258]. findings [674, 594]. fine [820]. fine-grained [820]. fingerprint [250]. finish [717]. Fink [171]. Finney [799]. firm [339]. firms [72, 388]. First [84, 166, 525]. Fisk [49]. Fitting [199, 397, 435]. fixed [461, 619, 667]. Flachaire [309]. flawed [445]. flaws [716]. flexibility [773]. Flexible [832, 435]. Florence [579, 402, 655]. Flourishing [632]. flow [98]. flows [794, 93, 95, 574]. fluctuations [339]. fluidity [616]. flux [522]. focus [721]. Foes [4]. Fokoué [80]. follow [826, 385, 321]. follow-up [826, 385]. follow-ups [321]. Food [197]. foot [261]. foot-and-mouth [261]. football [334, 719, 507, 460, 453]. footballers [390, 516]. Force [574, 788, 455, 65]. forced [704]. forecast [549]. Forecasting [499, 186, 747, 352, 659, 777, 749, 97, 454, 827, 434, 428, 589, 697, 372, 262, 141]. forecasts [665, 256, 269, 266, 786, 41]. foreign [771, 235, 676]. Foreman [319]. forensic [250]. Form [379]. formulae [176]. fortunes [320]. Foundations [714]. four [177]. four-block [177]. Fox [171]. fraction [190]. fractionally [659]. frailty [477]. framework [487, 345, 716, 717]. Francis [301, 706, 304]. Frank [709]. Frederick [800]. free [618, 39]. Freeman [682]. frequency [703, 476, 261, 512, 747, 406].

Frequentist [494, 871]. frequently [588, 578]. Friedman [83]. Friends [4]. Front [35, 63, 91, 119]. FRS [161]. FSS [245, 273]. fully [762, 692]. function [379, 435, 578, 662]. Functional [25, 600, 646, 499]. Fundamentals [399, 166]. Future [6, 788].

G [203, 208, 134, 89, 113, 24, 141, 290, 289, 54, 116, 26, 55, 111, 231, 133, 109, 204]. Galbraith [51]. Galton [64]. Gani [552]. Gaussian [739, 777, 439]. Gavin [551]. Gender [478, 9, 332, 255]. Gender-specific [255]. General [694, 92, 481, 698]. generalizability [668, 153]. Generalized [88, 451, 260, 648, 435, 725]. Generalizing [727, 854]. generated [221]. Generating [693]. generation [692]. Genetic [684]. Genetics [229]. GenitoriPiù [373]. Genomics [229, 56]. Gentleman [57]. genuine [190]. geocoded [693]. geodemographic [121]. Geoffrey [651, 682]. Geographers [54]. Geographic [114, 437, 531, 336]. Geographical [2, 345]. geographies [678, 120]. Geometry [202]. georeferenced [573]. George [301, 306]. Geostatistical [408, 451]. German [836, 44, 388, 237]. Germany [415]. gestational [701]. Getoor [110]. Getting [92]. ggplot2 [148, 147]. Gibbons [167]. Gini [342]. Girosi [141]. give [475]. given [821]. giving [239]. global [787, 822]. goal [390]. goals [460]. Godambe [582]. Goldstein [229]. Gompertz [65]. Good [633, 126, 431, 15]. Gopal [134, 136]. Gospodinov [280]. Goupy [82]. Government [780, 86, 177]. Gower [841]. gradient [416]. gradings [125]. graduates [784]. Graham [760, 624, 839]. grained [820]. grant [125]. grants [272]. granular [835]. Graphical [8, 538]. Graphics [764, 147, 148, 287, 401, 169, 736]. Gray [245]. Great [128, 475]. greatest [453]. Greenacre [312]. Greenfield [802].

Grenander [553]. gridded [432]. Griffin [104, 106]. gross [269, 747]. group [343, 334, 156, 392]. groups [541, 356]. growing [814]. Growth [325, 836, 577, 747, 339, 235]. Guerra [229]. Guide [143, 781, 89, 734, 290, 712, 144, 633].

H [23, 89, 285, 147, 113, 466, 80, 111]. H1N1v [297]. habit [130]. Hahne [57]. Hall [551]. Halley [191]. Halsey [466]. Hamada [111]. Handbook [859, 55, 284, 291, 809, 604, 201, 558, 634, 739, 847, 109]. Handedness [360]. Handling [393, 412, 524]. Hanson [227]. Happiness [632]. Harold [134, 135]. Harry [682]. Hastie [83]. Hate [470, 868]. Hawkes [790]. Hawkes-type [790]. Hayes [23]. Hazards [27]. head [523]. Heads [597]. Health [846, 745, 176, 611, 66, 541, 540, 396, 548, 131, 463, 797, 360, 416, 520, 431, 450, 391, 404, 248, 413, 704, 183, 176, 197]. Healthcare [543, 542, 126, 270, 544, 446, 233, 572, 183, 373]. Healy [583]. heat [384]. heat-or-eat [384]. Heavy [711, 544]. heavy-tailed [544]. Hendry [282]. Henry [301, 304, 190]. herd [129]. heroin [381]. heroin-related [381]. Herriot [646]. Herson [58]. heterogeneity [669, 69]. heterogeneous [385, 351, 183]. heteros*edasticity [331]. heteroscedastic [41]. hidden [462]. Hierarchical [172, 427, 650, 500, 235, 356, 484, 372, 377]. High [862, 7, 469, 361, 212, 827, 362, 271, 406]. High-Dimensional [862]. higher [663, 361, 121, 338, 716]. Highly [311, 212]. Hilda [273, 47]. Hinkley [801]. Hirotugu [45]. History [806, 327]. Hockey [230]. Hoff [84]. Hogg [205, 206]. Holland [708]. Home [464, 719, 68]. homelessness [758]. homes [362]. Hon [245]. Hong [788]. Honour [282]. Horgan [85]. Horton [287]. hospital [542, 254, 372]. hospitals [541, 681, 589]. house [68]. Household [6, 591, 506, 218, 220, 188, 475, 410, 323, 671, 395]. households [661]. Hox [201]. Huber [57, 142]. Hugh [837]. Human [632, 448, 403, 479]. Humanizing [687]. hurdle [336, 446]. Hurricane [219, 734]. hybrid [671]. Hyperactivity [7]. Hypothesis [471, 870, 126].

Iacus [168]. Ian [759]. Ideas [855, 227, 556]. Identification [781, 457, 403, 351, 452]. Identifying [577, 316, 390]. identity [547]. Ieno [89, 116]. ignorable [337, 726, 39]. II [556, 606]. III [75]. illegal [664]. illicit [758]. illness [297]. Illusion [230]. illustrations [405]. imagery [65]. Imai [317]. immersive [762]. immigration [97, 94]. immunodeficiency [448, 403]. impact [244, 359, 255]. Implementation [539]. implications [479, 788, 120, 594]. imply [526]. import [597]. improve [94, 321, 720, 666, 596]. Improved [270]. Improving [647, 372]. imputation [830, 124, 524]. incentives [541, 673, 720, 674]. incidence [407, 662]. include [397]. including [372]. Income [506, 647, 614, 448, 828, 343, 416, 43, 414]. inconsistently [385]. incorporating [755, 381]. increase [323]. increases [243]. increasing [666]. Index [117, 232, 314, 368, 425, 496, 564, 656, 742, 852, 741, 342, 445, 697]. India [752, 720]. indicator [157]. indicators [769, 270, 127, 455, 404, 645]. indices [216]. indirect [677]. individual [748]. indoor [450]. induced [570]. industries [795]. inequalities [797, 415, 413]. Inequality [188, 448, 130, 343, 124, 532, 676, 572]. inequity [183]. Inertia [1]. infant [373]. infection [403, 479, 748]. infectious [568, 436, 234]. Inference [9, 729, 163, 83, 428, 280, 168, 818, 458, 527, 124, 520, 351, 439, 724, 140, 167]. inferences [149]. inflated [821, 569]. inflation [499, 269, 353]. influences [220, 773]. influenza [297]. influenza-like [297]. inform [192]. Information [566, 587, 609, 635, 658, 672, 688, 715, 744, 761, 783, 813, 774, 2, 229, 13, 542, 755, 371, 381, 648, 486, 372, 12, 446]. Information-anchored [774]. informative [154, 387, 718]. inhomogeneous [768]. initiatives [437]. injecting [381, 405]. injuries [830]. INLA [812]. innovations [593, 433]. insect [459]. insecticide [396]. insecticide-treated-net [396]. Insight [687]. institutes [645]. Institutions [235, 388]. instrument [550]. instrumental [619, 729, 596]. instruments [518, 695]. insurance [388, 832, 503, 775, 262, 203]. intactness [522]. Integrated [546]. integration [255]. Intelligence [412]. intensity [703, 216, 418]. intensive [257, 615]. interaction [786, 254]. interactions [397]. interference [364]. interfirm [819]. Intergenerational [343]. Intermediate [858, 557]. International [721, 296, 98, 464]. Interpretable [610]. interrelated [482]. interval [665, 252]. interval-censored [252]. Intervals [88, 259]. intervention [450]. interventions [396, 385]. interview [675]. Interviewer [324, 430, 595, 326, 220, 327, 594, 325, 666]. interviewers [745]. interviewing [666]. Introducing [289]. Introduction [423, 472, 85, 227, 288, 179, 617, 869, 285, 110, 81, 24, 82, 174]. Introductory [283]. invalid [695]. inverse [727, 854, 439]. Investigating [222, 152, 212]. investment [771, 695, 235]. invitations [324]. involvement [361]. Ireland [295, 10, 120]. irreversible [406]. Irving [15]. Isaac [340]. isolated [333]. Issue [566, 587, 609, 635, 658, 672, 688, 715, 744, 761, 783, 813]. issues [405]. Italian [541, 824, 373]. Italy [542, 254]. Item [809, 485, 328, 726]. Izenman [59].

J [205, 27, 140, 311, 28, 301, 307, 586, 653, 23, 165, 291, 287, 285, 281, 113, 59, 206, 24, 142, 282, 173, 82, 200, 201, 54, 51, 83, 85, 116, 169, 58, 164, 167, 56, 109, 115, 174, 145]. J.-P [109]. Jamaican [214]. James [301, 840, 707, 302, 706]. January [300]. Jeffers [301, 303]. Jeffrey [842]. Jim [586]. JMP [82]. job [773]. John [134, 301, 585, 583, 731, 15, 705, 160, 194, 803, 138, 330, 159, 193, 303, 841, 621]. Johnson [227]. Johnstone [267]. Joint [252, 460, 703, 817, 723, 479, 751]. Jon [104, 107]. Jones [60]. Joseph [48, 552]. journal [212]. journals [485, 498]. Julian [160, 194, 161, 195]. jumping [478].

K. [17]. Kalof [81]. Kanji [134, 136]. Katrina [219]. Keen [169]. Kendall [202]. Kennedy [170]. kernel [573]. Khuri [112]. Kiefer [140]. King [141]. Kirchgassner [24]. Kleinberg [311]. Kleinman [287]. Klep [113]. know [373]. knowledge [97]. Kong [788]. Korn [203]. Kreiss [109]. Kroisandt [203]. Kuhnert [79]. Kulkarni [208, 204].

L [23, 171, 110, 81, 282, 82, 54, 21, 61, 310, 53]. L. [530]. Laan [56]. Labour [574, 754, 550, 66, 788, 455]. lag [429]. lagged [520]. Lajos [554]. landmarks [257]. Lanjouw [644]. Large [281, 346, 747, 789, 408, 822, 779]. Lash [171]. last [828]. late [418]. Latent [185, 726, 7, 777, 836, 485, 830, 768, 500, 190, 475, 748, 618]. Lattice [31]. law [823, 568, 791]. Lawley [276, 278]. Lawrence [28]. Lawson [172]. Lazar [25]. lead [361]. leaders [530]. league [192, 434]. Leamer [143]. learn [70]. Learning [83, 399, 350, 418, 344, 28, 110, 59, 80, 738]. Least [199]. leavers [365]. Ledolter [205, 206]. Lee [173]. leg [719]. Legislative [674, 462]. legislators [772]. Lehmann [74]. lending [817]. lens [772]. Leo [74]. LeSage [174]. Leslie [194, 195]. less [526]. level [829, 322, 511, 328, 755, 722, 431, 407, 649]. level-of-effort [322]. levels [749]. Lewis [19, 264, 263, 253]. Liberty [36]. Lies [686]. Life [806, 696, 163, 576, 191, 611, 389, 67, 415, 515, 39]. lifetime [572]. like [297]. likelihood [691, 79, 845]. limitations [192]. limited [100, 101]. line [221]. Linear [175, 199, 738, 288, 447, 451, 397, 189, 649, 406, 725, 262, 782, 112]. Linehan [623]. liner [791]. Lines [199]. link [614]. linkage [728]. linked [338, 835]. linking [522]. liquid [378]. literacy [645]. literature [321]. Liverpool [123]. Living [506, 545]. Livingstone [144]. loans [37]. Local [86, 432, 94, 514, 548, 177, 722]. location [595]. Log [266, 447]. log-linear [447]. Log-optimal [266]. Logistic [13, 756, 268, 752, 12, 594]. logit [461, 549]. Lomax [290]. London [379, 449]. Long [2, 746, 365, 548, 525, 463, 678]. long-run [746, 548, 525]. Long-Term [2, 365, 463, 678]. Longitudinal [360, 313, 513, 354, 257, 703, 337, 577, 150, 341, 364, 252, 723, 154, 216, 701, 356, 680, 415, 578, 259, 217, 430, 775]. Look [10, 191]. lose [597]. loss [155, 821, 262]. losses [611]. Lotteries [198]. lottery [222]. low [271]. loyalty [334]. LRAM [273]. Luke [298]. Luxembourg [676]. luxury [126]. Lynn [421].

M [140, 28, 87, 171, 312, 47, 175, 113, 142, 200, 54, 85, 116, 111, 288, 56, 168]. MA [759]. Maarseveen [113]. Machine [28, 80, 738, 399, 344]. Maclean [759]. Macmillan [276, 277]. Macroeconomic [143]. mail [776]. Making [259, 359, 383, 775, 673]. malaria [410]. Malawi [396]. Mallin [651]. Malvern [225]. man [44]. Management [287, 361, 549]. Managing [633]. Manchester [123]. Manifold [59]. Manual [29]. manuscripts [212]. Many [10]. map [722]. Mapping [172, 459, 643]. March [306]. margin [615]. Marijuana [7]. Mark [552, 298]. market [754, 66, 541, 352]. Marketing [687]. Markets [311]. Markov [663, 462]. Marriott [301, 304]. Martz [111]. Mary [273]. Master [340, 569]. Mastronardi [175]. Mata [735]. match [434]. matches [827, 507]. matching [437, 152]. material [768]. maternal [701]. Mathematical [855, 556]. mating [363]. Matrices [175, 779, 439, 281, 779]. Matrix [782, 175]. Matter [34, 62, 90, 118, 35, 63, 91, 119, 188, 388]. matters [504, 515]. Matthew [298]. may [397]. MBAs [143]. MBE [759, 273, 300]. McDonald [301, 305]. McIntosh [319]. mean [644, 362]. mean-squared [644]. means [646]. measured [578]. measurement [211, 828, 646, 326, 785, 573, 520, 244, 120, 164]. measurements [524, 532]. Measures [348, 21, 448, 784, 456, 150, 42, 127, 244, 694, 515]. Measuring [124, 11, 589, 768, 449, 507, 567]. Mechanics [52]. mechanisms [316]. mediating [746]. mediation [293, 825]. Medicine [53]. medium [819]. medium-sized [819]. Meesters [89]. Meeting [687, 178, 765]. member [221]. membership [392]. Mendelian [684]. Mendenhall [75]. meningitis [407]. mental [416]. Mesopotamia [214]. mesothelioma [428]. Messenger [11]. meta [126, 71, 378, 69, 702, 667, 229, 79]. meta-analyses [69]. meta-analysis [71, 378, 702, 667, 229, 79]. meta-regression [126]. Metcalfe [283]. meteorological [372]. method [644, 244, 482, 344, 181]. methodological [405]. Methodology [282, 394, 96, 420, 190, 112]. Methods [203, 87, 670, 292, 712, 173, 289, 845, 684, 26, 539, 21, 13, 164, 209, 872, 241, 341, 544, 524, 596, 12, 233, 234, 871, 175, 30, 84, 32, 494, 444, 280]. metric [457, 733]. Mexico [332, 796]. Michael [583, 653, 838]. Microbiology [29]. microdata [697]. Micromaps [310]. MIDAS [429, 429]. migrant [836]. migration [98, 128, 96, 786, 93, 95, 546, 574]. Mike [628]. Mikosch [109]. mill [487]. Millennium [517, 513]. Mind [113]. mine [487]. minimum [375, 243]. Mining [80, 83, 344]. minorities [573]. minority [592]. Mint [340]. misclassified [829, 237]. Missing [387, 403, 393, 337, 397, 459, 70, 524, 718, 39, 607]. Misspecification [756]. Mitchell [627]. Mixed [512, 313, 648, 595, 476, 71, 429, 611, 364, 747, 481, 756, 824, 390, 649, 725, 116]. mixed-mode [481]. mixture [38, 821]. mixtures [618, 439]. mobile [645]. mobility [362, 40, 271, 67]. mode [481]. Model [755, 112, 829, 155, 836, 461, 222, 38, 571, 595, 616, 817, 476, 787, 768, 500, 177, 458, 827, 797, 154, 479, 794, 434, 691, 618, 833, 455, 722, 643, 390, 428, 678, 336, 659, 751, 351, 372, 446, 503, 41, 217, 594, 377, 725, 262, 790, 163]. model-assisted [458]. Model-based [755]. Modeling [172, 87, 286, 860, 850, 812, 539, 290, 208, 490, 204]. Modelling [140, 664, 261, 770, 2, 33, 122, 436, 221, 739, 404, 93, 676, 821, 447, 612, 508, 212, 703, 327, 568, 793, 252, 748, 726, 462, 449, 356, 389, 697, 757, 752, 671, 590, 460, 598, 498, 430, 699, 796, 546, 569, 815, 685]. Models [203, 806, 782, 313, 738, 116, 7, 61, 164, 665, 777, 502, 663, 580, 485, 762, 96, 456, 37, 382, 454, 641, 512, 451, 397, 364, 723, 408, 270, 726, 615, 700, 432, 648, 832, 756, 824, 640, 679, 816, 484, 189, 427, 578, 549, 344, 821, 649, 392, 65, 642, 610, 376, 528, 178, 288, 88, 170]. models-estimation [454]. moderate [789]. Modern [734, 113, 24, 845, 848, 280, 115, 30, 59, 608, 848]. modification [681]. Molchanov [202]. monetary [787]. money [340]. Monitoring [58, 773, 259]. Monte [289, 544, 203]. Montford [230]. month [417]. month-of-birth [417]. monthly [157]. Moore [134, 190, 133]. Morris [104, 105, 288]. Morrison [586]. mortality [614, 448, 777, 130, 186, 696, 214, 681, 833, 428, 100, 101, 65, 483]. Mother [525]. Moulton [23]. Moustaki [51]. mouth [261]. MRI [25]. much [340, 70]. Muller [530]. Multi [508, 8]. Multi-Dimensional [8]. Multi-ethnic [508]. multigroup [726]. Multilevel [650, 382, 462, 201, 578, 580, 212, 327, 397, 449, 268, 189, 430, 594]. multimodal [756]. multinomial [543, 455]. Multiperil [775]. Multiple [680, 392, 830, 451, 124, 524, 42, 385, 122, 507, 380, 414, 228, 56]. multiple-classification [392]. multiple-imputation [124, 524]. Multiple-membership [392]. multiple-response [380]. multiplex [580]. multirate [588]. Multisite [825]. Multistate [806]. multisurvey [475]. Multivariable [286]. Multivariate [646, 511, 286, 31, 285, 59, 789, 862, 51, 469, 571, 456, 641, 397, 573, 797, 462, 692]. mustard [814].

N [140, 28, 301, 89, 287, 175, 199, 282, 116, 303, 280, 25, 32]. Nankivell [329]. Nason [26]. nation [431, 487]. National [846, 360, 86, 325, 726, 645, 259, 176]. nationwide [396]. natural [674]. Nature [536, 492, 797, 865]. Navajo [487]. nearest [572]. neighbour [572]. neighbourhood [326, 67]. Neil [624]. Nelder [160, 159]. neonatal [701]. nested [643]. net [396]. Netherlands [113, 830]. Network [535, 458, 612, 155, 580, 702, 392]. networks [580, 433, 570, 311]. Neubauer [231]. Neumann [251]. neutral [750]. never [545]. Newell [585]. Newton [340]. Nicholas [707]. night [92]. Nightingale [655, 579, 402]. Nixon [626]. noise [151]. Non [818, 719, 296, 739, 642, 322, 242, 417, 337, 190, 397, 522, 772, 544, 826, 42, 726, 692, 728, 321, 184, 271, 757, 825, 526, 718, 430, 377, 395, 406, 262, 335, 39, 309]. non-cognitive [417]. non-employment [271]. non-exchangeability [242]. Non-Gaussian [739]. non-genuine [190]. non-ignorable [337, 726, 39]. non-intactness [522]. non-linear [397, 406, 262]. non-linkage [728]. non-normal [544]. Non-parametric [818, 719, 642, 772, 692, 377, 309]. non-random [825, 335]. non-respondents [395]. Non-response [296, 322, 826, 42, 726, 321, 184, 825, 526, 718, 430]. non-state [757]. Nonlinear [284]. Nonparametric [291, 32, 167]. normal [641, 463, 544, 435]. normal-weight [463]. Norman [276, 278]. Northern [10, 120]. note [192]. notices [741]. novel [335]. November [306]. nowcasting [476]. number [830]. numbers [637]. Numerical [856, 491]. nurse [382]. nurse-reported [382]. nurses [578]. nursing [681].

obese [463]. obesity [295]. Obituaries [104, 16, 46, 160, 194, 134, 264, 276, 301, 357, 367, 386, 411, 440, 533, 509, 488]. Obituary [552, 554, 551, 553, 585, 586, 583, 601, 582, 584, 760, 759, 653, 624, 652, 708, 709, 731, 245, 620, 628, 625, 225, 651, 466, 622, 421, 730, 623, 707, 705, 621, 626, 706, 627, 682, 741]. objective [638]. observability [599]. Observational [146, 465, 524, 681, 217]. observations [326, 325]. Occupation [675]. off [384]. Official [113, 717]. officiating [464]. old [567]. older [337, 475]. Olley [527]. on-line [221]. one [224, 821]. onset [185, 758, 701]. open [356]. operation [745]. operational [791]. opinion [659, 427, 698]. opinions [256]. opportunities [823, 792]. opportunity [130]. Optimal [391, 266, 152]. Option [606]. O’Quigley [27]. order [663]. ordered [461, 456, 351]. ordinal [387]. origin [794, 439]. origin-destination [794]. originators [123]. origins [123]. Orleans [219]. orthography [331]. other [42, 439]. our [772]. out-of-school [350]. outbreak [436]. outbreaks [234]. outcome [456, 252]. outcomes [447, 514, 793, 154, 596, 42, 700, 385, 716, 356, 578, 598, 414]. outdoor [678]. output [749, 120, 41]. overflow [155]. Overlapping [3]. overseas [239]. Owen [329]. Oxford [284].

P [30, 146, 171, 84, 285, 113, 114, 283, 142, 173, 289, 623, 26, 164, 133, 32, 109]. P. [134]. P.-J [173]. Pace [174]. Pact [36]. Page [530]. Paige [264, 265]. pain [257, 415]. Pakes [527]. Pan [2]. Pan-European [2]. pandemic [297]. Panel [559, 861, 403, 754, 745, 520, 832, 244, 756, 431, 187, 836]. panels [395, 528]. Panik [87]. Paper [319, 251, 176, 317, 267, 249]. papers [540, 815, 690, 639]. papilloma [479]. paradata [322, 745, 328, 220, 315]. paradigm [250]. parameter [665, 787]. Parametric [291, 818, 719, 772, 544, 692, 377, 642, 376, 309]. Parental [187, 414]. Parsimonious [663]. partial [403, 599]. Partially [452, 693]. participant [748]. Participating [754]. participation [66, 664, 121, 338, 647]. partitioning [73]. partnership [68]. partnerships [447]. party [187]. pass [597, 747]. pass-through [597]. passenger [794]. passing [516]. passive [131]. Path [40]. patient [456, 382, 254, 751]. patient-reported [456]. pattern [616]. Patterned [779]. Patterns [143, 310, 525, 449, 100, 101, 546]. Patterson [60]. Paulson [29]. pay [361, 541, 271]. payment [593, 820]. payments [262]. Pearce [276, 275]. Pearl [145]. Pearson [65]. pedestrian [364]. peer [817, 773]. peer-to-peer [817]. penalized [691, 642]. penalty [427]. People [470, 320, 573, 463, 868, 187]. perform [394, 294, 420]. performance [121, 673, 785, 480, 125, 270]. Perils [510]. perinatal [155]. Period [781, 428]. periodic [790]. persistence [343]. Persistent [517, 479]. personality [418]. perspective [448, 501, 52]. Perspectives [202, 354, 333]. Peter [276, 277, 49, 264, 551, 601, 263, 77]. Petroleum [173]. Phantoms [545]. Pharmaceutical [228]. Pharmacology [60]. phase [484]. Philip [18]. Phillips [653]. Philosophical [492, 865]. phone [645]. Physical [205, 206]. physician [751]. Pickle [310]. Pigeonholes [814]. Pike [731]. PISA [296]. pitfall [769]. Plackett [19]. Plain [210]. planning [155, 346, 259]. players [222]. plays [132]. Point [665, 346, 724, 790]. Poisson [434, 336, 439]. Poland [722, 574]. Police [507, 531]. Policy [492, 787, 865]. Political [522, 772, 530, 659]. Politics [846, 849]. poll [92]. Polling [767, 349, 766]. polls [659, 698]. pollutants [404]. Pollution [796, 678, 450]. polynomials [429]. polytomous [485]. pooled [748]. poor [591]. poorest [591]. Population [2, 605, 600, 256, 545, 831, 475, 465, 481, 432]. populations [483]. Portfolios [61]. post [356]. post-treatment [356]. potentials [334, 510]. Poverty [643, 746, 644, 517, 722, 216]. power [568, 766, 733]. Prabakhar [582]. Practical [781, 834, 285, 712, 608, 539, 224, 144]. Practice [282, 312, 33, 607, 633]. predict [220]. Predicting [121, 353, 261, 500]. Prediction [456, 816, 13, 270, 12, 407, 610, 83]. predictions [259]. predictive [212, 818, 577, 344, 766]. Predictors [7, 578]. predicts [344]. Preface [540, 815, 690, 639]. preference [222, 770]. pregnancy [701]. Premier [434]. preschool [592]. prescriptions [751]. presence [477, 459, 692, 69, 825, 39]. presidential [767]. prevalence [403, 129, 451, 381, 122]. prevention [346]. price [445, 697, 68]. prices [379, 597, 697]. Pricing [61]. Primer [163]. Principles [33, 80]. prior [381, 482]. Private [780, 446, 691, 531]. Private-sector [780]. Probabilistic [41, 242, 270, 406, 766]. probabilities [348, 766]. Probability [807, 166, 537, 776, 727, 854, 269, 266, 435, 816, 253, 791, 549, 22, 85, 537, 399]. probit [599, 456]. Problem [781, 298]. Problems [228, 550, 785, 383]. procedure [334]. Procedures [56, 291]. Process [811, 724, 790]. processes [185, 807]. product [269, 747, 751]. production [717]. productivity [527]. Professor [133, 45]. Profile [457, 79, 212, 185, 295]. profiles [788]. Profit [61]. Program [647]. Programme [721, 296, 396, 259]. Programmers [735]. programmes [260]. Progress [113, 211, 73]. Promote [632]. promotion [618]. promptly [358]. propensity [260, 726, 153, 430]. properly [358]. properties [326]. property [379, 775]. Proportion [10]. Proportional [27]. prospective [541, 234]. protected [573]. protecting [482]. proxy [42]. psychiatry [70]. Psychometric [685]. Public [599, 693, 427, 333]. publication [212, 69]. published [342]. Publishing [246]. pupils [721].

Qualitative [151]. Quality [231, 542, 371, 127, 321, 589, 404, 67, 720, 325, 666]. quantify [505]. Quantifying [131, 364, 250]. Quantile [243, 619, 722, 581, 503, 513]. Quantitative [171, 423]. quarter [239]. quasi [544]. quasi-Monte-Carlo [544]. question [43]. questioning [244]. questionnaire [513]. questionnaires [776]. questions [380, 180]. Quiñonero [28]. Quiñonero-Candela [28].

R [205, 203, 301, 23, 146, 287, 175, 560, 33, 206, 52, 57, 491, 290, 163, 200, 229, 83, 85, 227, 303, 79, 109, 168, 174, 856, 873, 858, 863, 864, 866, 89, 31, 165, 283, 289, 116, 169, 26, 736, 115, 734, 468, 538, 561, 401, 493, 557]. racial [615]. Random [96, 779, 281, 807, 536, 403, 756, 825, 718, 513, 335]. random-effect [756]. Random-effects [96, 513]. Randomised [23]. Randomization [596, 684, 729]. Randomization-based [596]. Randomized [608, 727, 854, 71, 393, 723, 450, 268, 153]. Ranking [485, 380]. rankings [347]. Rao [730]. Rare [13, 816, 12]. Rasbash [104, 107]. rate [523, 597, 618, 753, 516, 775]. rates [665, 777, 241, 499, 660, 381, 435, 255, 570, 645, 753]. rating [663, 818, 233]. Ratings [347, 453]. ratios [483]. Rattanasiri [79]. Re [395, 667, 674]. re-election [674]. Re-engaging [395]. re-evaluation [667]. Reaction [420]. real [271]. reality [762]. really [828]. Realtime [476, 749, 407]. Reasoning [311, 805]. receipt [244]. recessionary [749]. recidivism [344, 610]. reciprocal [520]. recommendations [702]. Reconsidering [550]. reconstruction [483]. record [831, 327, 529, 728]. recorded [385]. recovery [219]. recreational [508]. redesigns [511]. Redfern [18]. redistribution [349]. redrawn [349]. reduce [341, 321]. reducing [450]. reduction [511]. Reese [111]. refereeing [44]. Referees [495, 563, 654, 740, 851]. referenced [451]. reflections [273, 680]. refreshment [593]. regimes [662]. regional [128, 188, 381]. regions [541]. register [363, 835, 237]. Regression [30, 199, 494, 59, 738, 812, 13, 88, 115, 619, 126, 375, 176, 793, 611, 747, 362, 544, 437, 543, 832, 531, 643, 272, 680, 12, 503, 41, 513, 871, 87]. regressions [429, 268, 752, 27]. regressive [499, 787, 454, 512, 794, 659, 549]. regulation [233]. related [457, 130, 541, 381, 183]. relating [405]. Relational [110]. relationship [382, 480, 477, 520, 414]. relative [530, 236, 795]. Relevant [782]. Reliability [111, 850]. reliable [43]. reliance [346]. remaining [365]. remittances [836]. renal [154]. repayment [661]. Repeated [3, 506, 524, 505]. Report [366, 419, 370, 99, 223, 299]. reported [456, 382, 385]. reporting [182, 218, 436, 183]. reports [447, 518]. representative [127, 323]. representativeness [668]. repurchases [72]. requests [751]. resampling [572]. Research [713, 86, 633, 246, 315, 176, 125, 636, 473]. Researchers [33, 580]. Reservation [66]. Residential [67, 120]. residual [377]. Resource [177, 178, 176]. Resources [173]. respond [721]. Respondent [570, 754, 438, 458, 240]. Respondent-driven [570, 438, 458, 240]. respondents [745, 395]. Response [809, 539, 630, 485, 14, 322, 328, 176, 826, 42, 726, 323, 296, 321, 184, 351, 825, 102, 570, 526, 718, 590, 430, 380, 666, 180]. responses [397, 380]. responsive [575]. responsiveness [248]. result [420, 342]. results [394, 519, 434, 720, 153, 373]. Retail [593, 820, 445]. retentivity [96]. Rethinking [560, 864]. Retirement [745, 548]. retrospective [408]. Returns [238]. Review [205, 27, 140, 143, 203, 311, 711, 806, 781, 28, 172, 87, 208, 534, 23, 30, 89, 146, 171, 210, 312, 846, 286, 782, 31, 84, 165, 291, 197, 287, 734, 535, 687, 292, 811, 175, 199, 285, 471, 779, 110, 147, 281, 198, 81, 470, 807, 113, 114, 442, 33, 59, 206, 685, 809, 850, 60, 313, 444, 472, 713, 733, 780, 805, 810, 849, 29, 52, 57, 80, 283, 468, 538, 24, 141, 142, 282, 290, 412, 536, 712, 738, 173, 82, 163, 200, 229, 289, 812, 845, 201, 309, 735]. Review [54, 86, 51, 684, 83, 844, 85, 227, 116, 169, 228, 26, 539, 443, 739, 55, 111, 202, 686, 736, 808, 848, 58, 21, 61, 230, 231, 79, 112, 164, 166, 167, 280, 288, 469, 537, 847, 737, 88, 25, 56, 144, 32, 22, 109, 168, 170, 204, 284, 115, 174, 209, 310, 145, 53, 714, 405, 321, 399, 515, 575, 234]. Reviews [441, 207, 20, 50, 78, 108, 139, 162, 196, 226, 279, 308, 398, 422, 489, 467, 555, 602, 631, 683, 710, 732, 843, 778, 804, 490, 559, 603, 607, 423, 562, 606, 494, 148, 560, 604, 605, 491, 424, 632, 608, 561, 401, 492, 493, 556, 558, 634, 557, 400, 633]. revisiting [342]. revolving [822]. reweighting [215]. rhetoric [772]. Rice [249]. Richard [225, 840, 626]. Rigby [54]. right [92, 154, 751, 187]. right-wing [187]. Risk [603, 823, 502, 394, 478, 750, 242, 37, 815, 420, 252, 487, 408, 479, 700, 669, 693, 798, 549, 217, 819, 711]. risks [38, 579, 723, 728]. risky [346]. road [830, 571]. Robert [194, 289, 193, 195, 19]. Roberts [201]. Robone [249]. Robust [644, 142]. Rod [301, 305]. Rodney [707]. Roger [104, 103, 245]. Rogerson [114]. Role [492, 746, 529, 40, 865]. Roma [592]. Romer [583]. Ronchetti [142]. Rosenbaum [146]. rotating [506, 528]. Royal [340, 273]. Rozovskii [284]. RPIJ [445]. Rule [472, 869]. rules [391]. run [746, 548, 525]. runs [158]. Ryan [30].

S [286, 113, 60, 29, 57, 283, 55, 111, 202, 79, 167, 280, 56, 168, 115, 667]. S. [586]. safeguard [237]. Safety [58, 382]. Saharan [407]. sake [68]. sale [190]. Sample [268, 776, 784, 721, 378, 465, 615, 588, 184, 825, 505]. samples [593, 131, 296]. Sampling [231, 727, 854, 429, 438, 458, 240, 391, 570, 718]. Samuel [134, 137]. Sarkar [31]. SAS [87, 55, 287]. Satisfaction [323, 576]. save [340]. Saveliev [116]. scale [595, 180]. Scaling [8, 457]. Scarring [365]. scenario [788]. scheduling [152]. Schilling [231]. schizophrenia [252]. School [7, 350, 599, 365, 192, 679, 581, 73]. school-leavers [365]. schoolchildren [296]. schools [721, 449, 824, 333]. Schumacker [290]. Schwaighofer [28]. Science [856, 873, 687, 468, 491, 51, 85, 492, 474, 347, 865, 605]. Sciences [424, 712, 163, 286]. Scientific [113, 144, 485]. Scientists [844, 227, 205, 206]. score [260]. scored [158]. scores [153, 513]. scoring [817, 390, 820]. Scotland [797, 297]. scraped [697]. screening [378, 190, 479, 233, 259]. Seasonal [613, 786, 617]. second [719]. second-leg [719]. secondary [449]. sectional [354, 150, 339, 701]. sector [780]. security [36, 535]. seeds [814]. segregation [449, 120]. select [212]. Selected [855, 556, 97, 510]. Selection [481, 222, 480, 615, 691, 387, 530, 184]. self [510]. self-selected [510]. semi [333]. semi-isolated [333]. Semiparametric [514, 824, 544, 669, 786, 356, 221, 752, 725, 790]. Semiseparable [175]. Senate [674]. Senegal [645]. sensitive [573, 181]. Sensitivity [702, 774, 240]. Separating [669, 594]. Sequence [529, 515]. Sequences [807, 8]. sequencing [216]. Sequential [870, 185, 391, 752, 471]. Series [165, 283, 24, 511, 650, 588, 452, 374, 482, 406, 109, 604]. Serious [735, 830]. service [541, 590, 176]. set [668]. sets [747, 728]. setting [543]. settings [529]. several [342]. severity [571]. sex [483, 546]. sexual [447]. shape [642]. shape-constrained [642]. Shaped [846]. share [72]. Sharing [633, 521]. Sheather [115]. Shephard [282]. Sheskin [291]. Shift [28]. shifts [243]. ships [791]. shocks [789, 68]. Short [548, 372]. Short- [548]. short-term [372]. Shortages [5]. Should [4, 588]. sibship [480]. signal [151]. Silverstein [281]. Similar [13, 12]. similarities [792]. simple [506]. Simpson [837]. Simulating [865, 492]. simulation [492, 865, 594, 168]. single [481, 43, 404]. single- [481]. Singular [175, 353]. sites [221]. size [550, 334, 831, 480, 410, 268]. sized [819]. Skerrett [251]. skew [641, 435]. skewed [640]. skills [417]. slavery [214]. Slope [154]. Small [641, 455, 640, 718, 215, 639, 505, 528, 646, 829, 346, 650, 514, 500, 648, 122, 647, 643, 486, 484, 649, 717, 819, 642, 725]. Smeeton [32]. Smith [116, 249]. smoking [130, 131, 385]. smoking-related [130]. Smoothing [696, 270, 693, 484]. soccer [44]. Social [350, 687, 81, 113, 33, 362, 713, 712, 51, 254, 343, 616, 315, 40, 668, 570, 473, 392]. socialization [522]. societies [123]. Society [113, 814, 273, 247]. Socio [836, 415, 520]. Socio-Economic [836, 415, 520]. sociodemographic [645]. Sociology [605]. Solutions [400]. solving [695]. some [211]. Somerville [264, 265]. sources [755]. space [511, 650, 571, 528]. span [834]. Spanish [600, 457]. spanking [258]. sparse [383]. Spatial [172, 349, 236, 590, 174, 599, 364, 459, 336, 404, 757, 699]. Spatially [508, 502, 451]. Spatiotemporal [660, 794, 448, 762, 797, 678, 452, 407, 790]. Species [242]. specific [435, 255, 753, 694]. specification [446, 376]. Spectral [281]. spectrum [353, 588]. speech [324]. speed [41]. spending [679]. Spillovers [787]. splines [642]. split [378]. split-sample [378]. sports [379]. Sprent [32, 601]. SPSS [424]. squared [644]. Squares [199]. stability [547]. stadia [379]. stadium [507]. stage [334, 185, 152]. Staines [285]. Stamhuis [113]. Stan [864, 560]. stand [247]. Stanley [276, 275, 706]. start [717]. state [511, 650, 768, 757, 427, 796, 528, 86]. Statistical [612, 84, 291, 287, 734, 273, 110, 113, 689, 858, 59, 444, 52, 424, 608, 83, 864, 355, 716, 634, 55, 167, 25, 233, 32, 234, 498, 872, 690, 568, 636, 123, 816, 645, 333, 344, 65, 149, 64, 292, 811, 114, 560, 33, 738, 173, 557, 209]. statistician [132]. Statisticians [227]. Statistics [205, 207, 506, 855, 846, 474, 113, 206, 60, 54, 86, 561, 169, 26, 556, 866, 359, 736, 848, 867, 868, 537, 22, 399, 567, 641, 638, 772, 402, 655, 41, 717, 498, 81, 142, 228, 210, 562, 470, 845, 443, 537, 737]. status [403, 520]. Staying [68]. Steele [51]. Stein [620]. Stella [301, 300]. Stephen [622]. Stephens [838]. stereoscopic [762]. Stick [230]. Stochastic [208, 580, 256, 606, 807, 202, 168, 204, 476, 789, 833, 217, 170]. store [699]. Stories [143, 846]. strategic [519]. strategy [158]. stratification [350]. stratified [255]. strengths [513]. stress [578]. stroke [39]. stroke-free [39]. Structural [290, 186, 454, 512, 692, 578]. Structure [6, 1, 616, 570, 406]. structures [377]. Student [721, 296, 673, 716, 679]. studies [346, 378, 465, 524, 510, 404, 484, 146, 57]. Study [492, 130, 508, 318, 681, 244, 385, 69, 865, 189, 95, 405, 517, 513]. sub [407]. sub-Saharan [407]. Subba [730]. subgroups [577]. Subject [741]. Subjective [504, 638, 520, 766, 180]. subnational [757]. Subtypes [7]. success [324]. sudden [373]. sugar [214]. Sugiyama [28]. suicide [255]. summarization [257]. superior [681]. supermarket [699]. supervised [452]. supply [550]. support [588]. surface [432]. surgical [681]. surrogate [346]. Surveillance [114, 502, 568, 459, 297, 233]. Survey [518, 846, 360, 325, 828, 576, 754, 241, 511, 218, 745, 784, 595, 324, 315, 500, 769, 42, 323, 244, 756, 640, 825, 671, 526, 720, 575, 594, 395, 666, 650, 745, 647, 574]. survey-based [769]. Surveys [3, 518, 506, 322, 328, 721, 150, 341, 451, 475, 826, 510, 481, 42, 151, 122, 296, 486, 668, 792, 430, 335, 505, 213, 591]. Survival [29, 38, 154, 408, 236, 598, 699]. Surviving [214]. Sutradhar [313]. Swiss [255]. syndrome [373]. Syndromic [297]. synoptic [298]. Synthesis [385]. synthetic [692, 693, 482, 694]. system [541, 831]. Systematic [780]. Systems [2, 818, 248, 208, 175, 204].

T [30, 171, 210, 76, 17, 81, 83, 227, 109, 209].

T. [584]. table [191]. tables [98, 192, 696]. Tail [711, 618]. tail-free [618]. tailed [544]. Tailored [323]. tails [597]. Takács [554]. taking [478]. Tallis [628]. Tamhane [228]. Tango [209]. targeted [826]. targeting [475]. tariffs [541]. Tarling [33]. Taskar [110]. Tata [730]. Taylor [568]. Teacher [5]. teaching [716, 737]. team [453]. teams [334]. Techniques [59, 9]. technologies [391]. technology [486]. teen [660]. Tell [686]. temperatures [613]. temporal [261]. temporally [502]. tennis [827]. tenure [238]. Terence [625]. Term [2, 365, 463, 678, 372]. terms [397]. terrorism [700, 757]. terrorist [612]. test [190, 158]. Testing [471, 228, 56, 870, 406, 376, 129, 424]. Thatcher [104, 103]. Their [4, 5, 492, 865]. Theory [711, 87, 80, 22, 774, 809, 400, 714]. therapy [356]. there [384]. thieves [507]. Think [470, 868]. Thomas [842, 292, 623, 61]. Those [735]. three [747, 722, 668, 395]. three-level [722]. three-pass [747]. thresholds [351]. Tibshirani [83]. Time [453, 37, 604, 283, 24, 109, 293, 665, 650, 571, 750, 787, 327, 454, 477, 618, 455, 588, 701, 374, 255, 482, 95, 699, 406, 165]. times [328, 220, 342, 385, 590]. timing [834]. Timothy [625]. Tingley [317]. Todman [285]. together [68]. Tom [104, 106]. Tool [8]. Tools [712]. Topics [855, 556, 443, 399]. topological [616]. total [39]. trachomatis [217]. Tracking [697]. trade [384]. trade-off [384]. traditional [817]. traffic [571]. training [677, 260]. traits [418]. trajectories [389, 515]. transaction [689]. transactions [819]. Transfers [10]. Transition [7]. transitions [663]. transmission [217]. transmitting [459]. transplant [154]. transportation [439]. Travel [650]. treated [396, 465]. Treatment [11, 383, 834, 71, 465, 258, 356, 702, 662]. tree [577]. tree-based [577]. trend [786]. trend-seasonal [786]. trends [239, 660, 255, 427, 676]. trial [346, 257, 252, 723, 70, 596]. Trials [608, 727, 854, 703, 393, 268, 792, 153, 23, 58]. triple [831]. triple-record [831]. Trust [637]. Turnout [9, 437]. Tutorial [472, 869]. twin [550]. two [334, 152, 484, 483]. two-phase [484]. two-sex [483]. two-stage [334, 152]. twofold [643]. Tying [11]. type [238, 830, 790]. types [795].

UK [567, 239, 384, 514, 375, 517, 771, 92, 150, 748, 668, 716, 68, 698, 574]. Ulf [553]. unbalanced [747]. Uncertainties [492, 865]. Uncertainty [536, 764, 127]. undecided [767]. under-age [268]. Understanding [872, 595, 444]. Unemployed [4, 365, 677]. unemployment [514, 187]. unintended [823]. Union [506]. unisex [823]. unit [829, 722]. united [247]. units [349]. university [272, 598]. unknown [428]. Unpacking [576]. unreliable [545]. Unrestricted [429]. unsecured [661]. unsupervised [824]. ups [321]. uranium [487]. urban [700]. Urdan [210]. USA [749, 660]. Usage [11, 593]. Use [846, 7, 13, 298, 749, 814, 393, 315, 722, 758, 693, 12, 153, 180]. user [543, 221]. user-generated [221]. users [381, 405]. Using [403, 322, 856, 94, 328, 220, 2, 9, 42, 13, 549, 777, 502, 222, 97, 830, 595, 727, 854, 593, 338, 734, 71, 327, 641, 577, 451, 125, 827, 364, 797, 154, 124, 270, 381, 260, 618, 684, 272, 659, 816, 55, 697, 693, 189, 645, 12, 79, 215, 41, 820, 217, 513, 594, 642, 775, 725, 376, 491]. utility [694, 325]. utilization [541, 446]. Uusipaikka [88].

V [205, 208, 206, 283, 231, 204]. Validating [180]. validation [818, 244]. validity [212]. Valuation [606]. Value [175, 348, 679, 581]. value-added [679]. valued [604]. Values [13, 393, 337, 12, 39]. Vandebril [175]. variability [404, 236]. variable [619, 475, 748, 726]. variables [729, 596, 40]. Variance [342, 506, 511, 189]. Variants [684]. Variation [418, 336]. variations [156]. varieties [355]. varying [665, 502, 453, 750, 37, 787, 454, 477]. Vaughan [301, 306]. vector [787, 454, 512, 659]. vehicle [830]. versus [650, 378]. via [457, 485, 456, 573, 588, 581, 752, 821]. victimization [776]. Victor [801]. Victorian [253]. Vidyadhar [582]. view [762, 475]. viewpoint [293]. vignette [181]. Vignettes [248, 179, 184, 180, 376]. virtual [762]. virus [448, 403, 479]. visits [336]. Visualization [31, 763, 765, 808]. Visualizing [762, 753, 310]. Vivian [301, 300]. vol [606, 175, 113]. Volatility [606, 476, 789]. vols [809, 556]. volume [117, 232, 314, 369, 368, 426, 425, 497, 496, 565, 657, 743, 853]. voodoo [347]. Vote [10, 349]. Voter [9, 767, 437]. voting [519]. vouchers [677]. vulnerability [700].

W [264, 584, 89, 281, 57, 263, 227, 202, 230, 310]. Wage [388, 375, 332, 243]. wages [66, 615]. Wales [213, 678, 547]. Walker [104, 116, 105, 77]. Wallenius [770]. Walter [76, 708]. Want [735]. War [402, 655, 846]. waste [487]. wavelet [617, 26]. Ways [10, 536]. weather [504]. web [697]. web-scraped [697]. weight [463, 250, 389, 189]. weighted [457]. weights [727, 854]. welfare [769, 680]. wellbeing [567, 504]. Werner [708]. where [294, 578]. whether [751]. Which [721, 344, 237]. Whitehead [709]. Who [470, 735, 868]. whose [246]. Wickham [147]. Widening [338]. Wilfred [301, 307]. Wilfrid [48]. Wilkinson [624]. William [579, 75]. Wilson [111]. win [597]. wind [41]. Windows [11]. wing [187]. within [343, 616, 388]. within-group [343]. without [822, 666]. Wolters [24]. Wood [842]. Woods [273]. work [182, 834, 152, 113]. workdays [152]. Worker [773]. workers [676, 373]. workflow [763]. workplace [345]. workweeks [152]. World [311]. worldwide [757]. Wroe [421]. Wyndham [264, 265].

Yamada [114]. Yamamoto [317]. year [828, 100, 101]. year-of-birth [100, 101]. years [375, 591]. young [748, 187]. Youth [360]. Yule [680].

Z [281]. Zealand [9]. Zero [569, 821]. zeroand-one [821]. Zero-inflated [569]. zeros [692]. Zhang [80]. Zuur [89, 116].

References

Driffield:2010:ECS

[1] Nigel Driffield and Sarmistha Pal. Evolution of capital structure in East Asia — corporate inertia or endeavours? Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):1–29, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622576.

Gregory:2010:MLT

[2] Ian N. Gregory, Jordi Marti-Henneberg, and Francisco J. Tapiador. Modelling long-term pan-European population change from 1870 to 2000 by using geographical information systems. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):31–50, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622577.

Chipperfield:2010:EER

[3] James Chipperfield and Philip Bell. Embedded experiments in repeated and overlapping surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):51–66, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/20622578.

Behncke:2010:UTC

[4] Stefanie Behncke, Markus Frölich, and Michael Lechner. Unemployed and their caseworkers: Should they be friends or foes? Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):67–92, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622579.

Bourdon:2010:TST

[5] Jean Bourdon, Markus Frölich, and Katharina Michaelowa. Teacher shortages, teacher contracts and their effect on education in Africa. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 93–116, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622580.

Alho:2010:FHS

[6] Juha Alho and Nico Keilman. On future household structure. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 117–143, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622581.

Reboussin:2010:LTM

[7] Beth A. Reboussin and Nicholas S. Ialongo. Latent transition models with latent class predictors: Attention deficit hyperactivity disorder subtypes and high school marijuana use. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 145–164, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622582.

Piccarreta:2010:ESG

[8] Raffaella Piccarreta and Orna Lior. Exploring sequences: A graphical tool based on multi-dimensional scaling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 165–184, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622583.

Hudson:2010:EIT

[9] Irene L. Hudson, Linda Moore, Eric J. Beh, and David G. Steel. Ecological inference techniques: An empirical evaluation using data describing gender and voter turnout at New Zealand elections, 1893–1919. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):185–213, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622584.

OKelly:2010:HMW

[10] Michael O’Kelly, John Doyle, and Philip J. Boland. How many ways can you look at a proportion?: Crosscommunity vote transfers in Northern Ireland after the Belfast Agreement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 215–235, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622585.

Lee:2010:MUE

[11] Myoung jae Lee. Measuring the usage effects of tying a messenger to windows: A treatment effect approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 237–253, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622586.

Seed:2010:CUC

[12] Paul Seed. Correspondence: The use of cost information when defining critical values for prediction of rare events by using logistic regression and similar methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):255–256, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Seed:2010:UCI

[13] Paul Seed, Richard Berk, Lawrence Sherman, Geoffrey Barnes, Ellen Kurtz, and Lindsay Ahlman. The use of cost information when defining critical values for prediction of rare events by using logistic regression and similar methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 255–258, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622587.

Berk:2010:CAR

[14] Richard Berk, Lawrence Sherman, Geoffrey Barnes, Ellen Kurtz, and Lindsay Ahlman. Correspondence: Authors’ response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):256–258, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lindley:2010:IJG

[15] Dennis V. Lindley. Irving John Good, 1916–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):259–261, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622588.

Anonymous:2010:Oa

[16] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 259–267, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Clarke:2010:KB

[17] Geoff Clarke and Tony Haws. K. T. Boyd, 1923–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):262, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622589.

Fox:2010:PR

[18] John Fox and Roger Thatcher. Philip Redfern, 1922–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):262–265, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/20622590.

Diggle:2010:RLP

[19] Peter Diggle, Robin Henderson, and John Matthews. Robert Lewis Plackett, 1920–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):265–267, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622591.

Anonymous:2010:BRa

[20] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):269, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rosenbaum:2010:BRB

[21] Janet Rosenbaum. Book review: Bayesian Methods for Measures of Agreement, by L. Broemeling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):270, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622592.

Stoyanov:2010:BRBa

[22] Jordan Stoyanov. Book review: Asymptotic Theory of Statistics and Probability, by A. DasGupta. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 270–271, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622593.

Batistatou:2010:BRB

[23] Eva Batistatou. Book review: Cluster Randomised Trials, by R. J. Hayes; L. H. Moulton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):271, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622594.

Kumar:2010:BRB

[24] Kuldeep Kumar. Book review: Introduction to Modern Time Series Analysis, by G. Kirchgassner; J. Wolters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 271–272, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622595.

Smart:2010:BRB

[25] Trevor Smart. Book review: The Statistical Analysis of Functional MRI Data, by N. A. Lazar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):272–273, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622596.

Paparoditis:2010:BRB

[26] Efstathios Paparoditis. Book review: Wavelet Methods in Statistics with R, by G. P. Nason. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):273, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622597.

Aalabaf-Sabaghi:2010:BRB

[27] Morteza Aalabaf-Sabaghi. Book review: Proportional Hazards Regressions, by J. O’Quigley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):273–274, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622598.

Adams:2010:BRB

[28] Niall Adams. Book review: Dataset Shift in Machine Learning, by J. Quiñonero-Candela; M. Sugiyama; A. Schwaighofer; N. D. Lawrence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):274, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622600.

Kleine:2010:BRBa

[29] Liliana López Kleine. Book review: Biostatistics and Microbiology: A Survival Manual, by D. S. Paulson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):274, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622599.

Bedogni:2010:BRBa

[30] Giorgio Bedogni. Book review: Modern Regression Methods, by T. P. Ryan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):275, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622601.

Cheshire:2010:BRBa

[31] James Cheshire. Book review: Lattice: Multivariate Data Visualization with R, by D. Sarkar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1):275–276, January 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622602.

Stoimenova:2010:BRB

[32] Eugenia Stoimenova. Book review: Applied Nonparametric Statistical Methods, by P. Sprent; N. C. Smeeton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):276, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622603.

Hochkirchen:2010:BRBa

[33] Thomas Hochkirchen. Book review: Statistical Modelling for Social Researchers: Principles and Practice, by R. Tarling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(1): 276–277, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622604.

Anonymous:2010:BMa

[34] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):??, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622605.

Anonymous:2010:FMa

[35] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):??, January 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622575.

Goddard:2010:ELS

[36] Ivor J. Goddard. Editorial: Liberty and security: the Faustian Pact. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 279–281, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666245.

Crook:2010:TVD

[37] Jonathan Crook and Tony Bellotti. Time varying and dynamic models for default risk in consumer loans. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 283–305, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666246.

Basu:2010:BCS

[38] Sanjib Basu and Ram C. Tiwari. Breast cancer survival, competing risks and mixture cure model: a Bayesian analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):307–329, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666247.

vandenHout:2010:ESF

[39] Ardo van den Hout and Fiona E. Matthews. Estimating stroke-free and total life expectancy in the presence of non-ignorable missing values. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 331–349, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666248.

Kuha:2010:PAD

[40] Jouni Kuha and John H. Goldthorpe. Path analysis for discrete variables: the role of education in social mobility. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 351–369, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666249.

Thorarinsdottir:2010:PFW

[41] Thordis L. Thorarinsdottir and Tilmann Gneiting. Probabilistic forecasts of wind speed: ensemble model output statistics by using heteroscedastic censored regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):371–388, April 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40666250.

Kreuter:2010:UPM

[42] F. Kreuter, K. Olson, J. Wagner, T. Yan, T. M. Ezzati-Rice, C. Casas-Cordero, M. Lemay, A. Peytchev, R. M. Groves, and T. E. Raghunathan. Using proxy measures and other correlates of survey outcomes to adjust for nonresponse: examples from multiple surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):389–407, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666251.

Micklewright:2010:HRI

[43] John Micklewright and Sylke V. Schnepf. How reliable are income data collected with a single question? Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 409–429, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666252.

Buraimo:2010:MRB

[44] Babatunde Buraimo, David Forrest, and Robert Simmons. The 12th man?: refereeing bias in English and German soccer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):431–449, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666253.

Tong:2010:PHA

[45] Howell Tong. Professor Hirotugu Akaike, 1927–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):451–454, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666254.

Anonymous:2010:Ob

[46] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):

451–460, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fieller:2010:HMD

[47] Nick Fieller. Hilda M. Davies, 1914– 2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):454–455, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666255.

Flournoy:2010:WJD

[48] Nancy Flournoy. Wilfrid Joseph Dixon, 1915–2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):455–457, April 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40666256.

Aitken:2010:PF

[49] Colin Aitken and David J. Finney. Peter Fisk, 1929–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):458–460, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666257.

Anonymous:2010:BRb

[50] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):461, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Liao:2010:BRB

[51] Tim Futing Liao. Book review: Analysis of Multivariate Social Science Data, by D. J. Bartholomew; F. Steele; I.

Moustaki; J. I. Galbraith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):462, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666258.

Kleine:2010:BRBb

[52] Liliana López Kleine. Book review: Computational Biology: a Statistical Mechanics Perspective, by R. Blossey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):462–463, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666259.

Wade:2010:BRB

[53] A. Wade. Book review: Bayesian Biostatistics and Diagnostic Medicine, by L. D. Broemeling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):463, April 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40666260.

Lewis:2010:BRBa

[54] Daniel Lewis. Book review: Elementary Statistics for Geographers, 3rd edn, by J. E. Burt; G. M. Barber; D. L. Rigby. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):464, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666261.

Pinon:2010:BRB

[55] Antonio Pinon. Book review: A Handbook of Statistical Analyses using SAS, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2): 464–465, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666262.

Sperrin:2010:BRB

[56] Matthew Sperrin. Book review: Multiple Testing Procedures with Applications to Genomics, by S. Dudoit; M. J. van der Laan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):465, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666263.

Kleine:2010:BRBc

[57] Liliana López Kleine. Book review: Bioconductor Case Studies, by F. Hahne; W. Huber; R. Gentleman; S. Falcon. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):465–466, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666264.

Rehman:2010:BRB

[58] Andrea M. Rehman. Book review: Data and Safety Monitoring Committees in Clinical Trials, by J. Herson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):466–467, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666265.

Hochkirchen:2010:BRBb

[59] Thomas Hochkirchen. Book review: Modern Multivariate Statistical Techniques: Regression, Classification, and Manifold Learning, by A. J. Izenman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):467, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666266.

Jaki:2010:BRB

[60] Thomas Jaki. Book review: Bioequivalence and Statistics in Clinical Pharmacology, by S. Patterson; B. Jones. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):467–468, April 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666267.

Savvopoulos:2010:BRB

[61] Anastasios Savvopoulos. Book review: Consumer Credit Models: Pricing, Profit and Portfolios, by L. C. Thomas. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):468, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666268.

Anonymous:2010:BMb

[62] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):??, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        40666269.

Anonymous:2010:FMb

[63] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(2):??, April 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666244.

Stigler:2010:DGS

[64] Stephen M. Stigler. Darwin, Galton and the Statistical Enlightenment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (3):469–482, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666271.

Turner:2010:CIS

[65] Elizabeth L. Turner and James A. Hanley. Cultural imagery and statistical models of the force of mortality: Addison, Gompertz and Pearson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 483–500, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666272.

Brown:2010:RWL

[66] Sarah Brown, Jennifer Roberts, and Karl Taylor. Reservation wages, labour market participation and health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 501–529, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666273.

Rabe:2010:RMQ

[67] Birgitta Rabe and Mark Taylor. Residential mobility, quality of neighbourhood and life course events. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 531–555, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666274.

Rainer:2010:STS

[68] Helmut Rainer and Ian Smith. Staying together for the sake of the home?: house price shocks and partnership dissolution in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):557–574, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666275.

Peters:2010:APB

[69] Jaime L. Peters, Alex J. Sutton, David R. Jones, Keith R. Abrams, Lesley Rushton, and Santiago G. Moreno. Assessing publication bias in metaanalyses in the presence of betweenstudy heterogeneity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):575–591, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/40666276.

Jackson:2010:HMC

[70] Dan Jackson, Ian R. White, and Morven Leese. How much can we learn about missing data?: an exploration of a clinical trial in psychiatry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 593–612, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666277.

Dias:2010:EAB

[71] S. Dias, N. J. Welton, V. C. C. Marinho, G. Salanti, J. P. T. Higgins, and A. E. Ades. Estimation and adjustment of bias in randomized evidence by using mixed treatment comparison metaanalysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):613–629, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40666278.

Bhargava:2010:EAD

[72] Alok Bhargava. An econometric analysis of dividends and share repurchases by US firms. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):631–656, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40666279.

Rasbash:2010:CEP

[73] Jon Rasbash, George Leckie, Rebecca Pillinger, and Jennifer Jenkins. Children’s educational progress: partitioning family, school and area effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (3):657–682, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666280.

Brillinger:2010:ELL

[74] David R. Brillinger. Erich Leo Lehmann, 1917–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):683–686, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/40666281.

Rao:2010:WMI

[75] P. V. Rao. William Mendenhall III, 1925–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):686–687, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/40666282.

Casella:2010:WF

[76] George Casella, Chuck McCulloch, Steve Schwager, and Marty Wells. Walter T. Federer, 1915–2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):687–688, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40666283.

Ross:2010:PW

[77] Gavin Ross and Peter Thursfield. Peter Walker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):688–689, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/40666284.

Anonymous:2010:BRc

[78] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):690, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2010:BRBa

[79] Shalabh. Book review: Meta-analysis of Binary Data using Profile Likelihood, by D. Böhning; R. Kuhnert; S. Rattanasiri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):691, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666285.

Kleine:2010:BRBd

[80] Liliana López Kleine. Book review: Principles and Theory for Data Mining and Machine Learning, by B. Clarke; E. Fokoué; H. H. Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 691–692, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666286.

Gonzalez-Izquierdo:2010:BRB

[81] Arturo González-Izquierdo. Book review: Introduction to Social Statistics, by T. Dietz; L. Kalof. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):692, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40666287.

Lazic:2010:BRB

[82] S. E. Lazic and Roche Basel. Book review: Introduction to Design of Experiments with JMP Examples, 3rd edn, by J. Goupy; L. Creighton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):

692–693, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666288.

Lu:2010:BRB

[83] Z. Q. John Lu. Book review: The Elements of Statistical Learning: Data Mining, Inference, and Prediction, 2nd edn, by T. Hastie; R. Tibshirani; J. Friedman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):693–694, July 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/40666289.

Cheshire:2010:BRBb

[84] James Cheshire. Book review: A First Course in Bayesian Statistical Methods, by P. D. Hoff. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):694–695, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        40666290.

McElduff:2010:BRB

[85] Fiona McElduff. Book review: Probability with R: an Introduction with Computer Science Applications, by J. M. Horgan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):695, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40666291.

Lewis:2010:BRBb

[86] Daniel Lewis. Book review: State and Local Government Statistics at a Crossroads National Research Council.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (3):695–696, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666292.

Andersen:2010:BRB

[87] Per Kragh Andersen. Book review: Regression Modeling: Methods, Theory, and Computation with SAS, by M. Panik. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):696, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666293.

Skovgaard:2010:BRB

[88] L. T. Skovgaard. Book review: Confidence Intervals in Generalized Regression Models, by E. Uusipaikka. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3): 696–697, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666294.

Bedogni:2010:BRBb

[89] Giorgio Bedogni. Book review: A Beginner’s Guide to R, by A. F. Zuur; E. N. Ieno; E. H. W. G. Meesters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (3):697–698, July 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666295.

Anonymous:2010:BMc

[90] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):??, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        40666296.

Anonymous:2010:FMc

[91] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):??, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        40666270.

Fisher:2010:EGI

[92] Stephen D. Fisher, Jouni Kuha, and Clive Payne. Editorial: Getting it right on the night, again — the 2010 UK general election exit poll. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 699–701, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40925394.

Raymer:2010:EMM

[93] James Raymer and Peter W. F. Smith. Editorial: Modelling migration flows. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 703–705, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40925395.

Boden:2010:UAD

[94] Peter Boden and Phil Rees. Using administrative data to improve the estimation of immigration to local areas in England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):707–731, October 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40925396.

Smith:2010:CAM

[95] Peter W. F. Smith, James Raymer, and Corrado Giulietti. Combining available migration data in England to study economic activity flows over time. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 733–753, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925397.

Congdon:2010:REM

[96] Peter Congdon. Random-effects models for migration attractivity and retentivity: a Bayesian methodology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 755–774, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925398.

Bijak:2010:BFI

[97] Jakub Bijak and Arkadiusz Wiśniowski. Bayesian forecasting of immigration to selected European countries by using expert knowledge. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):775–796, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40925399.

Abel:2010:EIM

[98] Guy J. Abel. Estimation of international migration flow tables in Europe.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 797–825, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925400.

Hand:2010:RC

[99] David J. Hand, Martin Dougherty, Sarah Simpson, Gerald Goodall, Trevor Lewis, Sheila Bird, Andrew Garratt, Jill Leyland, Colin Aitken, Ivor Goddard, Ron Smith, Paul Gentry, Adrian Bowman, Nicola Emmerson, Andy Garrett, John McColl, Penny Bidgood, Debra Hurcomb, Martin Owen, Robin Henderson, Frank Duckworth, Julian Champkin, Mary Sweetland, Jenny Church, Liam Murray, Heiko Grossmann, Richard Chandler, Vic Barnett, Julian Stander, Stephen Walters, Liz Harris, G. Robin Henderson, Richard Samworth, Antony Fielding, Moira Mugglestone, Jenny Lannon, David Martin, Maurice Merchant, Daphne Kounali, Chris Brignell, Mairi Spowage, Ben Torsney, Claus Dieter Mayer, Ting-Li Su, Paul Baxter, John Logsdon, Tony Hildreth, Jemma Hopewell, Fazil Baksh, Sara Hilditch, Paul Smith, John Eales, Heather Turner, Tom Grimes, David Hand, and Sayer Vincent. Report of the Council for 2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 827–914, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925401.

Murphy:2010:CDY

[100] Michael Murphy. Correspondence: Detecting year-of-birth mortality patterns with limited data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):915–920, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Murphy:2010:DYB

[101] Michael Murphy and Stephen Richards.

Detecting year-of-birth mortality patterns with limited data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 915–924, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925402.

Richards:2010:CAR

[102] Stephen Richards. Correspondence: Author’s response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):920–924, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Craig:2010:ART

[103] John Craig and John Fox. Arthur Roger Thatcher, 1926–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 925–927, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925403.

Anonymous:2010:OAR

[104] Anonymous. Obituaries: Arthur Roger Thatcher; Morris Walker; Tom Griffin; Jon Rasbash. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):925–932, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Priestley:2010:MW

[105] Maurice Priestley and Nick Fieller. Morris Walker, 1921–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 928–929, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40925404.

Wroe:2010:TG

[106] David Wroe. Tom Griffin, 1937–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 929–931, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40925405.

Goldstein:2010:JR

[107] Harvey Goldstein, Fiona Steele, and William Browne. Jon Rasbash. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 931–932, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/40925406.

Anonymous:2010:BRd

[108] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):933, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stoyanov:2010:BRBb

[109] Jordan Stoyanov. Book review: Handbook of Financial Time Series, by T. G. Andersen; R. A. Davis; J.-P. Kreiss; T. Mikosch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):934, October 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/40925407.

Ginestet:2010:BRB

[110] Cedric Ginestet. Book review: Introduction to Statistical Relational Learning, by L. Getoor; B. Taskar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 934–935, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925408.

Ravi:2010:BRB

[111] Sreenivasan Ravi. Book review:

Bayesian Reliability, by M. S. Hamada; A. G. Wilson; C. S. Reese; H. F. Martz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (4):935, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925409.

Shalabh:2010:BRBb

[112] Shalabh. Book review: Linear Model Methodology, by A. I. Khuri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 935–936, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925410.

Hand:2010:BRBa

[113] David J. Hand. Book review: The Statistical Mind in Modern Society: The Netherlands 1850–1940: vol. 1, Official Statistics, Social Progress and Modern Enterprise; vol. 2, Statistics and Scientific Work, by P. M. M. Klep; J. G. S. J. van Maarseveen; I. H. Stamhuis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 936–937, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925411.

Hand:2010:BRBb

[114] David J. Hand. Book review: Statistical Detection and Surveillance of Geographic Clusters, by P. Rogerson; I. Yamada. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):937–938, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925412.

Ugarte:2010:BRB

[115] M. Dolores Ugarte. Book review: A Modern Approach to Regression with R, by S. J. Sheather. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):938, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40925413.

Militino:2010:BRB

[116] Ana F. Militino. Book review: Mixed Effects Models and Extensions in Ecology with R, by A. F. Zuur; E. N. Ieno; N. J. Walker; A. A. Saveliev; G. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 938–939, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40925414.

Anonymous:2010:IAI

[117] Anonymous. Index of authors: Index of authors, volume 173, 2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4): 941–945, October 2010. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2010:BMd

[118] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):??, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40925415.

Anonymous:2010:FMd

[119] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(4):??, October 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
40925393.

Shuttleworth:2011:EIC

[120] I. G. Shuttleworth, C. D. Lloyd, and D. J. Martin. Exploring the implications of changing census output geographies for the measurement of residential segregation: the example of Northern Ireland 1991–2001. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):1–16, January 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013389.

Brunsdon:2011:PPH

[121] Chris Brunsdon, Paul Longley, Alex Singleton, and David Ashby. Predicting participation in higher education: a comparative evaluation of the performance of geodemographic classifications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):17–30, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013390.

Manzi:2011:MBC

[122] Giancarlo Manzi, David J. Spiegelhalter, Rebecca M. Turner, Julian Flowers, and Simon G. Thompson. Modelling bias in combining small area prevalence estimates from multiple surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 31–50, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013391.

OBrien:2011:OOE

[123] Christopher O’Brien. The origins and originators of early statistical societies: a comparison of Liverpool and Manchester. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):51–62, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013392.

Jenkins:2011:MIU

[124] Stephen P. Jenkins, Richard V. Burkhauser, Shuaizhang Feng, and Jeff Larrimore. Measuring inequality using censored data: a multiple-imputation approach to estimation and inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (1):63–81, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013393.

Goldstein:2011:ERP

[125] Harvey Goldstein. Estimating research performance by using research grant award gradings. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):83–93, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013394.

Costa-Font:2011:BHL

[126] Joan Costa-Font, Marin Gemmill, and Gloria Rubert. Biases in the healthcare luxury good hypothesis?: a metaregression analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):95–107, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013395.

Lee:2011:CRA

[127] Duncan Lee, Claire Ferguson, and E. Marian Scott. Constructing representative air quality indicators with measures of uncertainty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):109– 126, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013396.

Andrews:2011:DRM

[128] Martyn Andrews, Ken Clark, and William Whittaker. The determinants of regional migration in Great Britain: a duration approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):127–153, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013397.

Faes:2011:EHP

[129] Christel Faes, Marc Aerts, Saskia Litière, Estelle Méroc, Yves Van der Stede, and Koen Mintiens. Estimating herd prevalence on the basis of aggregate testing of animals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 155–174, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013398.

Balia:2011:CHS

[130] Silvia Balia and Andrew M. Jones. Catching the habit: a study of inequality of opportunity in smoking-related mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):175–194, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013399.

Frijters:2011:QCP

[131] Paul Frijters, Michael A. Shields, Stephen Wheatley Price, and Jenny Williams. Quantifying the cost of passive smoking on child health: evidence from children’s cotinine samples. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 195–212, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013400.

Tibshirani:2011:SPD

[132] Ryan J. Tibshirani, Andrew Price, and Jonathan Taylor. A statistician plays darts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):213–226, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013401.

Smith:2011:PPG

[133] T. M. F. Smith and Sonja Moore. Professor P. G. Moore, 1928–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 227–230, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013402.

Anonymous:2011:OPG

[134] Anonymous. Obituaries: P. G. Moore; David Harold Blackwell; Gopal K. Kanji; Andrew Samuel Christopher Ehrenberg; John Arthur Bound. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 227–238, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brillinger:2011:DHB

[135] David R. Brillinger. David Harold Blackwell, 1919–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):230– 233, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013403.

Gilchrist:2011:GKK

[136] Warren Gilchrist. Gopal K. Kanji, 1938– 2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):233–235, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013404.

Goodhardt:2011:ASC

[137] Gerald Goodhardt and Chris Chatfield. Andrew Samuel Christopher Ehrenberg, 1926–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):235–237, January 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013405.

Goodhardt:2011:JAB

[138] Gerald Goodhardt. John Arthur Bound, 1924–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):237–238, January 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013406.

Anonymous:2011:BRa

[139] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (1):239, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2011:BRBa

[140] Morteza Aalabaf-Sabaghi. Book review: Economic Modelling and Inference, by B. J. Christensen; N. M. Kiefer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (1):240, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013407.

Kumar:2011:BRBa

[141] Kuldeep Kumar. Book review: Demographic Forecasting, by F. Girosi; G. King. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):240–241, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013408.

Kumar:2011:BRBb

[142] Kuldeep Kumar. Book review: Robust Statistics, 2nd edn, by P. J. Huber; E. M. Ronchetti. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):241–242, January 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013409.

Aalabaf-Sabaghi:2011:BRBb

[143] Morteza Aalabaf-Sabaghi. Book review:

Macroeconomic Patterns and Stories: a Guide for MBAs, by E. E. Leamer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (1):242, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013410.

Sprent:2011:BRB

[144] P. Sprent. Book review: A Practical Guide to Scientific Data Analysis, by D. Livingstone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):242–243, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013411.

VanderWeele:2011:BRB

[145] Tyler VanderWeele. Book review: Causality, 2nd edn, by J. Pearl. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 243–244, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013412.

Bedogni:2011:BRBa

[146] Giorgio Bedogni. Book review: Design of Observational Studies, by P. R. Rosenbaum. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):244–245, January 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013413.

Ginestet:2011:BRB

[147] Cedric Ginestet. Book review: ggplot2: Elegant Graphics for Data Analysis, by H. Wickham. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1):245, January 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013414.

Ginestet:2011:BRG

[148] Cedric Ginestet. Book reviews: ggplot2: Elegant graphics for data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(1): 245–246, January 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wild:2011:TMA

[149] C. J. Wild, M. Pfannkuch, M. Regan, and N. J. Horton. Towards more accessible conceptions of statistical inferences [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2): 247–295, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014400.

Francesconi:2011:CEM

[150] Marco Francesconi, Holly Sutherland, and Francesca Zantomio. A comparison of earnings measures from longitudinal and cross-sectional surveys: evidence from the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):297–326, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23014401.

Lui:2011:QBS

[151] Silvia Lui, James Mitchell, and Martin Weale. Qualitative business surveys: signal or noise? Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):327–348, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23014402.

Lesnard:2011:ISW

[152] Laurent Lesnard and Man Yee Kan.

Investigating scheduling of work: a two-stage optimal matching analysis of workdays and workweeks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):

349–368, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014403.

Stuart:2011:UPS

[153] Elizabeth A. Stuart, Stephen R. Cole, Catherine P. Bradshaw, and Philip J. Leaf. The use of propensity scores to assess the generalizability of results from randomized trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2): 369–386, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014404.

Jaffa:2011:SEB

[154] Miran A. Jaffa, Robert F. Woolson, and Stuart R. Lipsitz. Slope estimation for bivariate longitudinal outcomes adjusting for informative right censoring by using a discrete survival model: application to the renal transplant cohort. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (2):387–402, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014405.

Asaduzzaman:2011:OLN

[155] Md Asaduzzaman and Thierry J. Chaussalet. An overflow loss network model for capacity planning of a perinatal network. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):403–417, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014406.

Plewis:2011:CVE

[156] Ian Plewis. Contextual variations in ethnic group differences in educational attainments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):419–437, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23014407.

Frale:2011:EMI

[157] Cecilia Frale, Massimiliano Marcellino, Gian Luigi Mazzi, and Tommaso Proietti. EUROMIND: a monthly indicator of the euro area economic conditions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):439–470, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014408.

Scarf:2011:DRS

[158] Philip Scarf, Xin Shi, and Sohail Akhtar.

On the distribution of runs scored and batting strategy in test cricket. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2): 471–497, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014409.

McCullagh:2011:JAN

[159] Peter McCullagh, Bob Gilchrist, and Roger Payne. John Ashworth Nelder, 1924–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):499–502, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23014410.

Anonymous:2011:OJA

[160] Anonymous. Obituaries: John Ashworth Nelder; Julian Besag. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2): 499–504, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Diggle:2011:JBF

[161] Peter J. Diggle and Peter J. Green. Julian Besag FRS, 1945–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2): 502–504, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23014411.

Anonymous:2011:BRb

[162] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):505, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2011:BRBa

[163] S. E. Lazic and F. Hoffmann-La Roche. Book review: Model based Inference in the Life Sciences: a Primer on Evidence, by D. R. Anderson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):506, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23014412.

Shalabh:2011:BRBa

[164] Shalabh. Book review: Measurement Error: Models, Methods and Applications, by J. P. Buonaccorsi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):

506–507, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014413.

Corbyn:2011:BRB

[165] Judith Corbyn. Book review: Time Series Analysis with Applications in R, 2nd edn by J. D. Cryer; K. Chan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):507, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014414.

Shalabh:2011:BRBb

[166] Shalabh. Book review: Fundamentals of Probability: a First Course, by A. Dasgupta. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):507–508, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014415.

Shalabh:2011:BRBc

[167] Shalabh. Book review: Nonparametric Statistical Inference, 5th edn, by J. Dickinson Gibbons; S. Chakraborti. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (2):508–509, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014416.

Stoyanov:2011:BRBa

[168] Jordan Stoyanov. Book review: Simulation and Inference for Stochastic Differential Equations: with R Examples, by S. M. Iacus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):509, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23014417.

Pacheco:2011:BRBa

[169] Gonzalo Durán Pacheco and F. Hoffmann-La Roche Basel. Book review: Graphics for Statistics and Data Analysis with R, by K. J. Keen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):509–510, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23014418.

Stoyanov:2011:BRBb

[170] Jordan Stoyanov. Book review: Stochastic Financial Models, by D. Kennedy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (2):510–511, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014419.

Bedogni:2011:BRBb

[171] Giorgio Bedogni. Book review: Applying Quantitative Bias Analysis to Epidemiologic Data, by T. L. Lash; M. P. Fox; A. K. Fink. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):511–512, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23014420.

Alexander:2011:BRB

[172] Neal Alexander. Book review: Bayesian Disease Mapping: Hierarchical Modeling in Spatial Epidemiology, by A. Lawson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):512–513, April 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23014421.

Lark:2011:BRB

[173] R. M. Lark. Book review: Statistical Methods for Estimating Petroleum Resources, by P.-J. Lee. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):513, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23014422.

Ugarte:2011:BRBa

[174] M. Dolores Ugarte. Book review: Introduction to Spatial Econometrics, by J. LeSage; R. K. Pace. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):513–514, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23014423.

Gillard:2011:BRBa

[175] Jonathan Gillard. Book review: Matrix Computations and Semiseparable Matrices: vol. 1, Linear Systems; vol. 2, Eigenvalue and Singular Value Methods, by R. Vandebril; M. Van Barel; N. Mastronardi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):514–515, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23014424.

Galbraith:2011:ARN

[176] Jane Galbraith and Mervyn Stone. The abuse of regression in the National Health Service allocation formulae: response to the Department of Health’s 2007 ‘Resource allocation research paper’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):517–528, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013508.

Gibson:2011:RAE

[177] Alex Gibson and Sheena Asthana. Resource allocation for English local government: a critique of the four-block model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):529–546, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013509.

Anonymous:2011:DMR

[178] Anonymous. Discussion on the meeting on ‘Resource allocation models’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):547–567, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013510.

Chevalier:2011:IAV

[179] Arnaud Chevalier and Antony Fielding. An introduction to anchoring vignettes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):569–574, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013511.

vanSoest:2011:VUA

[180] Arthur van Soest, Liam Delaney, Colm Harmon, Arie Kapteyn, and James P. Smith. Validating the use of anchoring vignettes for the correction of response scale differences in subjective questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):575–595, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013512.

Vonkova:2011:AVM

[181] Hana Voňková and Patrick Hullegie. Is the anchoring vignette method sensitive to the domain and choice of the vignette? Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):597–620, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013513.

Angelini:2011:DRW

[182] Viola Angelini, Danilo Cavapozzi, and Omar Paccagnella. Dynamics of reporting work disability in Europe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 621–638, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013514.

dUva:2011:ERI

[183] Teresa Bago d’Uva, Maarten Lindeboom, Owen O’Donnell, and Eddy van Doorslaer. Education-related inequity in healthcare with heterogeneous reporting of health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):639–664, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013515.

Paccagnella:2011:AVS

[184] Omar Paccagnella. Anchoring vignettes with sample selection due to nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):665–687, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013516.

Chung:2011:LCP

[185] Hwan Chung, James C. Anthony, and Joseph L. Schafer. Latent class profile analysis: an application to stage sequential processes in early onset drinking behaviours. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):689–712, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013517.

Coelho:2011:FME

[186] Edviges Coelho and Luis C. Nunes. Forecasting mortality in the event of a structural change. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):713–736, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23013518.

Siedler:2011:PUY

[187] Thomas Siedler. Parental unemployment and young people’s extreme rightwing party affinity: evidence from panel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):737–758, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013519.

Grab:2011:IBF

[188] Johannes Gräb and Michael Grimm. Inequality in Burkina Faso — to what extent do household, community and regional factors matter? Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 759–784, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013520.

Saville:2011:AVC

[189] Benjamin R. Saville, Amy H. Herring, and Jay S. Kaufman. Assessing variance components in multilevel linear models using approximate Bayes factors: a case-study of ethnic disparities in birth weight. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):785–804, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013521.

Gastwirth:2011:EFN

[190] Joseph L. Gastwirth, Wesley O. Johnson, and Hiro Hikawa. Estimating the fraction of ‘non-genuine’ artwork by Henry Moore for sale on eBay: application of latent class screening test methodology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):805–822, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23013522.

Bellhouse:2011:NLH

[191] David R. Bellhouse. A new look at Halley’s life table. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):823–832, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23013523.

Leckie:2011:NLS

[192] George Leckie and Harvey Goldstein. A note on ‘the limitations of school league tables to inform school choice’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):833–836, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013524.

Read:2011:JRA

[193] Kenneth Read. John Robert Ashford, 1929–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):837–838, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013525.

Anonymous:2011:OJR

[194] Anonymous. Obituaries: John Robert Ashford; Julian Robert Leslie. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 837–839, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Seneta:2011:JRL

[195] Eugene Seneta. Julian Robert Leslie, 1948–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):838–839, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013526.

Anonymous:2011:BRc

[196] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):841, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dasgupta:2011:BRB

[197] Aisha Dasgupta. Book review: Food, Economics, and Health, by A. Bhargava. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):842, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013527.

Goda:2011:BRB

[198] David Goda. Book review: Lotteries for Education, by C. Boyle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 842–843, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013528.

Gillard:2011:BRBb

[199] Jonathan Gillard. Book review: Circular and Linear Regression: Fitting Circles and Lines, by Least Squares by N. Chernov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):843, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013529.

Lazic:2011:BRBb

[200] S. E. Lazic. Book review: The R Book, by M. J. Crawley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):843–844, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013530.

Leckie:2011:BRB

[201] George Leckie. Book review: Handbook of Advanced Multilevel Analysis, by J. J. Hox; J. K. Roberts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 844–845, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013531.

Ravi:2011:BRB

[202] Sreenivasan Ravi. Book review: New Perspectives in Stochastic Geometry, by W. S. Kendall; I. Molchanov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3): 845–846, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013532.

Aalabaf-Sabaghi:2011:BRBc

[203] Morteza Aalabaf-Sabaghi. Book review: Monte Carlo Methods and Models in Finance and Insurance, by R. Korn; E. Korn; G. Kroisandt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):846, July 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23013533.

Stoyanov:2011:BRBc

[204] Jordan Stoyanov. Book review: Modeling and Analysis of Stochastic Systems, 2nd edn, by V. G. Kulkarni. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):847, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23013534.

Aachen:2011:BRB

[205] Thomas Aachen. Book review: Applied Statistics for Engineers and Physical Scientists, 3rd edn, by J. Ledolter; R. V. Hogg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):847–848, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23013535.

Hochkirchen:2011:BRB

[206] Thomas Hochkirchen. Book review: Applied Statistics for Engineers and Physical Scientists, 3rd edn, by J. Ledolter; R. V. Hogg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):847–848, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2011:BRBd

[207] Morteza Aalabaf-Sabaghi. Book reviews: Applied Statistics for Business and Economics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):848, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2011:BRBa

[208] Anonymous. Book review: Modeling and Analysis of Stochastic Systems, 2nd edn, by V. G. Kulkarni. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):848, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013536.

Ugarte:2011:BRBb

[209] M. Dolores Ugarte. Book review: Statistical Methods for Disease Clustering, by T. Tango. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(3):848–849, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23013537.

Bedogni:2011:BRBc

[210] Giorgio Bedogni. Book review: Statistics in Plain English, by T. C. Urdan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):849–850, July 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23013538.

Anand:2011:EMP

[211] Paul Anand, Martine Durand, and James Heckman. Editorial: The measurement of progress — some achievements and challenges. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):851– 855, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409685.

Bornmann:2011:MMA

[212] Lutz Bornmann, Rüdiger Mutz, Werner Marx, Hermann Schier, and Hans-Dieter Daniel. A multilevel modelling approach to investigating the predictive validity of editorial decisions: do the editors of a high profile journal select manuscripts that are highly cited after publication? Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 857–879, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409686.

Brown:2011:DCC

[213] James Brown, Owen Abbott, and Paul A. Smith. Design of the 2001 and 2011 Census Coverage Surveys for England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):881–906, October 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409687.

Forster:2011:SSM

[214] Martin Forster and S. D. Smith. Surviving slavery: mortality at Mesopotamia, a Jamaican sugar estate, 1762–1832. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 907–929, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409688.

Tanton:2011:SAE

[215] Robert Tanton, Yogi Vidyattama, Binod Nepal, and Justine McNamara. Small area estimation using a reweighting algorithm. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):931–951, October 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/41409689.

Mendola:2011:CIS

[216] Daria Mendola, Annalisa Busetta, and Anna Maria Milito. Combining the intensity and sequencing of the poverty experience: a class of longitudinal poverty indices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):953–973, October 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/41409690.

Tu:2011:SMA

[217] Wanzhu Tu, Pulak Ghosh, and Barry P. Katz. A stochastic model for assessing Chlamydia trachomatis transmission risk by using longitudinal observational data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):975–989, October 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409691.

Breunig:2011:ESD

[218] Robert Breunig and Rebecca McKibbin. The effect of survey design on household reporting of financial difficulty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 991–1005, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409692.

LeSage:2011:NOB

[219] James P. LeSage, R. Kelley Pace, Nina Lam, Richard Campanella, and Xingjian Liu. New Orleans business recovery in the aftermath of Hurricane Katrina. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1007–1027, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409693.

Durrant:2011:UPP

[220] Gabriele B. Durrant, Julia D’Arrigo, and Fiona Steele. Using paradata to predict best times of contact, conditioning on household and interviewer influences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1029–1049, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409694.

Park:2011:MMB

[221] Young-Hoon Park, Chang Hee Park, and Pulak Ghosh. Modelling member behaviour in on-line user-generated content sites: a semiparametric Bayesian approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):1051–1069, October 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409695.

Baker:2011:IBC

[222] Rose Baker and Ian G. McHale. Investigating the behavioural characteristics of lottery players by using a combination preference model for conscious selection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):1071–1086, October 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/41409696.

Isham:2011:RC

[223] Valerie Isham, Martin Dougherty, Sarah Simpson, Trevor Lewis, Stephen Senn, Andrew Garratt, Jill Leyland, Colin Aitken, Ron Smith, Martin Griffiths, Paul Gentry, Nicola Emmerson, Andy Garrett, Gerald Goodall, John Mc-Coll, Penny Bidgood, Debra Hurcomb, Neville Davies, Jon Forster, Martin Owen, Frank Duckworth, Julian Champkin, Mary Sweetland, Jenny Church, Liam Murray, Heiko Grossman, Richard Chandler, Vic Barnett, Julian Stander, Stephen Walters, Liz Harris, Carolyn Craggs, Richard Samworth, Antony Fielding, Jenny Lannon, David Martin, Maurice Marchant, G. Robin Henderson, Daphne Kounali, Mairi Spowage, Ben Torsney, Malcolm Hall, Matthew Sperrin, Paul Baxter, John Logsdon, Laura Bonnett, Alan Gordon, Jemma Hopewell, Sara Hilditch, John Eales, Peter Kimani, Tom Grimes, Sayer Vincent, and Helen Elliott. Report of the Council for 2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1087–1174, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409697.

Sadhukhan:2011:CPE

[224] Sudipta Sadhukhan, Zhihui Liu, and James A. Hanley. Correspondence: Practical and efficient estimates of one’s accuracy in darts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):1175–1177, October 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409698.

Harrison:2011:ORM

[225] Anne Harrison. Obituary: Richard Malvern Allen, 1933–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):1178– 1179, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409699.

Anonymous:2011:BRd

[226] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1180, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Metcalfe:2011:BRB

[227] Andrew V. Metcalfe. Book review:

Bayesian Ideas and Data Analysis — an Introduction for Scientists and Statisticians, by R. Christensen; W. Johnson; A. Branscum; T. E. Hanson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1181, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409700.

Pacheco:2011:BRBb

[228] Gonzalo Durán Pacheco and F. Hoffmann-La Roche. Book review: Multiple Testing Problems in Pharmaceutical Statistics, by A. Dmitrienko; A. C. Tamhane; F. Bretz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4):1181–1182, October 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/41409701.

Lazic:2011:BRBc

[229] S. E. Lazic and F. Hoffmann-La Roche. Book review: Meta-analysis and Combining Information in Genetics and Genomics, by R. Guerra; D. R. Goldstein. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1182–1183, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409702.

Shade:2011:BRBa

[230] John Shade. Book review: The Hockey Stick Illusion, by A. W. Montford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1183–1185, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409703.

Shade:2011:BRBb

[231] John Shade. Book review: Acceptance Sampling in Quality Control, by E. G. Schilling; D. V. Neubauer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1185–1186, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409704.

Anonymous:2011:IAI

[232] Anonymous. Index of authors: Index of authors, volume 174, 2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(4): 1187–1192, October 2011. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Spiegelhalter:2012:SMH

[233] David Spiegelhalter, Christopher Sherlaw-Johnson, Martin Bardsley, Ian Blunt, Christopher Wood, and Olivia Grigg. Statistical methods for healthcare regulation: rating, screening and surveillance [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):1–47, January 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409707.

Unkel:2012:SMP

[234] Steffen Unkel, C. Paddy Farrington, Paul H. Garthwaite, Chris Robertson, and Nick Andrews. Statistical methods for the prospective detection of infectious disease outbreaks: a review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (1):49–82, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409708.

McCloud:2012:IFD

[235] Nadine McCloud and Subal C. Kumbhakar. Institutions, foreign direct investment and growth: a hierarchical Bayesian approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):83–105, January 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409709.

Saez:2012:SVR

[236] Marc Saez, Maria Antónia Barceló, Carmen Martos, Carme Saurina, Rafael Marcos-Gragera, Gemma Renart, Ricardo Ocaña-Riola, Cristina Feja, and Tomás Alcalá. Spatial variability in relative survival from female breast cancer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 107–134, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409710.

Wichert:2012:WFS

[237] Laura Wichert and Ralf A. Wilke. Which factors safeguard employment?: an analysis with misclassified German register data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):135–151, January 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        41409711.

Amann:2012:RTT

[238] Roland A. Amann and Tobias J. Klein. Returns to type or tenure? Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 153–166, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409712.

Atkinson:2012:CGO

[239] Anthony B. Atkinson, Peter G. Backus, John Micklewright, Cathy Pharoah, and Sylke V. Schnepf. Charitable giving for overseas development: UK trends over a quarter century. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):167–190, January 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
41409713.

Lu:2012:SRD

[240] Xin Lu, Linus Bengtsson, Tom Britton, Martin Camitz, Beom Jun Kim, Anna Thorson, and Fredrik Liljeros. The sensitivity of respondent-driven sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 191–216, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409714.

Blom:2012:ECC

[241] Annelies G. Blom. Explaining crosscountry differences in survey contact rates: application of decomposition methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):217–242, January 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/41409715.

Craig:2012:SNE

[242] Peter S. Craig, Graeme L. Hickey, Robert Luttik, and Andy Hart. Species non-exchangeability in probabilistic ecotoxicological risk assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 243–262, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/41409716.

Stewart:2012:QEC

[243] Mark B. Stewart. Quantile estimates of counterfactual distribution shifts and the effect of minimum wage increases on the wage distribution. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 263–287, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409717.

Lynn:2012:IQM

[244] Peter Lynn, Annette Jäckle, Stephen P. Jenkins, and Emanuela Sala. The impact of questioning method on measurement error in panel survey measures of benefit receipt: evidence from a validation study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1):289–308, January 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/41409718.

Currie:2012:ORG

[245] Iain Currie. Obituary: Roger Gray, BSc, FFA(Hon), CStat, FSS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(1): 309–310, January 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/41409719.

Day:2012:EPR

[246] Simon Day. Editorial: Publishing research — for whose benefit? Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2): 311–312, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251239.

Isham:2012:ESU

[247] Valerie Isham. The evolving Society:

united we stand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):315–335, April 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251240.

Rice:2012:VHS

[248] Nigel Rice, Silvana Robone, and Peter C. Smith. Vignettes and health systems responsiveness in cross-country comparative analyses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):337–357, April 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/

stable/23251241. See discussion [249].

vanSoest:2012:DPR

[249] Arthur van Soest, C. Harmon, Mary Sweetland, D. R. Cox, Omar Paccagnella, Nicole Valentine, Elvan Ceyhan, Paul Allin, Chris Skinner, Andrew Gelman, Emilio Porcu, Ramón Giraldo, Carlos Alonso Malavar, José E. Chacón, Jesús Montanero, Luisa Corrado, Melvyn Weeks, Venkata K. Jandhyala, Stergios B. Fotopoulos, Thomas King, and Jorge Mateu. Discussion on the paper by Rice, Robone and Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (2):357–369, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251886.

Neumann:2012:QWE

[250] C. Neumann, I. W. Evett, and J. Skerrett. Quantifying the weight of evidence from a forensic fingerprint comparison: a new paradigm. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2): 371–396, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251242. See dis-

cussion [251].

Aitken:2012:DPN

[251] C. G. G. Aitken, David J. Balding, Bernard Silverman, Michael Risinger, Simon A. Cole, David A. Stoney, Ian Hunt, Leonard G. Butt, José E. Chacón, Nick Fieller, Stephen E. Fienberg, J. L. Gastwirth, Q. Pan, Ramón Giraldo, Carlos Alonso Malaver, Emilio Porcu, Thomas Hotz, Axel Munk, Venkata K. Jandhyala, Stergios B. Fotopoulos, Joseph B. Kadane, David H. Kaye, Steffen Lauritzen, Robert Cowell, Therese Graversen, Stephanie Llewelyn, Jorge Mateu, Robert S. Ramotowski, Bernard Robertson, G. A. Vignaux, Charles E. H. Berger, Michael J. Saks, Ashley M. Votruba, Hal Stern, William C. Thompson, and Max Welling. Discussion on the paper by Neumann, Evett and Skerrett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):396–415, April 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251887. See

[250].

Gueorguieva:2012:JML

[252] Ralitza Gueorguieva, Robert Rosenheck, and Haiqun Lin. Joint modelling of longitudinal outcome and interval-censored competing risk dropout in a schizophrenia clinical trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):417–433, April 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251243.

Seneta:2012:VPL

[253] Eugene Seneta. Victorian probability and Lewis Carroll. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):435–451, April 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251244.

Moscone:2012:SIP

[254] F. Moscone, E. Tosetti, and G. Vittadini. Social interaction in patients’ hospital choice: evidence from Italy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (2):453–472, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251245.

Riebler:2012:GSD

[255] Andrea Riebler, Leonhard Held, Håvard Rue, and Matthias Bopp. Genderspecific differences and the impact of family integration on time trends in agestratified Swiss suicide rates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):473–490, April 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/23251246. See [248].

Billari:2012:SPF

[256] F. C. Billari, R. Graziani, and E. Melilli. Stochastic population forecasts based on conditional expert opinions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2): 491–511, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251247.

Bunouf:2012:AIL

[257] P. Bunouf, J.-M. Grouin, G. Molenberghs, and G. Koch. Analysing intensive longitudinal data after summarization at landmarks: an application to daily pain evaluation in a clinical trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):513–534, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251248.

Lee:2012:FDT

[258] Myoung jae Lee and Fali Huang. Finding dynamic treatment effects under anticipation: the effects of spanking on behaviour. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):535–567, April 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23251249.

Sweeting:2012:MPC

[259] M. J. Sweeting and S. G. Thompson. Making predictions from complex longitudinal data, with application to planning monitoring intervals in a national screening programme. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):

569–586, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251250.

Kluve:2012:ECT

[260] Jochen Kluve, Hilmar Schneider, Arne Uhlendorff, and Zhong Zhao. Evaluating continuous training programmes by using the generalized propensity score. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (2):587–617, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251251.

Choi:2012:MPT

[261] Young Ku Choi, Wesley O. Johnson, Geoff Jones, Andres Perez, and Mark C. Thurmond. Modelling and predicting temporal frequency of foot-and-mouth disease cases in countries with endemic foot-and-mouth disease. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2): 619–636, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251252.

Zhang:2012:BNL

[262] Yanwei Zhang, Vanja Dukic, and James Guszcza. A Bayesian non-linear model for forecasting insurance loss payments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (2):637–656, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251253.

Lawrance:2012:PWL

[263] Tony Lawrance and Eddie McKenzie. Peter A. W. Lewis, 1932–2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2): 657–659, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251254.

Anonymous:2012:OPW

[264] Anonymous. Obituaries: Peter A. W. Lewis, 1932–2011; Wyndham Somerville Paige. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):657–660, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Springall:2012:WSP

[265] Tony Springall. Wyndham Somerville Paige. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(2):659–660, April 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251255.

Johnstone:2012:LOE

[266] D. J. Johnstone. Log-optimal economic evaluation of probability forecasts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):661–681, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251168. See dis-

cussion [267].

Satchell:2012:DPJ

[267] Steve Satchell, Christopher A. T. Ferro, Peter W. F. Smith, Michael P. Clements, Robert L. Winkler, Michael Crowe, Carlos Eduardo Alonso Malaver, Emilio Porcu, Ramón Moncho Giraldo, Jorge Mateu, Oscar O. Melo, and Carlos E. Melo. Discussion on the paper by Johnstone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):681–689, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251883. See

[266].

Reboussin:2012:SSE

[268] Beth A. Reboussin, John S. Preisser, Eun-Young Song, and Mark Wolfson. Sample size estimation for alternating logistic regressions analysis of multilevel randomized community trials of underage drinking. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):691–712, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251169.

Galbraith:2012:AGD

[269] John W. Galbraith and Simon van Norden. Assessing gross domestic product and inflation probability forecasts derived from Bank of England fan charts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (3):713–727, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251170.

Jones:2012:IPP

[270] Hayley E. Jones and David J. Spiegelhalter. Improved probabilistic prediction of healthcare performance indicators using bidirectional smoothing models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):729–747, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23251171.

Pavlopoulos:2012:HRM

[271] Dimitris Pavlopoulos, Ruud Muffels, and Jeroen K. Vermunt. How real is mobility between low pay, high pay and non-employment? Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):749–773, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23251172.

Mealli:2012:EEU

[272] Fabrizia Mealli and Carla Rampichini. Evaluating the effects of university grants by using regression discontinuity designs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):775–798, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23251173.

Farewell:2012:HMW

[273] Vern Farewell, Tony Johnson, and Rosemary Gear. Hilda Mary Woods MBE, DSc, LRAM, FSS (1892–1971): reflections on a Fellow of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):799–811, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23251174.

Powell:2012:C

[274] Matthew Powell, Paul Allin, and Joe Grice. Correspondence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):813–814, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/23251175.

Sprent:2012:SCP

[275] P. Sprent, E. E. Bassett, and D. A.

Preece. Stanley Clifford Pearce, 1914– 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):815–817, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251176.

Anonymous:2012:OSC

[276] Anonymous. Obituaries: Stanley Clifford Pearce; Anthony Peter Macmillan Coxon; Derrick Norman Lawley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (3):815–819, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jones:2012:APM

[277] Charles Jones. Anthony Peter Macmillan Coxon, 1938–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):817–818, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/23251177.

Finney:2012:DNL

[278] David J. Finney. Derrick Norman Lawley, 1915–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):818–819, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251178.

Anonymous:2012:BRa

[279] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):821, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2012:BRBa

[280] Shalabh. Book review: Methods for Estimation and Inference in Modern Econometrics, by S. Anatolyev; N. Gospodinov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):822, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251179.

Ginestet:2012:BRB

[281] Cedric E. Ginestet. Book review: Spectral Analysis of Large Dimensional Random Matrices, 2nd edn, by Z. Bai; J. W. Silverstein. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):822–823, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251180.

Kumar:2012:BRBa

[282] Kuldeep Kumar. Book review: The Methodology and Practice of Econometrics: a Festschrift in Honour of David F. Hendry, by J. L. Castle; N. Shephard. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):823–824, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251181.

Kostenko:2012:BRB

[283] Andrey Kostenko. Book review: Introductory Time Series with R, by P. S. P. Cowpertwait; A. V. Metcalfe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):824, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251182.

Stoyanov:2012:BRB

[284] Jordan Stoyanov. Book review: The Oxford Handbook of Nonlinear Filtering, by D. Crisan; B. Rozovskii. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3): 824–825, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251183.

Gillard:2012:BRB

[285] Jonathan Gillard. Book review: Approaching Multivariate Analysis, a Practical Introduction, 2nd edn, by P. Dugard; J. Todman; H. Staines. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):825, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251184.

Chakrapani:2012:BRB

[286] Chuck Chakrapani. Book review: Multivariable Modeling and Multivariate Analysis for the Behavioral Sciences, by B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):825–826, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251185.

Denny:2012:BRB

[287] Frances Denny. Book review: SAS and R: Data Management, Statistical Analysis, and Graphics, by K. Kleinman; N. J. Horton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):826–827, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251186.

Shalabh:2012:BRBb

[288] Shalabh. Book review: Design of Experiments: an Introduction based on Linear Models, by M. D. Morris. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):827, July 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23251187.

Lazic:2012:BRB

[289] S. E. Lazic and F. Hoffmann-La Roche. Book review: Introducing Monte Carlo Methods with R, by C. P. Robert; G. Casella. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):827–828, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251188.

Kumar:2012:BRBb

[290] Kuldeep Kumar. Book review: A Beginner’s Guide to Structural Equation Modeling, 3rd edn, by R. E. Schumacker; R. G. Lomax. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):828–829, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251189.

Cortina-Borja:2012:BRB

[291] Mario Cortina-Borja. Book review:

Handbook of Parametric and Nonparametric Statistical Procedures, 5th edn, by D. J. Sheskin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3):829, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23251190.

Fanshawe:2012:BRB

[292] Tom Fanshawe. Book review: Statistical Methods in Environmental Epidemiology, by D. C. Thomas. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(3): 829–830, July 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23251191.

Aalen:2012:CMT

[293] Odd O. Aalen, Kjetil Røysland, Jon Michael Gran, and Bruno Ledergerber. Causality, mediation and time: a dynamic viewpoint. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):831–861, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355305.

Badunenko:2012:WWH

[294] Oleg Badunenko, Daniel J. Henderson, and Subal C. Kumbhakar. When, where and how to perform efficiency estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):863–892, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/23355306.

Madden:2012:POI

[295] David Madden. A profile of obesity in Ireland, 2002–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):893–914, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355307.

Micklewright:2012:NRB

[296] John Micklewright, Sylke V. Schnepf, and Chris Skinner. Non-response biases in surveys of schoolchildren: the case of the English Programme for International Student Assessment (PISA) samples. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):915–938, October 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23355308.

Kavanagh:2012:SSI

[297] Kimberley Kavanagh, Chris Robertson, Heather Murdoch, George Crooks, and Jim McMenamin. Syndromic surveillance of influenza-like illness in Scotland during the influenza A H1N1v pandemic and beyond. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):939–958, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355309.

Abakuks:2012:SPM

[298] Andris Abakuks. The synoptic problem:

on Matthew’s and Luke’s use of Mark. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 959–975, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355310.

Isham:2012:RC

[299] Valerie Isham. Report of the Council for 2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):977–1056, October 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355311.

Phillpotts:2012:SVC

[300] Greg Phillpotts. Stella Vivian Cunliffe MBE, January 12th, 1917–January 19th, 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1057–1059, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355312.

Anonymous:2012:OSV

[301] Anonymous. Obituaries: Stella Vivian Cunliffe; James Durbin; John N. R. Jeffers; Francis Henry Charles Marriott; Rod McDonald; George Vaughan Dyke; Wilfred J. Corlett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1057–1070, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koopman:2012:JDF

[302] Siem Jan Koopman. James Durbin, FBA, 1923–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1060–1064, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355313.

Rennolls:2012:JRJ

[303] Keith Rennolls. John N. R. Jeffers, 1926–2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1064–1067, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355314.

Armitage:2012:FHC

[304] Peter Armitage and John F. Bithell. Francis Henry Charles Marriott, 1926– 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1067–1068, October 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/23355315.

Goldstein:2012:RM

[305] Harvey Goldstein. Rod McDonald, 1928–2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1068–1069, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355316.

Preece:2012:GVD

[306] Donald Preece, Geoff Clarke, and David Bartholomew. George Vaughan Dyke, November 10th, 1921–March 29th, 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1069–1070, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355317.

Anonymous:2012:WJC

[307] Anonymous. Wilfred J. Corlett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1070, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355318.

Anonymous:2012:BRb

[308] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1071, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Leong:2012:BRB

[309] Chee Kian Leong. Book review:

Non-parametric Econometrics, by I. Ahamada; E. Flachaire. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4):1072, October 2012. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/23355319.

Ugarte:2012:BRB

[310] M. Dolores Ugarte. Book review: Visualizing Data Patterns with Micromaps, by D. B. Carr; L. W. Pickle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1072–1073, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355320.

Aalabaf-Sabaghi:2012:BRB

[311] Morteza Aalabaf-Sabaghi. Book review: Networks, Crowds and Markets: Reasoning about a Highly Connected World, by D. Easley; J. Kleinberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1073, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355321.

Beh:2012:BRB

[312] Eric J. Beh. Book review: Biplots in Practice, by M. Greenacre. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1073–1074, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355322.

Jowaheer:2012:BRB

[313] Vandna Jowaheer. Book review: Dynamic Mixed Models for Familial Longitudinal Data, by B. C. Sutradhar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1074–1075, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355323.

Anonymous:2012:IAI

[314] Anonymous. Index of authors: Index of authors, volume 175, 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175(4): 1077–1081, October 2012. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Durrant:2013:EUP

[315] Gabriele Durrant and Frauke Kreuter.

Editorial: The use of paradata in social survey research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):1–3, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355174.

Imai:2013:EDI

[316] Kosuke Imai, Dustin Tingley, and Teppei Yamamoto. Experimental designs for identifying causal mechanisms. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176 (1):5–32, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355175. See dis-

cussion [317].

Mealli:2013:DPI

[317] Fabrizia Mealli, Carlo Berzuini, Guanglei Hong, Brian L. Egleston, Roland R. Ramsahai, Theis Lange, Andrew Gelman, Manabu Kuroki, Jeffrey M. Albert, John G. Bullock, Donald P. Green, Vanessa Didelez, David Draper, Adam N. Glynn, Booil Jo, David A. Kenny, Víctor Leiva, Emilio Porcu, N. T. Longford, David P. MacKinnon, Jorge Mateu, Oscar O. Melo, Carlos E. Melo, Alessandra Mattei, T. S. Richardson, J. M. Robins, Donald B. Rubin, Marc Saez, Michael E. Sobel, Tyler J. VanderWeele, and Richard A. Emsley. Discussion on the paper by Imai, Tingley and Yamamoto. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176 (1):32–51, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355176. See

[316].

Best:2013:BAC

[318] Nicky Best, Deborah Ashby, Frank Dunstan, David Foreman, and Neil McIntosh. A Bayesian approach to complex clinical diagnoses: a case-study in child abuse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):53–83, January 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355177. See

discussion [319].

Dawid:2013:DPB

[319] A. P. Dawid, Tim J. Cole, Rosalind L. Smyth, Ian Hunt, Daniel Farewell, Silvia Liverani, Marta Blangiardo, Woody Caan, Jo Sibert, Sheila M. Bird, Christopher Verity, D. S. Coad, L. I. Pettit, Carlos Díaz-Avalos, Jorge Mateu, Michael R. Elliott, Andrew Gelman, Thomas A. Louis, Emilio Porcu, Alessandro Zini, and Marc Saez. Discussion on the paper by Best, Ashby, Dunstan, Foreman and McIntosh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176 (1):83–96, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355178. See

[318].

Dorling:2013:FCF

[320] Danny Dorling. Fairness and the changing fortunes of people in Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1): 97–128, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355179.

Olson:2013:DNR

[321] Kristen Olson. Do non-response followups improve or reduce data quality?: a review of the existing literature. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1): 129–145, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355180.

Biemer:2013:ULE

[322] Paul P. Biemer, Patrick Chen, and Kevin Wang. Using level-of-effort paradata in non-response adjustments with application to field surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1): 147–168, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355181.

Luiten:2013:TFD

[323] Annemieke Luiten and Barry Schouten.

Tailored fieldwork design to increase representative household survey response: an experiment in the survey of consumer satisfaction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):169–189, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355182.

Conrad:2013:ISS

[324] Frederick G. Conrad, Jessica S. Broome, José R. Benkí, Frauke Kreuter, Robert M. Groves, David Vannette, and Colleen McClain. Interviewer speech and the success of survey invitations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1): 191–210, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355183.

West:2013:EQU

[325] Brady T. West. An examination of the quality and utility of interviewer observations in the national survey of family growth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):211–225, January 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355184.

Casas-Cordero:2013:AME

[326] C. Casas-Cordero, F. Kreuter, Y. Wang, and S. Babey. Assessing the measurement error properties of interviewer observations of neighbourhood characteristics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):227–249, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355185.

Durrant:2013:AIC

[327] Gabriele B. Durrant, Julia D’Arrigo, and Fiona Steele. Analysing interviewer call record data by using a multilevel discrete time event history modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):251–269, January 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355186.

Couper:2013:UPE

[328] Mick P. Couper and Frauke Kreuter.

Using paradata to explore item level response times in surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1): 271–286, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355187.

Harrison:2013:ON

[329] Anne Harrison and John Walton. Owen Nankivell, 1927–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):287– 288, January 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355188.

Haigh:2013:JAB

[330] John Haigh. John Alfred Bather, 1936– 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(1):288–290, January 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355189.

Paloyo:2013:EOH

[331] Alfredo R. Paloyo. Editorial: On the orthography of heteros*edasticity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 291–293, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355192.

Popli:2013:GWD

[332] Gurleen K. Popli. Gender wage differentials in Mexico: a distributional approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):295–319, February 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355193.

Smallman-Raynor:2013:ESI

[333] Matthew Smallman-Raynor and Andrew D. Cliff. Epidemics in semi-isolated communities: statistical perspectives on acute childhood diseases in English public boarding schools, 1930–1939. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 321–346, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355194.

Brandes:2013:GSL

[334] Leif Brandes, Egon Franck, and Philipp Theiler. The group size and loyalty of football fans: a two-stage estimation procedure to compare customer potentials across teams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):347–369, February 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355195.

deMunnik:2013:AAN

[335] Daniel de Munnik, Mark Illing, and David Dupuis. Assessing the accuracy of non-random business conditions surveys: a novel approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 371–388, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355196.

Neelon:2013:SPH

[336] Brian Neelon, Pulak Ghosh, and Patrick F. Loebs. A spatial Poisson hurdle model for exploring geographic variation in emergency department visits. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 389–413, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355197.

Falcaro:2013:ACL

[337] Milena Falcaro, Neil Pendleton, and Andrew Pickles. Analysing censored longitudinal data with non-ignorable missing values: depression in older age. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 415–430, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355198.

Chowdry:2013:WPH

[338] Haroon Chowdry, Claire Crawford, Lorraine Dearden, Alissa Goodman, and Anna Vignoles. Widening participation in higher education: analysis using linked administrative data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 431–457, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355199.

Holly:2013:AFC

[339] Sean Holly, Ivan Petrella, and Emiliano Santoro. Aggregate fluctuations and the cross-sectional dynamics of firm growth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 459–479, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355200.

Belenkiy:2013:MRM

[340] Ari Belenkiy. The Master of the Royal Mint: how much money did Isaac Newton save Britain? Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):481–498, February 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355201.

Fumagalli:2013:EMR

[341] Laura Fumagalli, Heather Laurie, and Peter Lynn. Experiments with methods to reduce attrition in longitudinal surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):499–519, February 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/23355202.

Langel:2013:VEG

[342] Matti Langel and Yves Tillé. Variance estimation of the Gini index: revisiting a result several times published. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 521–540, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355203.

Blanden:2013:IPI

[343] Jo Blanden, Paul Gregg, and Lindsey Macmillan. Intergenerational persistence in income and social class: the effect of within-group inequality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2): 541–563, February 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/23355204.

Tollenaar:2013:WMP

[344] N. Tollenaar and P. G. M. van der Heijden. Which method predicts recidivism best?: a comparison of statistical, machine learning and data mining predictive models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):565–584, February 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
23355205.

Martin:2013:DGF

[345] David Martin, Samantha Cockings, and Andrew Harfoot. Development of a geographical framework for census workplace data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):585–602, February 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355206.

Baker:2013:RRS

[346] Stuart G. Baker and Barnett S. Kramer. The risky reliance on small surrogate end point studies when planning a large prevention trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(2):603–608, February 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        23355207.

Paruolo:2013:RRV

[347] Paolo Paruolo, Michaela Saisana, and Andrea Saltelli. Ratings and rankings: voodoo or science? Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):609–634, June 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Johnstone:2013:MEV

[348] D. J. Johnstone, S. Jones, V. R. R. Jose, and M. Peat. Measures of the economic value of probabilities of bankruptcy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176 (3):635–653, June 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pavia:2013:SVR

[349] Jose M. Pavía and Antonio López-Quílez. Spatial vote redistribution in redrawn polling units. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3): 655–678, June 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Andersson:2013:SSS

[350] Christian Andersson and Per Johansson.

Social stratification and out-of-school learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):679–701, June 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Peracchi:2013:HTO

[351] Franco Peracchi and Claudio Rossetti.

The heterogeneous thresholds ordered response model: identification and inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):703–722, June 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koop:2013:FEC

[352] Gary Koop and Lise Tole. Forecasting the European carbon market. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3): 723–741, June 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hassani:2013:PID

[353] Hossein Hassani, Abdol S. Soofi, and Anatoly Zhigljavsky. Predicting inflation dynamics with singular spectrum analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):743–760, June 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Backus:2013:BCB

[354] Peter Backus and David Clifford. Are big charities becoming more dominant?: cross-sectional and longitudinal perspectives. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):761–776, June 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nam:2013:SAV

[355] Christopher F. H. Nam, Sach Mukherjee, Marco Schilk, and Joybrato Mukherjee. Statistical analysis of varieties of English. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):777–793, June 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Paddock:2013:BHS

[356] Susan M. Paddock and Terrance D. Savitsky. Bayesian hierarchical semiparametric modelling of longitudinal post-treatment outcomes from open enrolment therapy groups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3):795–808, June 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2013:Oa

[357] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(3): 809–814, June 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bird:2013:ECD

[358] Sheila M. Bird. Editorial: Counting the dead properly and promptly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 815–817, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pullinger:2013:SMI

[359] John Pullinger. Statistics making an impact. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):819–840, October 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Johnston:2013:HHC

[360] David W. Johnston, Michael E. R.

Nicholls, Manisha Shah, and Michael A. Shields. Handedness, health and cognitive development: evidence from children in the national longitudinal survey of youth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):841–860, October 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bockerman:2013:DHI

[361] Petri Böckerman, Alex Bryson, and Pekka Ilmakunnas. Does high involvement management lead to higher pay? Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 861–885, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jerrim:2013:SMR

[362] John Jerrim and Anna Vignoles. Social mobility, regression to the mean and the cognitive development of high ability children from disadvantaged homes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):

887–906, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Frimmel:2013:AMD

[363] Wolfgang Frimmel, Martin Halla, and Rudolf Winter-Ebmer. Assortative mating and divorce: evidence from Austrian register data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):907–929, October 2013. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Graham:2013:QEA

[364] Daniel J. Graham, Emma J. McCoy, and David A. Stephens. Quantifying the effect of area deprivation on child pedestrian casualties by using longitudinal mixed models to adjust for confounding, interference and spatial dependence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 931–950, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cockx:2013:SER

[365] Bart Cockx and Matteo Picchio. Scarring effects of remaining unemployed for long-term unemployed school-leavers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 951–980, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2013:RC

[366] Anonymous. Report of the council for 2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):981–1071, October 2013. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2013:Ob

[367] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4):1073– 1086, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2013:IAI

[368] Anonymous. Index of authors: Index of authors, volume 176, 2013. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 1087–1090, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2013:CVC

[369] Anonymous. Contents of volume: Contents of volume 176, 2013. Journal of the Royal Statistical Society. Series A (Statistics in Society), 176(4): 1091–1092, October 2013. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:RER

[370] Anonymous. Report of the editors: Report of the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):1–2, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kenett:2014:IQ

[371] Ron S. Kenett and Galit Shmueli. On information quality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):3–38, January 2014.

CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sahu:2014:HBM

[372] Sujit K. Sahu, Bernard Baffour, Paul R. Harper, John H. Minty, and Christophe Sarran. A hierarchical Bayesian model for improving short-term forecasting of hospital demand by including meteorological information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):39–61, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

deLuca:2014:WDH

[373] Federico de Luca and Giovanna Boccuzzo. What do healthcare workers know about sudden infant death syndrome?: the results of the Italian campaign ‘GenitoriPiù’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):63–82, January 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ridley:2014:ATS

[374] Dennis Ridley and Pierre Ngnepieba. Antithetic time series analysis and the CompanyX data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):83–94, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dickens:2014:UMW

[375] Richard Dickens, Rebecca Riley, and David Wilkinson. The UK minimum wage at 22 years of age: a regression discontinuity approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):

95–114, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vanSoest:2014:TSP

[376] Arthur van Soest and Hana Vonkova.

Testing the specification of parametric models by using anchoring vignettes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 115–133, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Voleti:2014:NPM

[377] Sudhir Voleti and Pulak Ghosh. A nonparametric model of residual brand equity in hierarchical branding structures with application to US beer data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 135–152, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Epstein:2014:CAL

[378] David Epstein, Antonio Olry de Labry Lima, Leticia García Mochón, Jaime Espín Balbino, and Javier Esquivias. A comparison of the accuracy of liquid cytology versus conventional screening: a meta-analysis of split-sample studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 153–168, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ahlfeldt:2014:FFE

[379] Gabriel M. Ahlfeldt and Georgios Kavetsos. Form or function?: the effect of new sports stadia on property prices in London. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):169–190, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wang:2014:BRR

[380] Hsiuying Wang and Wei Heng Huang. Bayesian ranking responses in multipleresponse questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):191–208, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2014:EPI

[381] Ruth King, Sheila M. Bird, Antony M. Overstall, Gordon Hay, and Sharon J. Hutchinson. Estimating prevalence of injecting drug users and associated heroin-related death rates in England by using regional data and incorporating prior information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):209–236, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Diya:2014:MFA

[382] Luwis Diya, Baoyue Li, Koen Van den Heede, Walter Sermeus, and Emmanuel Lesaffre. Multilevel factor analytic models for assessing the relationship between nurse-reported adverse events and patient safety. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1):237–257, January 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Soares:2014:TCD

[383] Marta O. Soares, Jo C. Dumville, A. E. Ades, and Nicky J. Welton. Treatment comparisons for decision making: facing the problems of sparse and few data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 259–279, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Beatty:2014:THE

[384] Timothy K. M. Beatty, Laura Blow, and Thomas F. Crossley. Is there a ‘heat-or-eat’ trade-off in the UK? Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 281–294, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Madan:2014:SEH

[385] Jason Madan, Yen-Fu Chen, Paul Aveyard, Dechao Wang, Ismail Yahaya, Marcus Munafo, Linda Bauld, and Nicky Welton. Synthesis of evidence on heterogeneous interventions with multiple outcomes recorded over multiple follow-up times reported inconsistently: a smoking cessation case-study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 295–314, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:Oa

[386] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(1): 315–317, January 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Miranda:2014:MOC

[387] Alfonso Miranda and Sophia Rabe-Hesketh. Missing ordinal covariate with informative selection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 319–344, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Guertzgen:2014:WIW

[388] N. Guertzgen. Wage insurance within German firms: do institutions matter? Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 345–369, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pizzi:2014:MEL

[389] Costanza Pizzi, Tim J. Cole, Camila Corvalan, Isabel dos Santos Silva, Lorenzo Richiardi, and Bianca L. De Stavola. On modelling early life weight trajectories. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):371–396, February 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McHale:2014:MEM

[390] Ian G. McHale and Łukasz Szczepański. A mixed effects model for identifying goal scoring ability of footballers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 397–417, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pertile:2014:OBS

[391] Paolo Pertile, Martin Forster, and Davide La Torre. Optimal Bayesian sequential sampling rules for the economic evaluation of health technologies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 419–438, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tranmer:2014:MMM

[392] Mark Tranmer, David Steel, and William J. Browne. Multiple-membership multiple-classification models for social network and group dependences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 439–455, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Diaz-Ordaz:2014:HMV

[393] K. Díaz-Ordaz, Michael G. Kenward, and Richard Grieve. Handling missing values in cost effectiveness analyses that use data from cluster randomized trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 457–474, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Babuscia:2014:APE

[394] Alessandra Babuscia and Kar-Ming Cheung. An approach to perform expert elicitation for engineering design risk analysis: methodology and experimental results. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):475–497, February 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Watson:2014:RES

[395] Nicole Watson and Mark Wooden. Reengaging with survey non-respondents: evidence from three household panels. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 499–522, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Deuchert:2014:ENH

[396] Eva Deuchert and Conny Wunsch. Evaluating nationwide health interventions: Malawi’s insecticide-treated-net distribution programme. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):523–552, February 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goldstein:2014:FMM

[397] Harvey Goldstein, James R. Carpenter, and William J. Browne. Fitting multilevel multivariate models with missing data in responses and covariates that may include interactions and non-linear terms. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):553–564, February 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:BRa

[398] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2): 565, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stoyanov:2014:BRBa

[399] Jordan Stoyanov. Book review: Probability for Statistics and Machine Learning: Fundamentals and Advanced Topics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):566, February 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stoyanov:2014:BRBb

[400] Jordan Stoyanov. Book reviews: Exercises and Solutions in Biostatistical Theory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):566–567, February 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lhachimi:2014:BRB

[401] Stefan K. Lhachimi. Book reviews: R Graphics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):567, February 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McDonald:2014:FNS

[402] Lynn McDonald. Florence Nightingale, statistics and the Crimean War. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177 (3):569–586, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arpino:2014:UPD

[403] Bruno Arpino, Elisabetta De Cao, and Franco Peracchi. Using panel data for partial identification of human immunodeficiency virus prevalence when infection status is missing not at random. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177 (3):587–606, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Powell:2014:MSV

[404] Helen Powell and Duncan Lee. Modelling spatial variability in concentrations of single pollutants and composite air quality indicators in health effects studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3):607–623, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

White:2014:RMI

[405] Simon R. White, Sheila M. Bird, and Richard Grieve. Review of methodological issues in cost-effectiveness analyses relating to injecting drug users, and case-study illustrations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3): 625–642, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wild:2014:TNL

[406] Phillip Wild, John Foster, and Melvin J. Hinich. Testing for non-linear and time irreversible probabilistic structure in high frequency financial time series data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3):643–659, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stanton:2014:TRS

[407] Michelle C. Stanton, Lydiane Agier, Benjamin M. Taylor, and Peter J. Diggle. Towards realtime spatiotemporal prediction of district level meningitis incidence in sub-Saharan Africa.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 177 (3):661–678, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jiang:2014:GSM

[408] Huan Jiang, Patrick E. Brown, Håvard Rue, and Silvia Shimakura. Geostatistical survival models for environmental risk assessment with large retrospective cohorts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3):679–695, June 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Demiris:2014:EFC

[409] Nikolaos Demiris, Theodore Kypraios, and L. Vanessa Smith. On the epidemic of financial crises. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3):697–723, June 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hulden:2014:AHS

[410] Lena Huldén, Ross McKitrick, and Larry Huldén. Average household size and the eradication of malaria. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3): 725–742, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:Ob

[411] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3): 743–752, June 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2014:BRBa

[412] Kuldeep Kumar. Book review: Advances in Data Analysis, Data Handling and Business Intelligence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(3):753, June 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vignoles:2014:EEH

[413] Anna Vignoles and Arnaud Chevalier. Editorial: Educational and health inequalities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):755–756, October 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Washbrook:2014:DAR

[414] Elizabeth Washbrook, Paul Gregg, and Carol Propper. A decomposition analysis of the relationship between parental income and multiple child outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 757–782, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schurer:2014:SEI

[415] Stefanie Schurer, Michael A. Shields, and Andrew M. Jones. Socio-economic inequalities in bodily pain over the life cycle: longitudinal evidence from Australia, Britain and Germany. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 783–806, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Johnston:2014:IGC

[416] David W. Johnston, Carol Propper, Stephen E. Pudney, and Michael A. Shields. The income gradient in childhood mental health: all in the eye of the beholder? Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):807–827, October 2014. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Crawford:2014:DMB

[417] Claire Crawford, Lorraine Dearden, and Ellen Greaves. The drivers of monthof-birth differences in children’s cognitive and non-cognitive skills. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 829–860, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Thiel:2014:VLI

[418] Hendrik Thiel, Stephan L. Thomsen, and Bettina Büttner. Variation of learning intensity in late adolescence and the effect on personality traits. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 861–892, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:RCR

[419] Anonymous. Report of the council for 2013: Report of the council for 2013. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 893–980, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ganguly:2014:CRA

[420] T. Ganguly, K. J. Wilson, J. Quigley, R. M. Cooke, Alessandra Babuscia, and Kar-Ming Cheung. Correspondence: Reaction to ‘an approach to perform expert elicitation for engineering design risk analysis: methodology and experimental results’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):981–985, October 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kidgell:2014:ODC

[421] John Kidgell. Obituary: David Charles Lynn Wroe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):987–988, October 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:BRb

[422] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):989, October 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2014:BRB

[423] Sebastian Dietz. Book reviews: An Introduction to Quantitative Finance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177 (4):989, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2014:BRBb

[424] Kuldeep Kumar. Book reviews: SPSS for Applied Sciences: Basic Statistical Testing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4):990, October 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:IAI

[425] Anonymous. Index of authors: Index of authors, volume 177, 2014. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 991–994, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2014:CVC

[426] Anonymous. Contents of volume: Contents of volume 177, 2014. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(4): 995–996, October 2014. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shirley:2015:HME

[427] Kenneth E. Shirley and Andrew Gelman. Hierarchical models for estimating state and demographic trends in US death penalty public opinion. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (1):1–28, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Miranda:2015:IFA

[428] María Dolores Martínez Miranda, Bent Nielsen, and Jens Perch Nielsen. Inference and forecasting in the age–period– cohort model with unknown exposure with an application to mesothelioma mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):29–55, January 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Foroni:2015:UMD

[429] Claudia Foroni, Massimiliano Marcellino, and Christian Schumacher. Unrestricted mixed data sampling (MI-DAS): MIDAS regressions with unrestricted lag polynomials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 57–82, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vassallo:2015:IEN

[430] Rebecca Vassallo, Gabriele B. Durrant, Peter W. F. Smith, and Harvey Goldstein. Interviewer effects on nonresponse propensity in longitudinal surveys: a multilevel modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (1):83–99, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Moreno-Serra:2015:BHC

[431] Rodrigo Moreno-Serra and Peter C. Smith. Broader health coverage is good for the nation’s health: evidence from country level panel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 101–124, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lloyd:2015:LCS

[432] Christopher D. Lloyd. Local cost surface models of distance decay for the analysis of gridded population data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 125–146, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Greenan:2015:DID

[433] Charlotte C. Greenan. Diffusion of innovations in dynamic networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 147–166, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koopman:2015:DBP

[434] Siem Jan Koopman and Rutger Lit.

A dynamic bivariate Poisson model for analysing and forecasting match results in the English Premier League. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 167–186, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mazzuco:2015:FAS

[435] Stefano Mazzuco and Bruno Scarpa. Fitting age-specific fertility rates by a flexible generalized skew normal probability density function. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):187–203, January 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Noufaily:2015:MRD

[436] Angela Noufaily, Yonas Ghebremichael-Weldeselassie, Doyo Gragn Enki, Paul Garthwaite, Nick Andrews, André Charlett, and Paddy Farrington. Modelling reporting delays for outbreak detection in infectious disease data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):

205–222, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keele:2015:EGR

[437] Luke Keele, Rocío Titiunik, and José R. Zubizarreta. Enhancing a geographic regression discontinuity design through matching to estimate the effect of ballot initiatives on voter turnout. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 223–239, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gile:2015:DRD

[438] Krista J. Gile, Lisa G. Johnston, and Matthew J. Salganik. Diagnostics for respondent-driven sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1): 241–269, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Perrakis:2015:BIT

[439] Konstantinos Perrakis, Dimitris Karlis, Mario Cools, and Davy Janssens. Bayesian inference for transportation origin–destination matrices: the Poisson–inverse Gaussian and other Poisson mixtures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):271–296, January 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:O

[440] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):

297–299, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:BRa

[441] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (1):300, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hewson:2015:BRB

[442] Paul Hewson. Book review: Bayesian Data Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):301, January 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pilling:2015:BRS

[443] Mark Pilling. Book review: Statistics Topics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(1):301–302, January 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kaimi:2015:BRU

[444] Irene Kaimi. Book review: Understanding Advanced Statistical Methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (1):302, January 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Levell:2015:CIF

[445] Peter Levell. Is the Carli index flawed?:

assessing the case for the new retail price index RPIJ. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2):303–336, February 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shi:2015:PIH

[446] Peng Shi and Wei Zhang. Private information in healthcare utilization: specification of a copula-based hurdle model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2): 337–361, February 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Admiraal:2015:LLM

[447] Ryan Admiraal and Mark S. Handcock. A log-linear modelling approach to assessing the consistency of ego reports of dyadic outcomes with applications to fertility and sexual partnerships. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2): 363–382, February 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aktekin:2015:AII

[448] Tevfik Aktekin and Muzaffer Musal. Analysis of income inequality measures on human immunodeficiency virus mortality: a spatiotemporal Bayesian perspective. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2):383–403, February 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Leckie:2015:MMA

[449] George Leckie and Harvey Goldstein. A multilevel modelling approach to measuring changing patterns of ethnic composition and segregation among London secondary schools, 2001–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2): 405–424, February 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Peng:2015:EHB

[450] Roger D. Peng, Arlene M. Butz, Amber J. Hackstadt, D’Ann L. Williams, Gregory B. Diette, Patrick N. Breysse, and Elizabeth C. Matsui. Estimating the health benefit of reducing indoor air pollution in a randomized environmental intervention. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2):425–443, February 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Giorgi:2015:CDM

[451] Emanuele Giorgi, Sanie S. S. Sesay, Dianne J. Terlouw, and Peter J. Diggle. Combining data from multiple spatially referenced prevalence surveys using generalized linear geostatistical models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (2):445–464, February 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reich:2015:PSS

[452] Brian J. Reich and Michael D. Porter.

Partially supervised spatiotemporal clustering for burglary crime series identification. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2):465–480, February 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Baker:2015:TVR

[453] Rose D. Baker and Ian G. McHale. Time varying ratings in Association Football: the all-time greatest team is . . .. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(2): 481–492, February 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Eickmeier:2015:CTV

[454] Sandra Eickmeier, Wolfgang Lemke, and Massimiliano Marcellino. Classical time varying factor-augmented vector autoregressive models-estimation, forecasting and structural analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3): 493–533, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lopez-Vizcaino:2015:SAE

[455] Esther López-Vizcaíno, María José Lombardía, and Domingo Morales. Small area estimation of labour force indicators under a multinomial model with correlated time and area effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):535–565, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Conigliani:2015:PPR

[456] Caterina Conigliani, Andrea Manca, and Andrea Tancredi. Prediction of patientreported outcome measures via multivariate ordered probit models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3): 567–591, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Albarran:2015:PIW

[457] Irene Albarrán, Pablo Alonso, and Aurea Grané. Profile identification via weighted related metric scaling: an application to dependent Spanish children. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):593–618, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gile:2015:NMA

[458] Krista J. Gile and Mark S. Handcock.

Network model-assisted inference from respondent-driven sampling data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3): 619–639, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hong:2015:MSD

[459] Andrew E. Hong, Corentin M. Barbu, Dylan S. Small, Michael Z. Levy, and the Chagas Disease Working Group in Arequipa. Mapping the spatial distribution of a disease-transmitting insect in the presence of surveillance error and missing data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):641–658, June 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Titman:2015:JMG

[460] A. C. Titman, D. A. Costain, P. G. Ridall, and K. Gregory. Joint modelling of goals and bookings in association football. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):659–683, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Baetschmann:2015:CEF

[461] Gregori Baetschmann, Kevin E. Staub, and Rainer Winkelmann. Consistent estimation of the fixed effects ordered logit model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):685–703, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lagona:2015:MMM

[462] Francesco Lagona, Antonello Maruotti, and Fabio Padovano. Multilevel multivariate modelling of legislative count data, with a hidden Markov chain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):705–723, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hong:2015:BEL

[463] Hyokyoung Grace Hong, Yu Yue, and Pulak Ghosh. Bayesian estimation of long-term health consequences for obese and normal-weight elderly people. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3): 725–739, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sacheti:2015:HBO

[464] Abhinav Sacheti, Ian Gregory-Smith, and David Paton. Home bias in officiating: evidence from international cricket. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):741–755, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hartman:2015:SAT

[465] Erin Hartman, Richard Grieve, Roland Ramsahai, and Jasjeet S. Sekhon. From sample average treatment effect to population average treatment effect on the treated: combining experimental with observational studies to estimate population treatment effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3): 757–778, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Heath:2015:OHH

[466] Anthony Heath. Obituary: A. H. Halsey, FBA. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):779–780, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:BRb

[467] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):781, June 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kostenko:2015:BRC

[468] Andrey Kostenko. Book review: Computational Actuarial Science with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):782–783, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2015:BRA

[469] Shalabh. Book review: Analysis of Multivariate and High Dimensional Data.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (3):783–784, June 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Green:2015:BRS

[470] Nathan Green. Book review: Statistics for People Who (Think They) Hate Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):784, June 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gillard:2015:BRS

[471] Jonathan Gillard. Book review: Sequential Analysis: Hypothesis Testing and Changepoint Detection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):785, June 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keeble:2015:BRB

[472] Claire Keeble. Book review: Bayes’ Rule: a Tutorial Introduction to Bayesian Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(3):785, June 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shlomo:2015:EBD

[473] Natalie Shlomo and Harvey Goldstein. Editorial: Big data in social research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 787–790, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Diggle:2015:SDS

[474] Peter J. Diggle. Statistics: a data science for the 21st century. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 793–813, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hancock:2015:DHS

[475] Ruth Hancock, Marcello Morciano, Stephen Pudney, and Francesca Zantomio. Do household surveys give a coherent view of disability benefit targeting?: a multisurvey latent variable analysis for the older population in Great Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):815–836, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Carriero:2015:RNB

[476] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Realtime nowcasting with a Bayesian mixed frequency model with stochastic volatility. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 837–862, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gottard:2015:RBE

[477] Anna Gottard, Alessandra Mattei, and Daniele Vignoli. The relationship between education and fertility in the presence of a time varying frailty component. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 863–881, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Boheim:2015:GRT

[478] René Böheim and Mario Lackner. Gender and risk taking: evidence from jumping competitions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):883–902, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Katki:2015:JMP

[479] Hormuzd A. Katki, Li C. Cheung, Barbara Fetterman, Philip E. Castle, and Rajeshwari Sundaram. A joint model of persistent human papilloma virus infection and cervical cancer risk: implications for cervical cancer screening. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 903–923, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ghilagaber:2015:ASB

[480] Gebrenegus Ghilagaber and Linda Wänström. Adjusting for selection bias in assessing the relationship between sibship size and cognitive performance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 925–944, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Klausch:2015:SES

[481] Thomas Klausch, Joop Hox, and Barry Schouten. Selection error in single- and mixed-mode surveys of the Dutch general population. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):945–961, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schneider:2015:NMP

[482] Matthew J. Schneider and John M.

Abowd. A new method for protecting interrelated time series with Bayesian prior distributions and synthetic data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 963–975, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wheldon:2015:BRT

[483] Mark C. Wheldon, Adrian E. Raftery, Samuel J. Clark, and Patrick Gerland. Bayesian reconstruction of two-sex populations by age: estimating sex ratios at birth and sex ratios of mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 977–1007, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ross:2015:BHM

[484] Michelle Ross and Jon Wakefield.

Bayesian hierarchical models for smoothing in two-phase studies, with application to small area estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1009–1023, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bartolucci:2015:RSJ

[485] Francesco Bartolucci, Valentino Dardanoni, and Franco Peracchi. Ranking scientific journals via latent class models for polytomous item response data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1025–1049, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Militino:2015:DSA

[486] A. F. Militino, M. D. Ugarte, and T. Goicoa. Deriving small area estimates from information technology business surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1051–1067, October 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hund:2015:BFE

[487] Lauren Hund, Edward J. Bedrick, Curtis Miller, Gabriel Huerta, Teddy Nez, Sandy Ramone, Chris Shuey, Miranda Cajero, and Johnnye Lewis. A Bayesian framework for estimating disease risk due to exposure to uranium mine and mill waste on the Navajo Nation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1069–1091, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Smith:2015:O

[488] Martha K. Smith, Stuart Reid, Max Murray, and Crawford Revie. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178 (4):1093–1096, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:BR

[489] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1097, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2015:BRM

[490] Anoop Chaturvedi. Book reviews: Modeling Count Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1098–1099, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kostenko:2015:BRU

[491] Andrey Kostenko. Book reviews: Using R for Numerical Analysis in Science and Engineering. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1098, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Okkan:2015:BRS

[492] Umut Okkan and Guül nan. Book reviews: Simulating Nature: a Philosophical Study of Computer-simulation Uncertainties and Their Role in Climate Science and Policy Advice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1099–1100, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pallmann:2015:BRA

[493] Philip Pallmann. Book reviews: Analyzing Baseball Data with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4): 1099, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gillard:2015:BRB

[494] Jonathan Gillard. Book reviews:

Bayesian and Frequentist Regression Methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1100–1101, October 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:RR

[495] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1102– 1106, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:IAV

[496] Anonymous. Index of authors, volume 178, 2015. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1107–1111, October 2015. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2015:CV

[497] Anonymous. Contents of volume 178, 2015. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1112–1114, October 2015. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Varin:2016:SMC

[498] Cristiano Varin, Manuela Cattelan, and David Firth. Statistical modelling of citation exchange between statistics journals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1):1–63, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaudhuri:2016:FDI

[499] Kausik Chaudhuri, Minjoo Kim, and Yongcheol Shin. Forecasting distributions of inflation rates: the functional auto-regressive approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 65–102, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fabrizi:2016:HLC

[500] Enrico Fabrizi, Giorgio E. Montanari, and M. Giovanna Ranalli. A hierarchical latent class model for predicting disability small area counts from survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 103–131, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tzioumis:2016:DDD

[501] Konstantinos Tzioumis. Detecting discrimination: a dynamic perspective. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 133–152, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Assaf:2016:ABR

[502] Shireen Assaf, Stefano Campostrini, Fang Xu, and Carol Gotway Crawford. Analysing behavioural risk factor surveillance data by using spatially and temporally varying coefficient models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179 (1):153–175, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sriram:2016:BQR

[503] Karthik Sriram, Peng Shi, and Pulak Ghosh. A Bayesian quantile regression model for insurance company costs data.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 177–202, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Feddersen:2016:SWW

[504] John Feddersen, Robert Metcalfe, and Mark Wooden. Subjective wellbeing: why weather matters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 203–228, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vandenBrakel:2016:SAEa

[505] Jan A. van den Brakel, Bart Buelens, and Harm-Jan Boonstra. Small area estimation to quantify discontinuities in repeated sample surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 229–250, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Berger:2016:SVE

[506] Y. G. Berger and R. Priam. A simple variance estimator of change for rotating repeated surveys: an application to the European Union Statistics on Income and Living Conditions household surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1):251–272, January 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Marie:2016:PTS

[507] Olivier Marie. Police and thieves in the stadium: measuring the (multiple) effects of football matches on crime. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 273–292, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Berchuck:2016:SMA

[508] Samuel I. Berchuck, Joshua L. Warren, Amy H. Herring, Kelly R. Evenson, Kari A. B. Moore, Yamini K. Ranchod, and Ana V. Diez-Roux. Spatially modelling the association between access to recreational facilities and exercise: the ‘multiethnic study of atherosclerosis’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 293–310, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Champkin:2016:O

[509] Julian Champkin, Lord Layard, and Steven Gilmour. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(1): 311–318, January 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keiding:2016:PPS

[510] Niels Keiding and Thomas A. Louis.

Perils and potentials of self-selected entry to epidemiological studies and surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2):319–376, February 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bollineni-Balabay:2016:MSS

[511] Oksana Bollineni-Balabay, Jan van den Brakel, and Franz Palm. Multivariate state space approach to variance reduction in series with level and variance breaks due to survey redesigns. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 377–402, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Foroni:2016:MFS

[512] Claudia Foroni and Massimiliano Marcellino. Mixed frequency structural vector auto-regressive models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 403–425, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tzavidis:2016:LAS

[513] Nikos Tzavidis, Nicola Salvati, Timo Schmid, Eirini Flouri, and Emily Midouhas. Longitudinal analysis of the strengths and difficulties questionnaire scores of the Millennium Cohort Study children in England using M-quantile random-effects regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2):427– 452, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chambers:2016:SSA

[514] Ray Chambers, Nicola Salvati, and Nikos Tzavidis. Semiparametric small area estimation for binary outcomes with application to unemployment estimation for local authorities in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 453–479, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Studer:2016:WMD

[515] Matthias Studer and Gilbert Ritschard.

What matters in differences between life trajectories: a comparative review of sequence dissimilarity measures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 481–511, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Szczepanski:2016:BCR

[516] Lukasz Szczepański and Ian McHale. Beyond completion rate: evaluating the passing ability of footballers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 513–533, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dickerson:2016:PPC

[517] Andrew Dickerson and Gurleen K. Popli.

Persistent poverty and children’s cognitive development: evidence from the UK Millennium Cohort Study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 535–558, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Battistin:2016:SIR

[518] Erich Battistin and Mario Padula. Survey instruments and the reports of consumption expenditures: evidence from the consumer expenditure surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2):

559–581, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Herrmann:2016:DES

[519] Michael Herrmann, Simon Munzert, and Peter Selb. Determining the effect of strategic voting on election results. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2): 583–605, February 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kroger:2016:CME

[520] Hannes Kröger, Rasmus Hoffmann, and Eduwin Pakpahan. Consequences of measurement error for inference in crosslagged panel design — the example of the reciprocal causal relationship between subjective health and socioeconomic status. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(2):607–628, February 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hand:2016:EBD

[521] David J. Hand. Editorial: ‘big data’ and data sharing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):629–631, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hener:2016:PSF

[522] Timo Hener, Helmut Rainer, and Thomas Siedler. Political socialization in flux?: linking family non-intactness during childhood to adult civic engagement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):633–656, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brandolini:2016:BBH

[523] Andrea Brandolini and Eliana Viviano.

Behind and beyond the (head count) employment rate. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):657–681, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kalaycioglu:2016:CMI

[524] Oya Kalaycioglu, Andrew Copas, Michael King, and Rumana Z. Omar. A comparison of multiple-imputation methods for handling missing data in repeated measurements observational studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):683–706, June 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fruhwirth-Schnatter:2016:MLR

[525] Sylvia Frühwirth-Schnatter, Christoph Pamminger, Andrea Weber, and Rudolf Winter-Ebmer. Mothers’ long-run career patterns after first birth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3): 707–725, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schouten:2016:DMB

[526] Barry Schouten, Fannie Cobben, Peter Lundquist, and James Wagner. Does more balanced survey response imply less non-response bias? Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):

727–748, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hyytinen:2016:OPP

[527] Ari Hyytinen, Pekka Ilmakunnas, and Mika Maliranta. Olley–Pakes productivity decomposition: computation and inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):749–761, June 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vandenBrakel:2016:SAEb

[528] Jan A. van den Brakel and Sabine Krieg. Small area estimation with state space common factor models for rotating panels. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179 (3):763–791, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hanly:2016:SAC

[529] Mark Hanly, Paul Clarke, and Fiona Steele. Sequence analysis of call record data: exploring the role of different cost settings. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):793–808, June 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Muller:2016:BLP

[530] Daniel Muller and Lionel Page. Born leaders: political selection and the relative age effect in the US Congress: D. Muller and L. Page. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):809–829, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

MacDonald:2016:EPP

[531] John M. MacDonald, Jonathan Klick, and Ben Grunwald. The effect of private police on crime: evidence from a geographic regression discontinuity design. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):831–846, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rohde:2016:DME

[532] Nicholas Rohde. J-divergence measurements of economic inequality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3): 847–870, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:O

[533] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3): 871–875, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:BRB

[534] Anonymous. Book review: Books for review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):877, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2016:BRD

[535] Sebastian Dietz. Book review: Data Analysis for Network Cyber-security. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):878, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2016:BRC

[536] Kuldeep Kumar. Book review: A Certain Uncertainty: Nature’s Random Ways. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):878–879, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2016:BRP

[537] Shalabh. Book review: Probability and Statistics by Example 1: Basic Probability and Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):879–880, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kostenko:2016:BRG

[538] Andrey Kostenko. Book review: Graphical Data Analysis with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):880, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Penenberg:2016:BRE

[539] Darryl Neil Penenberg. Book review: Exposure–response modeling methods and practical implementation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3): 881–882, June 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chevalier:2016:PPH

[540] A. Chevalier, F. Moscone, and J. Mullahy. Preface to the papers on ‘health econometrics’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):883–884, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cappellari:2016:DMI

[541] Lorenzo Cappellari, Anna De Paoli, and Gilberto Turati. Do market incentives for hospitals affect health and service utilization?: evidence from prospective pay system–diagnosis-related groups tariffs in Italian regions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):885–905, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Berta:2016:ABA

[542] Paolo Berta, Gianmaria Martini, Francesco Moscone, and Giorgio Vittadini. The association between asymmetric information, hospital competition and quality of healthcare: evidence from Italy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):907–926, October 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koch:2016:HFC

[543] Steven F. Koch and Jeffrey S. Racine.

Healthcare facility choice and user fee abolition: regression discontinuity in a multinomial choice setting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 927–950, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jones:2016:QMC

[544] Andrew M. Jones, James Lomas, Peter T. Moore, and Nigel Rice. A quasi-Monte-Carlo comparison of parametric and semiparametric regression methods for heavy-tailed and non-normal data: an application to healthcare costs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 951–974, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cairns:2016:PNL

[545] Andrew J. G. Cairns, David Blake, Kevin Dowd, and Amy R. Kessler. Phantoms never die: living with unreliable population data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):975– 1005, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wisniowski:2016:IMA

[546] Arkadiusz Wiśniowski, Jonathan J. Forster, Peter W. F. Smith, Jakub Bijak, and James Raymer. Integrated modelling of age and sex patterns of European migration. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1007–1024, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Simpson:2016:SEI

[547] Ludi Simpson, Stephen Jivraj, and James Warren. The stability of ethnic identity in England and Wales 2001– 2011. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1025–1049, October 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fe:2016:SLR

[548] Eduardo Fé and Bruce Hollingsworth.

Short- and long-run estimates of the local effects of retirement on health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1051–1067, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Taylor:2016:UAR

[549] James W. Taylor and Keming Yu. Using auto-regressive logit models to forecast the exceedance probability for financial risk management. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1069–1092, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Braakmann:2016:REF

[550] Nils Braakmann and John Wildman.

Reconsidering the effect of family size on labour supply: the twin problems of the twin birth instrument. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1093–1115, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:OPG

[551] Anonymous. Obituary: Peter Gavin Hall AO, 1951–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1117– 1120, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:OJM

[552] Anonymous. Obituary: Joseph Mark Gani, 1924–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1120–1122, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:OUG

[553] Anonymous. Obituary: Ulf Grenander, 1923–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1122–1124, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:OLT

[554] Anonymous. Obituary: Lajos Takács, 1924–2015. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1124–1125, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:BRc

[555] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1127, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Penenberg:2016:BRM

[556] Darryl Neil Penenberg. Book reviews: Mathematical Statistics: Basic Ideas and Selected Topics, 2nd edn, vols I and II. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179 (4):1128–1129, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2016:BRS

[557] Shalabh. Book reviews: Statistical Analysis and Data Display: an Intermediate Course with Examples in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1129–1130, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pilling:2016:BRH

[558] Mark Pilling. Book reviews: Handbook of Cluster Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1130, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2016:BRA

[559] Anoop Chaturvedi. Book reviews: Analysis of Panel Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1130–1131, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hewson:2016:BRS

[560] Paul Hewson. Book reviews: Statistical Rethinking: a Bayesian Course with Examples in R and Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1131–1132, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lewis:2016:BRC

[561] Nando Lewis. Book reviews: Circular Statistics in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1132, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2016:BRS

[562] Sebastian Dietz. Book reviews: Statistics and Data Analysis for Financial Engineering. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1132–1133, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:R

[563] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1135– 1140, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:IA

[564] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4): 1141–1145, October 2016. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2016:CV

[565] Anonymous. Contents of volume 179, 2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1147–1148, October 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:IIa

[566] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (1):1–2, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Allin:2017:NSO

[567] Paul Allin and David J. Hand. New statistics for old? — Measuring the wellbeing of the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):3–43, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Enki:2017:TPL

[568] Doyo Gragn Enki, Angela Noufaily, Paddy Farrington, Paul Garthwaite, Nick Andrews, and Andre Charlett. Taylor’s power law and the statistical modelling of infectious disease surveillance data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):45–72, January 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Young:2017:ZIM

[569] Derek S. Young, Andrew M. Raim, and Nancy R. Johnson. Zero-inflated modelling for characterizing coverage errors of extracts from the US Census Bureau’s Master Address File. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 73–97, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rocha:2017:RDS

[570] Luis E. C. Rocha, Anna E. Thorson, Renaud Lambiotte, and Fredrik Liljeros. Respondent-driven sampling bias induced by community structure and response rates in social networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 99–118, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Boulieri:2017:STM

[571] Areti Boulieri, Silvia Liverani, Kees de Hoogh, and Marta Blangiardo. A space–time multivariate Bayesian model to analyse road traffic accidents by severity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):119–139, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wong:2017:AIL

[572] Albert Wong, Hendriek Boshuizen, Johan Polder, and José António Ferreira. Assessing the inequality of lifetime healthcare expenditures: a nearest neighbour resampling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 141–160, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gross:2017:EDE

[573] Marcus Groß, Ulrich Rendtel, Timo Schmid, Sebastian Schmon, and Nikos Tzavidis. Estimating the density of ethnic minorities and aged people in Berlin: multivariate kernel density estimation applied to sensitive georeferenced administrative data protected via measurement error. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):161–183, January 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wisniowski:2017:CLF

[574] Arkadiusz Wiśniowski. Combining Labour Force Survey data to estimate migration flows: the case of migration from Poland to the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 185–202, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tourangeau:2017:ARS

[575] Roger Tourangeau, J. Michael Brick, Sharon Lohr, and Jane Li. Adaptive and responsive survey designs: a review and assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):203–223, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Angelini:2017:UDL

[576] Viola Angelini, Marco Bertoni, and Luca Corazzini. Unpacking the determinants of life satisfaction: a survey experiment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 225–246, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Foster:2017:ISE

[577] Jared C. Foster, Danping Liu, Paul S.

Albert, and Aiyi Liu. Identifying subgroups of enhanced predictive accuracy from longitudinal biomarker data by using tree-based approaches: applications to fetal growth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):247–261, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Steele:2017:MSE

[578] Fiona Steele, Paul Clarke, George Leckie, Julia Allan, and Derek Johnston.

Multilevel structural equation models for longitudinal data where predictors are measured more frequently than outcomes: an application to the effects of stress on the cognitive function of nurses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):263–283, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Beyersmann:2017:FNW

[579] Jan Beyersmann and Christine Schrade. Florence Nightingale, William Farr and competing risks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):285–293, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Barbillon:2017:SBM

[580] Pierre Barbillon, Sophie Donnet, Emmanuel Lazega, and Avner Bar-Hen. Stochastic block models for multiplex networks: an application to a multilevel network of researchers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 295–314, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Page:2017:ECS

[581] Garritt L. Page, Ernesto San Martín, Javiera Orellana, and Jorge González. Exploring complete school effectiveness via quantile value added. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 315–340, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:OVP

[582] Anonymous. Obituary: Vidyadhar Prabakhar Godambe, 1926–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 341–342, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:OMJ

[583] Anonymous. Obituary: Michael John Romer Healy, 1923–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1): 342–344, January 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:OWA

[584] Anonymous. Obituary: T. W. Anderson, 1918–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):344–346, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:ODJ

[585] Anonymous. Obituary: David John Newell, 1929–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):346–348, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:OJJ

[586] Anonymous. Obituary: S. J. (Jim) Morrison, 1917–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(1):348–350, January 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:IIb

[587] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 351–352, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nason:2017:SWS

[588] Guy P. Nason, Ben Powell, Duncan Elliott, and Paul A. Smith. Should we sample a time series more frequently?: decision support via multirate spectrum estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):353–407, February 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Papanicolas:2017:MFQ

[589] Irene Papanicolas and Alistair McGuire. Measuring and forecasting quality in English hospitals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):409–432, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Taylor:2017:SME

[590] Benjamin M. Taylor. Spatial modelling of emergency service response times. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 433–453, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gazeley:2017:PPY

[591] Ian Gazeley, Hector Gutierrez Rufrancos, Andrew Newell, Kevin Reynolds, and Rebecca Searle. The poor and the poorest, 50 years on: evidence from British Household Expenditure Surveys of the 1950s and 1960s. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 455–474, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Felfe:2017:DPB

[592] Christina Felfe and Martin Huber. Does preschool boost the development of minority children?: the case of Roma children. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):475–502, February 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chen:2017:RPI

[593] Heng Chen, Marie-Hélène Felt, and Kim P. Huynh. Retail payment innovations and cash usage: accounting for attrition by using refreshment samples. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 503–530, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vassallo:2017:SIA

[594] Rebecca Vassallo, Gabriele Durrant, and Peter Smith. Separating interviewer and area effects by using a cross-classified multilevel logistic model: simulation findings and implications for survey designs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):531–550, February 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brunton-Smith:2017:DUI

[595] Ian Brunton-Smith, Patrick Sturgis, and George Leckie. Detecting and understanding interviewer effects on survey data by using a cross-classified mixed effects location–scale model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 551–568, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Keele:2017:RBI

[596] Luke Keele, Dylan Small, and Richard Grieve. Randomization-based instrumental variables methods for binary outcomes with an application to the ‘IM-PROVE’ trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):569–586, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brun-Aguerre:2017:HWT

[597] Raphael Brun-Aguerre, Ana-Maria Fuertes, and Matthew Greenwood-Nimmo. Heads I win; tails you lose: asymmetry in exchange rate passthrough into import prices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 587–612, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vallejos:2017:BSM

[598] Catalina A. Vallejos and Mark F. J. Steel. Bayesian survival modelling of university outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):613– 631, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brasington:2017:PSC

[599] David M. Brasington and Olivier Parent. Public school consolidation: a partial observability spatial bivariate probit approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):633–656, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Albarran-Lozano:2017:DES

[600] Irene Albarrán-Lozano, Pablo J. Alonso-González, and Ana Arribas-Gil. Dependence evolution in the Spanish disabled population: a functional data analysis approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):657–677, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:OPS

[601] Anonymous. Obituary: Peter Sprent, 1923–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):679–680, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:BRa

[602] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 681, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Craven:2017:BRA

[603] Matthew Craven. Book reviews: Adversarial Risk Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):682, February 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kadhem:2017:BRH

[604] Safaa Kadhem. Book reviews: Handbook of Discrete-valued Time Series. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 682–683, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2017:BRS

[605] Thomas King. Book reviews: Sociology as a Population Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 683, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2017:BRO

[606] Sebastian Dietz. Book reviews: Option Valuation under Stochastic Volatility, vol. II. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2):683–684, February 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

DiTanna:2017:BRM

[607] Gian Luca Di Tanna. Book reviews: Missing Data Analysis in Practice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(2): 684–685, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2017:BRM

[608] Stanley E. Lazic. Book reviews: Modern Adaptive Randomized Clinical Trials: Statistical and Practical Aspects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (2):685, February 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:IIc

[609] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3): 687–688, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zeng:2017:ICM

[610] Jiaming Zeng, Berk Ustun, and Cynthia Rudin. Interpretable classification models for recidivism prediction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (3):689–722, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gheorghe:2017:HLE

[611] Maria Gheorghe, Susan Picavet, Monique Verschuren, Werner B. F. Brouwer, and Pieter H. M. van Baal. Health losses at the end of life: a Bayesian mixed beta regression approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):723–749, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aitkin:2017:SMT

[612] Murray Aitkin, Duy Vu, and Brian Francis. Statistical modelling of a terrorist network. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):751–768, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Proietti:2017:SCC

[613] Tommaso Proietti and Eric Hillebrand. Seasonal changes in central England temperatures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):769–791, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ahammer:2017:EIM

[614] Alexander Ahammer, G. Thomas Horvath, and Rudolf Winter-Ebmer. The effect of income on mortality — new evidence for the absence of a causal link. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (3):793–816, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lee:2017:EIM

[615] Myoung jae Lee. Extensive and intensive margin effects in sample selection models: racial effects on wages. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (3):817–839, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bukodi:2017:PSF

[616] Erzsébet Bukodi, John H. Goldthorpe, and Jouni Kuha. The pattern of social fluidity within the British class structure: a topological model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):

841–862, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sayal:2017:IAW

[617] Homesh Sayal, John A. D. Aston, Duncan Elliott, and Hernando Ombao. An introduction to applications of wavelet benchmarking with seasonal adjustment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):863–889, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Li:2017:LPT

[618] Li Li and Ji-Hyun Lee. A latent promotion time cure rate model using dependent tail-free mixtures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3): 891–905, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Contoyannis:2017:DAD

[619] Paul Contoyannis and Jinhu Li. The dynamics of adolescent depression: an instrumental variable quantile regression with fixed effects approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3): 907–922, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Efron:2017:OCS

[620] Bradley Efron. Obituary: Charles Stein, 1920–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):923–925, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Titterington:2017:OJA

[621] Mike Titterington. Obituary: John Aitchison, 1926–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):925–927, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kadane:2017:OSF

[622] Joseph B. (‘Jay’) Kadane. Obituary: Stephen Fienberg, 1942–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3): 927–928, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Murphy:2017:OTP

[623] Donal Murphy. Obituary: Thomas P. Linehan, 1926–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):929–930, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brien:2017:OGN

[624] Chris Brien and Roger Payne. Obituary: Graham Neil Wilkinson 1927–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (3):930–931, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Haines:2017:OTT

[625] Linda Haines and Francesca Little. Obituary: Timothy Terence Dunne, 1948–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):931–932, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tom:2017:ORN

[626] B. D. M. Tom, S. G. Thompson, S. W.

Duffy, M. J. Sweeting, and D. I. Ohlssen. Obituary: Richard Nixon, 1972–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180 (3):932–933, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

White:2017:OAM

[627] Lynda White. Obituary: Ann Mitchell, 1935–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):933–934, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fisher:2017:OMT

[628] N. I. Fisher. Obituary: Mike Tallis, 1931–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):934–936, June 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Smith:2017:C

[629] Paul A. Smith. Correspondence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3): 937–938, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Angelini:2017:CAR

[630] Marco Bertoni (on behalf of Viola Angelini and Luca Corazzini). Correspondence: Authors’ response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):938, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:BRb

[631] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):939, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2017:BRH

[632] Kuldeep Kumar. Book reviews: Happiness Explained: What Human Flourishing Is and How We Can Promote It. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):940, June 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sutton:2017:BRM

[633] Carole Sutton. Book reviews: Managing and Sharing Research Data: a Guide to Good Practice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):940–941, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pilling:2017:BRH

[634] Mark Pilling. Book reviews: Handbook of Statistical Distributions with Applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(3):941, June 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:IId

[635] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):

943–944, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Goldstein:2017:EDE

[636] Harvey Goldstein. Editorial: The Digital Economy Act and statistical research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):945–946, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Spiegelhalter:2017:TN

[637] David Spiegelhalter. Trust in numbers.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 948–965, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gelman:2017:BSO

[638] Andrew Gelman and Christian Hennig. Beyond subjective and objective in statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):967–1033, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tzavidis:2017:PPS

[639] Nikos Tzavidis, Li-Chun Zhang, Danny Pfeffermann, and Partha Lahiri. Preface to the papers on ‘small area estimation’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1035–1037, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Moura:2017:SAM

[640] Fernando A. S. Moura, André Felipe Neves, and Denise Britz do N. Silva.

Small area models for skewed Brazilian business survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1039– 1055, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ferrante:2017:SDE

[641] Maria Rosaria Ferrante and Silvia Pacei. Small domain estimation of business statistics by using multivariate skew normal models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1057–1088, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wagner:2017:NPS

[642] Julian Wagner, Ralf Münnich, Joachim Hill, Johannes Stoffels, and Thomas Udelhoven. Non-parametric small area models using shape-constrained penalized B-splines. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1089–1109, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Marhuenda:2017:PMS

[643] Yolanda Marhuenda, Isabel Molina, Domingo Morales, and J. N. K. Rao. Poverty mapping in small areas under a twofold nested error regression model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1111–1136, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Das:2017:RMS

[644] Sumonkanti Das and Ray Chambers. Robust mean-squared error estimation for poverty estimates based on the method of Elbers, Lanjouw and Lanjouw. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1137–1161, October 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schmid:2017:CSI

[645] Timo Schmid, Fabian Bruckschen, Nicola Salvati, and Till Zbiranski. Constructing sociodemographic indicators for national statistical institutes by using mobile phone data: estimating literacy rates in Senegal. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1163–1190, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arima:2017:MFH

[646] Serena Arima, William R. Bell, Gauri S.

Datta, Carolina Franco, and Brunero Liseo. Multivariate Fay–Herriot Bayesian estimation of small area means under functional measurement error. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1191–1209, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Maples:2017:ISA

[647] Jerry J. Maples. Improving small area estimates of disability: combining the American Community Survey with the Survey of Income and Program Participation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1211–1227, October 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lombardia:2017:MGA

[648] María José Lombardía, Esther López-Vizcaíno, and Cristina Rueda. Mixed generalized Akaike information criterion for small area models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1229–1252, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Torkashvand:2017:CSA

[649] Elaheh Torkashvand, Mohammad Jafari Jozani, and Mahmoud Torabi. Clustering in small area estimation with area level linear mixed models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1253–1279, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bollineni-Balabay:2017:MHB

[650] Oksana Bollineni-Balabay, Jan van den Brakel, Franz Palm, and Harm Jan Boonstra. Multilevel hierarchical Bayesian versus state space approach in time series small area estimation: the Dutch Travel Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1281–1308, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Haws:2017:OGM

[651] Tony Haws. Obituary: Geoffrey Mallin Clarke, 1928–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1309–1311, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Carpenter:2017:OBC

[652] James Carpenter. Obituary: Bob Carpenter, 1929–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1311, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ansell:2017:OMJ

[653] Jake Ansell. Obituary: Michael J. Phillips, 1943–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1311– 1312, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:R

[654] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1313– 1317, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McDonald:2017:CFN

[655] Lynn McDonald. Corrigendum: Florence Nightingale, statistics and the Crimean War. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1319, October 2017. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:IA

[656] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4): 1321–1326, October 2017. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2017:CV

[657] Anonymous. Contents of volume 180, 2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 180(4):1327–1329, October 2017. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:IIa

[658] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181 (1):1–2, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nielsen:2018:FDP

[659] Morten Ørregaard Nielsen and Sergei S. Shibaev. Forecasting daily political opinion polls using the fractionally cointegrated vector auto-regressive model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181 (1):3–33, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Khan:2018:STT

[660] Diba Khan, Lauren M. Rossen, Brady Hamilton, Erin Dienes, Yulei He, and Rong Wei. Spatiotemporal trends in teen birth rates in the USA, 2003–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181 (1):35–58, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Grant:2018:RUD

[661] Charles Grant and Mario Padula. The repayment of unsecured debt by European households. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):59–83, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yavuz:2018:ECI

[662] Idil Yavuz, Yu Chng, and Abdus S.

Wahed. Estimating the cumulative incidence function of dynamic treatment regimes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):85–106, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Baena-Mirabete:2018:PHO

[663] S. Baena-Mirabete and P. Puig. Parsimonious higher order Markov models for rating transitions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):107–131, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brown:2018:MID

[664] Sarah Brown, Mark N. Harris, Preety Srivastava, and Xiaohui Zhang. Modelling illegal drug participation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1): 133–154, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Abbate:2018:PID

[665] Angela Abbate and Massimiliano Marcellino. Point, interval and density forecasts of exchange rates with time varying parameter models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):155–179, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

West:2018:CCI

[666] Brady T. West, Frederick G. Conrad, Frauke Kreuter, and Felicitas Mittereder. Can conversational interviewing improve survey response quality without increasing interviewer effects? Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1): 181–203, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rice:2018:REF

[667] Kenneth Rice, Julian P. T. Higgins, and Thomas Lumley. A re-evaluation of fixed effect(s) meta-analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1): 205–227, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Moore:2018:DSR

[668] Jamie C. Moore, Gabriele B. Durrant, and Peter W. F. Smith. Data set representativeness during data collection in three UK social surveys: generalizability and the effects of auxiliary covariate choice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):229–248, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mazza:2018:SRH

[669] Jacopo Mazza and Hans van Ophem. Separating risk from heterogeneity in education: a semiparametric approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1): 249–275, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

DiazOrdaz:2018:MEC

[670] K. DiazOrdaz, A. J. Franchini, and R. Grieve. Methods for estimating complier average causal effects for cost-effectiveness analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1): 277–297, January 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Roberts:2018:DEC

[671] Jennifer Roberts, Gurleen Popli, and Rosemary J. Harris. Do environmental concerns affect commuting choices?: hybrid choice modelling with household survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(1):299–320, January 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:IIb

[672] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 321–322, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chevalier:2018:MIC

[673] Arnaud Chevalier, Peter Dolton, and Melanie Lührmann. ‘Making it count’: incentives, student effort and performance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):323–349, February 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Titiunik:2018:LBA

[674] Rocío Titiunik and Andrew Feher. Legislative behaviour absent re-election incentives: findings from a natural experiment in the Arkansas Senate. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 351–378, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schierholz:2018:OCD

[675] Malte Schierholz, Miriam Gensicke, Nikolai Tschersich, and Frauke Kreuter. Occupation coding during the interview. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 379–407, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sologon:2018:MED

[676] Denisa M. Sologon and Philippe Van Kerm. Modelling earnings dynamics and inequality: foreign workers and inequality trends in Luxembourg, 1988–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 409–440, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Huber:2018:DIE

[677] Martin Huber, Michael Lechner, and Anthony Strittmatter. Direct and indirect effects of training vouchers for the unemployed. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):441–463, February 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Mukhopadhyay:2018:BSM

[678] Sabyasachi Mukhopadhyay and Sujit K. Sahu. A Bayesian spatiotemporal model to estimate long-term exposure to outdoor air pollution at coarser administrative geographies in England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 465–486, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nicoletti:2018:ESS

[679] Cheti Nicoletti and Birgitta Rabe. The effect of school spending on student achievement: addressing biases in valueadded models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):487–515, February 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Plewis:2018:MRL

[680] Ian Plewis. Multiple regression, longitudinal data and welfare in the 19th century: reflections on Yule (1899). Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 517–533, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lee:2018:DEM

[681] Kwonsang Lee, Dylan S. Small, Jesse Y.

Hsu, Jeffrey H. Silber, and Paul R. Rosenbaum. Discovering effect modification in an observational study of surgical mortality at hospitals with superior nursing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):535–546, February 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wood:2018:OGH

[682] J. T. Wood. Obituary: Geoffrey Harry Freeman, 1928–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 547, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:BRa

[683] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 548, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Little:2018:BRM

[684] Mark Little. Book review: Mendelian Randomization: Methods for using Genetic Variants in Causal Estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 549–550, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hutchison:2018:BRB

[685] Dougal Hutchison. Book review: Bayesian Psychometric Modelling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2): 550, February 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2018:BRE

[686] Allan Reese. Book review: Errors, Blunders, and Lies: How to Tell the Difference, 1st edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):550–551, February 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dietz:2018:BRH

[687] Sebastian Dietz. Book review: Humanizing Big Data: Marketing at the Meeting of Data, Social Science and Consumer Insight. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(2):551–552, February 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:IIc

[688] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 553–554, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hand:2018:SCA

[689] David J. Hand. Statistical challenges of administrative and transaction data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181 (3):555–605, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Drechsler:2018:PPD

[690] Jörg Drechsler and Natalie Shlomo.

Preface to the papers on ‘Data confidentiality and statistical disclosure control’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):607–608, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lei:2018:DPM

[691] Jing Lei, Anne-Sophie Charest, Aleksandra Slavkovic, Adam Smith, and Stephen Fienberg. Differentially private model selection with penalized and constrained likelihood. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):609–633, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Manrique-Vallier:2018:BNP

[692] Daniel Manrique-Vallier and Jingchen Hu. Bayesian non-parametric generation of fully synthetic multivariate categorical data in the presence of structural zeros. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):635–647, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Quick:2018:GPS

[693] Harrison Quick, Scott H. Holan, and Christopher K. Wikle. Generating partially synthetic geocoded public use data with decreased disclosure risk by using differential smoothing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 649–661, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Snoke:2018:GSU

[694] Joshua Snoke, Gillian M. Raab, Beata Nowok, Chris Dibben, and Aleksandra Slavkovic. General and specific utility measures for synthetic data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 663–688, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Galvao:2018:SEI

[695] Antonio F. Galvao, Gabriel Montes-Rojas, Jose Olmo, and Suyong Song. On solving endogeneity with invalid instruments: an application to investment equations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):689–716, June 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dodd:2018:SMD

[696] Erengul Dodd, Jonathan J. Forster, Jakub Bijak, and Peter W. F. Smith. Smoothing mortality data: the English Life Tables, 2010–2012. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 717–735, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Powell:2018:TMP

[697] Ben Powell, Guy Nason, Duncan Elliott, Matthew Mayhew, Jennifer Davies, and Joe Winton. Tracking and modelling prices using web-scraped price microdata: towards automated daily consumer price index forecasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 737–756, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sturgis:2018:ACE

[698] Patrick Sturgis, Jouni Kuha, Nick Baker, Mario Callegaro, Stephen Fisher, Jane Green, Will Jennings, Benjamin E. Lauderdale, and Patten Smith. An assessment of the causes of the errors in the 2015 UK general election opinion polls. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):757–781, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Warren:2018:FAS

[699] Joshua L. Warren and Penny Gordon-Larsen. Factors associated with supermarket and convenience store closure: a discrete time spatial survival modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):783–802, June 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Liu:2018:AMB

[700] Jingyu Liu, Walter W. Piegorsch, A. Grant Schissler, and Susan L. Cutter. Autologistic models for benchmark risk or vulnerability assessment of urban terrorism outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):803–823, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ortega-Villa:2018:EOT

[701] Ana Maria Ortega-Villa, Katherine L. Grantz, and Paul S. Albert. Estimating onset time from longitudinal and cross-sectional data with an application to estimating gestational age from longitudinal maternal anthropometry during pregnancy and neonatal anthropometry at birth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):825–842, June 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Phillippo:2018:STR

[702] David M. Phillippo, Sofia Dias, A. E.

Ades, Vanessa Didelez, and Nicky J. Welton. Sensitivity of treatment recommendations to bias in network metaanalysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):843–867, June 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Buta:2018:BJM

[703] Eugenia Buta, Stephanie S. O’Malley, and Ralitza Gueorguieva. Bayesian joint modelling of longitudinal data on abstinence, frequency and intensity of drinking in alcoholism trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 869–888, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zilic:2018:EFD

[704] Ivan Zilic. Effect of forced displacement on health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):889–906, June 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Swan:2018:ODJ

[705] Celia Swan and Martin Knott. Obituary: David John Bartholomew, 1931– 2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):907–909, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Walton:2018:OSF

[706] John Walton. Obituary: Stanley Francis James, 1927–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):909–910, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stillman:2018:ONR

[707] Eleanor Stillman. Obituary: Nicholas Rodney James Fieller, 1947–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 910–911, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chinn:2018:OWW

[708] Susan Chinn. Obituary: Walter Werner Holland, 1929–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 911–912, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Craig:2018:OFE

[709] John Craig and John Fox. Obituary: Frank Ernest Whitehead, 1930– 2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):913–914, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:BRb

[710] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):915, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2018:BRR

[711] Morteza Aalabaf-Sabaghi. Book review: Risk Theory: a Heavy Tail Approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):916, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2018:BRB

[712] Kuldeep Kumar. Book review: Big Data and Social Sciences: a Practical Guide to Methods and Tools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):916, June 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2018:BRB

[713] Thomas King. Book review: Bit by Bit: Social Research in the Digital Age. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3):917, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Walters:2018:BRF

[714] Peter Walters. Book review: Foundations of Ergodic Theory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(3): 917–919, June 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:IId

[715] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 921–922, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nason:2018:ESF

[716] Guy Nason. Editorial: Statistical flaws in the teaching excellence and student outcomes framework in UK higher education. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):923–925, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tzavidis:2018:SFF

[717] Nikos Tzavidis, Li-Chun Zhang, Angela Luna, Timo Schmid, and Natalia Rojas-Perilla. From start to finish: a framework for the production of small area official statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):927–979, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sverchkov:2018:SAE

[718] Michael Sverchkov and Danny Pfeffermann. Small area estimation under informative sampling and not missing at random non-response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 981–1008, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Geenens:2018:NPE

[719] Gery Geenens and Thomas Cuddihy. Non-parametric evidence of second-leg home advantage in European football. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1009–1031, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stecklov:2018:CII

[720] Guy Stecklov, Alexander Weinreb, and Calogero Carletto. Can incentives improve survey data quality in developing countries?: results from a field experiment in India. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1033–1056, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Durrant:2018:WSP

[721] Gabriele B. Durrant and Sylke V.

Schnepf. Which schools and pupils respond to educational achievement surveys?: a focus on the English Programme for International Student Assessment sample. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1057–1075, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Marchetti:2018:UTL

[722] Stefano Marchetti, Maciej Beresewicz, Nicola Salvati, Marcin Szymkowiak, and Lukasz Wawrowski. The use of a threelevel M-quantile model to map poverty at local administrative unit 1 in Poland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1077–1104, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hickey:2018:CJM

[723] Graeme L. Hickey, Pete Philipson, Andrea Jorgensen, and Ruwanthi Kolamunnage-Dona. A comparison of joint models for longitudinal and competing risks data, with application to an epilepsy drug randomized controlled trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1105–1123, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Taylor:2018:CIA

[724] Benjamin M. Taylor, Ricardo Andrade-Pacheco, and Hugh J. W. Sturrock. Continuous inference for aggregated point process data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1125–1150, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yu:2018:DBA

[725] Hongjian Yu, Yueyan Wang, Jean Opsomer, Pan Wang, and Ninez A. Ponce. A design-based approach to small area estimation using a semiparametric generalized linear mixed model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1151–1167, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kuha:2018:LVM

[726] Jouni Kuha, Myrsini Katsikatsou, and Irini Moustaki. Latent variable modelling with non-ignorable item nonresponse: multigroup response propensity models for cross-national analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1169–1192, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Buchanan:2018:GER

[727] Ashley L. Buchanan, Michael G. Hudgens, Stephen R. Cole, Katie R. Mollan, Paul E. Sax, Eric S. Daar, Adaora A. Adimora, Joseph J. Eron, and Michael J. Mugavero. Generalizing evidence from randomized trials using inverse probability of sampling weights. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1193–1209, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See corrigendum [854].

Moore:2018:CRL

[728] Jamie C. Moore, Peter W. F. Smith, and Gabriele B. Durrant. Correlates of record linkage and estimating risks of non-linkage biases in business data sets. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1211–1230, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kang:2018:IIV

[729] Hyunseung Kang, Laura Peck, and Luke Keele. Inference for instrumental variables: a randomization inference approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1231–1254, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Laycock:2018:OTS

[730] Patrick J. Laycock. Obituary: Tata Subba Rao, 1942–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1255– 1256, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Curnow:2018:ODJ

[731] Robert Curnow. Obituary: Derek John Pike, 1944–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1256–1257, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:BRc

[732] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1259, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2018:BRM

[733] Thomas King. Book review: Metric Power. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1260, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dialsingh:2018:BRH

[734] Isaac Dialsingh. Book review: Hurricane Climatology: a Modern Statistical Guide using R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1260–1261, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Leurent:2018:BRM

[735] Baptiste Leurent. Book review: The Mata Book: a Book for Serious Programmers and Those Who Want To Be. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181 (4):1261, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2018:BRG

[736] R. Allan Reese. Book review: Graphics for Statistics and Data Analysis with R, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1261–1262, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sheldon:2018:BRT

[737] Neil Sheldon. Book review: Teaching Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1263, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2018:BRS

[738] Kuldeep Kumar. Book review: Statistical Regression and Classification: from Linear Models to Machine Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1263–1264, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pilling:2018:BRH

[739] Mark Pilling. Book review: Handbook of Applied Modelling: Non-Gaussian and Correlated Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1264–1265, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:R

[740] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1266– 1270, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Farebrother:2018:SIO

[741] Richard W. Farebrother and Peter J.

Neal. Subject index of obituary notices 2006–2017. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1271–1273, October 2018. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:IA

[742] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4): 1275–1280, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2018:CV

[743] Anonymous. Contents of volume 181, 2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1281–1283, October 2018. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:IIa

[744] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (1):1–2, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bristle:2019:CPF

[745] Johanna Bristle, Martina Celidoni, Chiara Dal Bianco, and Guglielmo Weber. The contributions of paradata and features of respondents, interviewers and survey agencies to panel co-operation in the Survey of Health, Ageing and Retirement in Europe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):3–35, January 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bellani:2019:LRE

[746] Luna Bellani and Michela Bia. The long-run effect of childhood poverty and the mediating role of education. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 37–68, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hepenstrick:2019:FGD

[747] Christian Hepenstrick and Massimiliano Marcellino. Forecasting gross domestic product growth with large unbalanced data sets: the mixed frequency three-pass regression filter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 69–99, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Koukounari:2019:LVM

[748] Artemis Koukounari, Andrew J. Copas, and Andrew Pickles. A latent variable modelling approach for the pooled analysis of individual participant data on the association between depression and chlamydia infection in adolescence and young adulthood in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 101–130, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aristidou:2019:EUR

[749] Chrystalleni Aristidou, Kevin Lee, and Kalvinder Shields. Evaluating the use of realtime data in forecasting output levels and recessionary events in the USA. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 182(1): 131–163, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Casarin:2019:BTV

[750] Roberto Casarin, German Molina, and Enrique ter Horst. A Bayesian time varying approach to risk neutral density estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):165–195, January 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pareek:2019:AYD

[751] Bhuvanesh Pareek, Qiang Liu, and Pulak Ghosh. Ask your doctor whether this product is right for you: a Bayesian joint model for patient drug requests and physician prescriptions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 197–223, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rigon:2019:BSM

[752] Tommaso Rigon, Daniele Durante, and Nicola Torelli. Bayesian semiparametric modelling of contraceptive behaviour in India via sequential logistic regressions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 225–247, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shang:2019:VRC

[753] Han Lin Shang. Visualizing rate of change: an application to age-specific fertility rates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):249–262, January 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bach:2019:PPS

[754] Ruben L. Bach and Stephanie Eckman.

Participating in a panel survey changes respondents’ labour market behaviour. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 263–281, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Erciulescu:2019:MBC

[755] Andreea L. Erciulescu, Nathan B.

Cruze, and Balgobin Nandram. Modelbased county level crop estimates incorporating auxiliary sources of information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):283–303, January 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Marquart:2019:MMR

[756] Louise Marquart and Michele Haynes.

Misspecification of multimodal randomeffect distributions in logistic mixed models for panel survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 305–321, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Python:2019:BAM

[757] André Python, Janine B. Illian, Charlotte M. Jones-Todd, and Marta Blangiardo. A Bayesian approach to modelling subnational spatial dynamics of worldwide non-state terrorism, 2010– 2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):323–344, January 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McVicar:2019:EID

[758] Duncan McVicar, Julie Moschion, and Jan C. van Ours. Early illicit drug use and the age of onset of homelessness. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 345–372, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:OIM

[759] Anonymous. Obituary: Ian Maclean MBE, MA, 1930–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1):373– 375, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:ODG

[760] Anonymous. Obituary: Douglas Graham Altman, 1948–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(1): 375–376, January 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:IIb

[761] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 377–378, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Castruccio:2019:VSM

[762] Stefano Castruccio, Marc G. Genton, and Ying Sun. Visualizing spatiotemporal models with virtual reality: from fully immersive environments to applications in stereoscopic view. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 379–387, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gabry:2019:VBW

[763] Jonah Gabry, Daniel Simpson, Aki Vehtari, Michael Betancourt, and Andrew Gelman. Visualization in Bayesian workflow. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):389–402, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bowman:2019:GU

[764] Adrian W. Bowman. Graphics for uncertainty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):403–418, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:DMD

[765] Anonymous. Discussion on the meeting on ‘Data visualization’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 419–441, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

deBresser:2019:PPS

[766] Jochem de Bresser and Arthur van Soest. The predictive power of subjective probabilities: probabilistic and deterministic polling in the Dutch 2017 election. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):443–466, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bon:2019:PBU

[767] Joshua J. Bon, Timothy Ballard, and Bernard Baffour. Polling bias and undecided voter allocations: US presidential elections, 2004–2016. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 467–493, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dotto:2019:DIL

[768] Francesco Dotto, Alessio Farcomeni, Maria Grazia Pittau, and Roberto Zelli. A dynamic inhomogeneous latent state model for measuring material deprivation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):495–516, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hetschko:2019:EPS

[769] Clemens Hetschko, Louisa von Reumont, and Ronnie Schöb. Embedding as a pitfall for survey-based welfare indicators: evidence from an experiment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 517–539, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Grazian:2019:MPD

[770] Clara Grazian, Fabrizio Leisen, and Brunero Liseo. Modelling preference data with the Wallenius distribution. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 541–558, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Driffield:2019:BFI

[771] Nigel Driffield and Michail Karoglou.

Brexit and foreign investment in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 559–582, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Iliev:2019:PRT

[772] Iliyan R. Iliev, Xin Huang, and Yulia R. Gel. Political rhetoric through the lens of non-parametric statistics: are our legislators that different? Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 583–604, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Johansson:2019:WAP

[773] Per Johansson, Arizo Karimi, and J. Peter Nilsson. Worker absenteeism: peer influences, monitoring and job flexibility. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (2):605–621, February 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cro:2019:IAS

[774] Suzie Cro, James R. Carpenter, and Michael G. Kenward. Informationanchored sensitivity analysis: theory and application. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):623–645, February 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Yang:2019:MRM

[775] Lu Yang and Peng Shi. Multiperil rate making for property insurance using longitudinal data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):647–668, February 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Brick:2019:EEM

[776] J. Michael Brick and Sharon Lohr. Experimental evaluation of mail questionnaires in a probability sample on victimization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):669–687, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Alexopoulos:2019:BFM

[777] Angelos Alexopoulos, Petros Dellaportas, and Jonathan J. Forster. Bayesian forecasting of mortality rates by using latent Gaussian models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 689–711, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:BRa

[778] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 713, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gillard:2019:BRL

[779] Jonathan Gillard. Book review: Large Covariance and Autocovariance Matrices; Patterned Random Matrices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 714, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2019:BRB

[780] Tom King. Book review: Bulk Collection: Systematic Government Access to Private-sector Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 714–715, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2019:BRP

[781] Morteza Aalabaf-Sabaghi. Book review:

A Practical Guide to Age-Period-Cohort Analysis: the Identification Problem and Beyond. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2):715–716, February 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2019:BRL

[782] Anoop Chaturvedi. Book review: Linear Models and the Relevant Distributions and Matrix Algebra. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(2): 716, February 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:IIc

[783] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 717–718, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Britton:2019:CSS

[784] Jack Britton, Neil Shephard, and Anna Vignoles. A comparison of sample survey measures of earnings of English graduates with administrative data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (3):719–754, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Fisher:2019:CAP

[785] N. I. Fisher. A comprehensive approach to problems of performance measurement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):755–803, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Milivinti:2019:BSA

[786] Alice Milivinti and Giacomo Benini. A Bayesian semiparametric approach for trend-seasonal interaction: an application to migration forecasts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 805–830, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cuaresma:2019:SUM

[787] Jesús Crespo Cuaresma, Gernot Doppelhofer, Martin Feldkircher, and Florian Huber. Spillovers from US monetary policy: evidence from a time varying parameter global vector auto-regressive model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):831–861, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lloyd:2019:SAF

[788] Chris J. Lloyd, Raymond Kwok, and Paul S. F. Yip. A scenario analysis of future Hong Kong age and labour force profiles and its implications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 863–886, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Izzeldin:2019:MSV

[789] Marwan Izzeldin, Mike G. Tsionas, and Panayotis G. Michaelides. Multivariate stochastic volatility with large and moderate shocks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):887–917, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Zhuang:2019:SSH

[790] Jiancang Zhuang and Jorge Mateu. A semiparametric spatiotemporal Hawkestype point process model with periodic background for crime data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 919–942, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Song:2019:PDO

[791] Yunting Song and Nuo Wang. On probability distributions of the operational law of container liner ships. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (3):943–961, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rosenblum:2019:ADS

[792] Michael Rosenblum, Peter Miller, Benjamin Reist, Elizabeth A. Stuart, Michael Thieme, and Thomas A. Louis. Adaptive design in surveys and clinical trials: similarities, differences and opportunities for cross-fertilization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (3):963–982, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Geneletti:2019:BMB

[793] Sara Geneletti, Federico Ricciardi, Aidan G. O’Keeffe, and Gianluca Baio. Bayesian modelling for binary outcomes in the regression discontinuity design. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (3):983–1002, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kim:2019:SAR

[794] Keunseo Kim, Vinnam Kim, and Heeyoung Kim. Spatiotemporal autoregressive model for origin-destination air passenger flows. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1003–1016, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

vonAuer:2019:CIT

[795] Ludwig von Auer, Andranik Stepanyan, and Mark Trede. Classifying industries into types of relative concentration. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1017–1037, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

White:2019:PSM

[796] Philip A. White, Alan E. Gelfand, Eliane R. Rodrigues, and Guadalupe Tzintzun. Pollution state modelling for Mexico City. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1039–1060, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jack:2019:ECN

[797] Eilidh Jack, Duncan Lee, and Nema Dean. Estimating the changing nature of Scotland’s health inequalities by using a multivariate spatiotemporal model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1061–1080, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Rougier:2019:CRA

[798] Jonathan Rougier. Confidence in risk assessments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1081–1095, June 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:DF

[799] Anonymous. David Finney, 1917–2018.

Journal of the Royal Statistical Society.

Series A (Statistics in Society), 182(3): 1097–1099, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:AFB

[800] Anonymous. Alfred Frederick Bissell, 1927–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1100–1101, June 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:DVH

[801] Anonymous. David Victor Hinkley, 1944–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1101–1103, June 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:AAG

[802] Anonymous. Anthony Aynsley Greenfield, 1931–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1103–1105, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:JD

[803] Anonymous. John Disney, 1960–2018. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 1105–1106, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:BRb

[804] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1107, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2019:BREa

[805] Thomas King. Book review: Ethical Reasoning in Big Data: an Exploratory Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1108, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Aalabaf-Sabaghi:2019:BRM

[806] Morteza Aalabaf-Sabaghi. Book review:

Multistate Models for the Analysis of Life History Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1108–1109, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Haigh:2019:BRA

[807] John Haigh. Book review: Applied Probability: from Random Sequences to Stochastic Processes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1109, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2019:BRD

[808] Allan Reese. Book review: Data Visualization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1109–1110, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hutchison:2019:BRH

[809] Dougal Hutchison. Book review: Handbook of Item Response Theory, vols 1– 3. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (3):1110–1113, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2019:BREb

[810] Thomas King. Book review: The Ethics of Biomedical Big Data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 1113–1114, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gibson:2019:BRS

[811] Martin G. Gibson. Book review: Statistical Process Control. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3): 1114–1115, June 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2019:BRB

[812] Stanley Lazic. Book review: Bayesian Regression Modeling with INLA. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(3):1115, June 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:IId

[813] Anonymous. Issue information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1117–1118, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ashby:2019:PMS

[814] Deborah Ashby. Pigeonholes and mustard seeds: growing capacity to use data for society. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1121–1137, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Crook:2019:PPC

[815] Jonathan Crook, Tony Bellotti, Christophe Mues, and Ana-Maria Fuertes. Preface to the papers on ‘Credit risk modelling’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1139–1142, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ogundimu:2019:PDP

[816] Emmanuel O. Ogundimu. Prediction of default probability by using statistical models for rare events. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1143–1162, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Calabrese:2019:JSM

[817] Raffaella Calabrese, Silvia Angela Osmetti, and Luca Zanin. A joint scoring model for peer-to-peer and traditional lending: a bivariate model with copula dependence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1163–1188, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Coolen-Maturi:2019:NPP

[818] T. Coolen-Maturi and F. P. A. Coolen.

Non-parametric predictive inference for the validation of credit rating systems. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1189–1204, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Vinciotti:2019:EIF

[819] Veronica Vinciotti, Elisa Tosetti, Francesco Moscone, and Mark Lycett. The effect of interfirm financial transactions on the credit risk of small and medium-sized enterprises. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1205–1226, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tobback:2019:RCS

[820] Ellen Tobback and David Martens. Retail credit scoring using fine-grained payment data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1227–1246, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Tomarchio:2019:MLG

[821] Salvatore D. Tomarchio and Antonio Punzo. Modelling the loss given default distribution via a family of zeroand-one inflated mixture models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1247–1266, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Thackham:2019:EDC

[822] Mark Thackham and Jun Ma. Exposure at default without conversion factorsevidence from Global Credit Data for large corporate revolving facilities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1267–1286, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Andreeva:2019:LEO

[823] Galina Andreeva and Anna Matuszyk.

The law of equal opportunities or unintended consequences?: the effect of unisex risk assessment in consumer credit. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1287–1311, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Masci:2019:SME

[824] Chiara Masci, Anna Maria Paganoni, and Francesca Ieva. Semiparametric mixed effects models for unsupervised classification of Italian schools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1313–1342, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Qin:2019:MCM

[825] Xu Qin, Guanglei Hong, Jonah Deutsch, and Edward Bein. Multisite causal mediation analysis in the presence of complex sample and survey designs and non-random non-response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1343–1370, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kaputa:2019:EET

[826] Stephen J. Kaputa, Katherine Jenny Thompson, and Jennifer L. Beck. An embedded experiment for targeted nonresponse follow-up in establishment surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1371–1391, October 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gorgi:2019:AFT

[827] P. Gorgi, S. J. Koopman, and R. Lit. The analysis and forecasting of tennis matches by using a high dimensional dynamic model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1393–1409, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Angel:2019:WDY

[828] Stefan Angel, Franziska Disslbacher, Stefan Humer, and Matthias Schnetzer. What did you really earn last year?: explaining measurement error in survey income data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1411–1437, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arima:2019:ULS

[829] Serena Arima and Silvia Polettini. A unit level small area model with misclassified covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1439–1462, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Boeschoten:2019:ENS

[830] Laura Boeschoten, Ton de Waal, and Jeroen K. Vermunt. Estimating the number of serious road injuries per vehicle type in The Netherlands by using multiple imputation of latent classes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1463–1486, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chatterjee:2019:EPS

[831] Kiranmoy Chatterjee and Prajamitra Bhuyan. On the estimation of population size from a dependent triple-record system. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1487–1501, October 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lu:2019:FPR

[832] Yang Lu. Flexible (panel) regression models for bivariate count-continuous data with an insurance application. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1503–1521, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Li:2019:BAD

[833] Johnny Siu-Hang Li, Kenneth Q. Zhou, Xiaobai Zhu, Wai-Sum Chan, and Felix Wai-Hon Chan. A Bayesian approach to developing a stochastic mortality model for China. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1523–1560, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Angelov:2019:PCA

[834] Nikolay Angelov, Per Johansson, and Myoung jae Lee. Practical causal analysis for the treatment timing effect on doubly censored duration: effect of fertility on work span. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1561–1585, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lansley:2019:CLC

[835] Guy Lansley, Wen Li, and Paul A. Longley. Creating a linked consumer register for granular demographic analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1587–1605, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bacci:2019:LCG

[836] Silvia Bacci, Francesco Bartolucci, Giulia Bettin, and Claudia Pigini. A latent class growth model for migrants’ remittances: an application to the German Socio-Economic Panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1607– 1632, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Erskine:2019:EHS

[837] Ralph Erskine. Edward Hugh Simpson, 1922–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1633–1634, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lockhart:2019:MAS

[838] Richard Lockhart. Michael Arthur Stephens, 1927–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1634– 1636, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pickles:2019:GD

[839] Andrew Pickles. Graham Dunn, 1949– 2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1636–1637, October 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Henderson:2019:RJB

[840] Robin Henderson, John Matthews, and Darren Wilkinson. Richard James Boys, 1960–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1637–1639, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ross:2019:JCG

[841] Gavin Ross. John Clifford Gower, 1930– 2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1639–1641, October 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Muller:2019:JTW

[842] Warren Müller and Mark Westcott. Jeffrey Thomas Wood, 1942–2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1641–1642, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:BR

[843] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1643, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

McConway:2019:BRB

[844] Kevin McConway. Book review: Bayesian Data Analysis for Animal Scientists: the Basics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1644, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lazic:2019:BRL

[845] Stanley Lazic. Book review: Likelihood Methods in Biology and Ecology: a Modern Approach to Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1644–1645, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bird:2019:BRW

[846] Sheila Bird. Book review: War Stories from the Drug Survey (How Culture, Politics, and Statistics Shaped the National Survey on Drug Use and Health). Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1645–1646, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2019:BRH

[847] Shalabh. Book review: Handbook of Big Data Analytics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1646–1647, October 2019. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2019:BRM

[848] R. Allan Reese. Book review: Modern Statistics for Modern Biology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):

1647, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

King:2019:BRP

[849] Tom King. Book review: The Politics of Big Data: Big Data, Big Brother? Journal of the Royal Statistical Society. Series A (Statistics in Society), 182 (4):1648, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Iliadis:2019:BRR

[850] Ilias Iliadis. Book review: Reliability and Availability Engineering: Modeling, Analysis, and Applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1648–1649, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:R

[851] Anonymous. Referees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1651– 1656, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:IA

[852] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4): 1657–1662, October 2019. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2019:CV

[853] Anonymous. Contents of volume 182, 2019. Journal of the Royal Statistical Society. Series A (Statistics in Society), 182(4):1663–1665, October 2019. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Buchanan:2022:CGE

[854] Ashley L. Buchanan and Michael G. Hudgens. Corrigendum: Generalizing evidence from randomized trials using inverse probability of sampling weights. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1): 443–444, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [727].

Bickel:2015:MSB

[855] Peter J. Bickel and Kjell A. Doksum. Mathematical Statistics: Basic Ideas and Selected Topics. Texts in statistical science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2015. ISBN 1-4987-2380-2 (vol. 1), 1-4987-2268-7 (vol. 2). xxiv + 556 (vol. 1), xix + 465 (vol.

2) pp. LCCN QA276 .B47 2015.

Bloomfield:2014:URN

[856] Victor A. Bloomfield. Using R for Numerical Analysis in Science and Engineering. Chapman and Hall/CRC the R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2014. ISBN 1-4398-8448-X (hardcover). xxii + 335 pp. LCCN Q183.9 .B56

2014. URL http://www.crcnetbase.
com/isbn/9781439884492.

Gelman:2014:BDA

[857] Andrew Gelman, John B. Carlin, Hal S. Stern, David B. Dunson, Aki Vehtari, and Donald B. Rubin. Bayesian data analysis. Chapman and Hall/CRC texts in statistical science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2014. ISBN 1-4398-4095-4 (hardcover). xiv + 661 pp. LCCN QA279.5 .G45 2014. URL

http://images.tandf.co.uk/common/
jackets/websmall/978143984/9781439840955.
jpg.

Heiberger:2015:SAD

[858] Richard M. Heiberger and Burt Holland. Statistical Analysis and Data Display: an Intermediate Course with Examples in R. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2015. ISBN 1-4939-2121-5 (hardcover), 1-4939-2122-3 (e-book). xxxi + 898 pp. LCCN QA276.4 .H445 2015.

Hennig:2016:HCA

[859] Christian M. Hennig, Marina Meila, Fionn Murtagh, and Roberto Rocci, editors. Handbook of Cluster Analysis, volume 9 of Chapman and Hall/CRC handbooks of modern statistical methods. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4665-5188-7. xx + 753 pp. LCCN QA278 .H3456 2016.

Hilbe:2014:MCD

[860] Joseph M. Hilbe. Modeling Count Data. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02833-7 (hardcover), 1-107-61125-3 (paperback). xv + 283 pp. LCCN QA278 .H56 2014.

Hsiao:2014:APD

[861] Cheng Hsiao. Analysis of Panel Data.

Econometric society monographs. Cambridge University Press, Cambridge, UK, third edition, 2014. ISBN 1-107-03869-3 (hardcover), 1-107-65763-6 (paperback). xxi + 538 pp. LCCN HB139 .H75 2014.

Koch:2014:AMH

[862] Inge Koch. Analysis of Multivariate and High-Dimensional Data. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-88793-3 (hardcover). xxv + 504 pp. LCCN QA278 .K5935 2014.

Marchi:2014:ABD

[863] Max Marchi and Jim Albert. Analyzing baseball data with R. R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-7022-9 (paperback). xvii + 333 pp. LCCN GV877 .M353 2014. URL

http://images.tandf.co.uk/common/
jackets/websmall/978146657/9781466570221.
jpg.

McElreath:2016:SRB

[864] Richard McElreath. Statistical rethinking: a Bayesian course with examples in R and Stan, volume 122 of Chapman and Hall/CRC texts in statistical science series. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4822-5344-5 (hardcover). xvii + 469 pp. LCCN QA279.5 .M3975 2016.

Petersen:2012:SNP

[865] Arthur C. (Arthur Caesar) Petersen. Simulating nature: a philosophical study of computer-simulation uncertainties and their role in climate science and policy advice. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2012. ISBN 1-4665-0062-X. xvi + 208 pp. LCCN QC981.45 .P48

2012. URL http://www.crcnetbase.
com/isbn/9781466500624.

Pewsey:2013:CSR

[866] Arthur Pewsey, Markus Neuhäuser, and Graeme D. Ruxton. Circular Statistics in R. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967113-3. xiv + 183 pp. LCCN QA276

.P45 2013. URL http://www.loc.
gov/catdir/enhancements/fy1410/
2013940576-b.html;   http://www.
loc.gov/catdir/enhancements/fy1410/
2013940576-d.html;   http://www.
loc.gov/catdir/enhancements/fy1410/
2013940576-t.html.

Ruppert:2011:SDA

[867] David Ruppert. Statistics and Data Analysis for Financial Engineering. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-7786-4. xxii + 638 pp. LCCN HG176.7

.R87 2011. URL http://www.loc.
gov/catdir/enhancements/fy1201/
2011282136-b.html;   http://www.
loc.gov/catdir/enhancements/fy1201/
2011282136-d.html;   http://www.
loc.gov/catdir/enhancements/fy1201/
2011282136-t.html.

Salkind:2014:SPW

[868] Neil J. Salkind. Statistics for people who (think they) hate statistics. SAGE Publications, Inc., Thousand Oaks, CA, USA, fifth edition, 2014. ISBN 1-4522-7771-0 (paperback), 1-4833-1288-7 (web PDF). xxvii + 483 pp. LCCN HA29 .S2365 2014.

Stone:2013:BRT

[869] James V. Stone. Bayes’ rule: a tutorial introduction to Bayesian analysis. Sebtel Press, Lexington, KY, USA, 2013. ISBN 0-9563728-4-8 (paperback). 170 pp. LCCN QA279.5 .S766 2013.

Tartakovsky:2015:SAH

[870] Alexander Tartakovsky, I. V. (Igor Vladimirovich) Nikiforov, and M. (Michèle) Basseville. Sequential Analysis: Hypothesis Testing and Changepoint Detection, volume 136 of Monographs on statistics and applied probability. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4398-3820-8 (hardcover), 1-4398-3821-6 (e-book). xxiii + 579 pp. LCCN QA279.7 .T37

                                        2015. URL http://www.crcnetbase.
                                        com/isbn/9781439838211.

Wakefield:2013:BFR

[871] Jon Wakefield. Bayesian and frequentist regression methods. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4419-0924-9. ISSN 0172-7397. xix + 697 pp. LCCN QA278.2 .W35 2013.

Westfall:2013:UAS

[872] Peter H. Westfall and Kevin S. S. Henning. Understanding Advanced Statistical Methods. Texts in statistical science series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2013. ISBN 1-4665-1210-5 (hardcover). xxv + 543 pp. LCCN QA276 .W4537 2013. URL

                                        http://images.tandf.co.uk/common/
                                        jackets/websmall/978146651/9781466512108.
                                        jpg.

Charpentier:2015:CAS

[873] Arthur Charpentier, editor. Computational Actuarial Science with R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-9259-1 (hardcover). xxxi + 618 pp. LCCN HG8781 .C637 2014.